### v2

# HyperMARL: Adaptive Hypernetworks for Multi-Agent RL

Abstract
Adaptability is critical in cooperative multi-agent reinforcement learning (MARL), where agents must learn specialised or homogeneous behaviours for diverse tasks. While parameter sharing methods are sample-efficient, they often encounter gradient interference among agents, limiting their behavioural diversity. Conversely, non-parameter sharing approaches enable specialisation, but are computationally demanding and sample-inefficient. To address these issues, we propose HyperMARL, a parameter sharing approach that uses hypernetworks to dynamically generate agent-specific actor and critic parameters, without altering the learning objective or requiring preset diversity levels. By decoupling observation- and agent-conditioned gradients, HyperMARL empirically reduces policy gradient variance and facilitates specialisation within FuPS, suggesting it can mitigate cross-agent interference. Across multiple MARL benchmarks involving up to twenty agents—and requiring homogeneous, heterogeneous, or mixed behaviours—HyperMARL consistently performs competitively with fully shared, non-parameter-sharing, and diversity-promoting baselines, all while preserving a behavioural diversity level comparable to non-parameter sharing. These findings establish hypernetworks as a versatile approach for MARL across diverse environments. [^1] All code will be made available soon.
Multi-Agent Reinforcement Learning, Parameter Sharing, Specialisation, Behavioural Diversity
1 Introduction
Figure 1: HyperMARL Policy Architecture. Common Agent-ID conditioned shared MARL policy (left) vs HyperMARL (right), which uses a hypernetwork to generate agent-specific policies and decouples observation- and agent-conditioned gradients.
Specialist and generalist behaviours are critical to collective intelligence. In biological and social systems, individual specialisation enhances performance by allowing agents to excel at specific tasks (Woolley et al., 2015; Smith et al., 2008; Surowiecki, 2004), while generalist (homogeneous) behaviours promote adaptability and efficiency (Kassen, 2002; Williams & O’Reilly III, 1998). In Multi-Agent Reinforcement Learning (MARL), tasks similarly vary and require adaptability (Li et al., 2021; Bettini et al., 2024; Albrecht et al., 2024): some demand specialised agent behaviours, others benefit from homogeneous strategies, and some may require a blend of both (mixed behaviours).
Thus optimal performance is tied to being able to represent the required behaviours. No parameter sharing (NoPS) (Lowe et al., 2017) enables specialisation by using separate policy and critic networks per agent, but introduces significant computational overhead and sample inefficiency (Christianos et al., 2021). Conversely, full parameter sharing (FuPS) (Tan, 1993; Gupta et al., 2017; Foerster et al., 2016) tackles these inefficiencies by training a single shared network, but struggles in scenarios requiring diverse or specialised behaviours (Kim & Sung, 2023; Fu et al., 2022; Li et al., 2021).
Balancing FuPS efficiency with the capacity for diverse behaviours therefore remains a central open problem in MARL. Prior works have explored intrinsic-rewards (Li et al., 2021; Jiang & Lu, 2021), role-based allocations (Wang et al., 2020a, b), specialised architectures (Kim & Sung, 2023; Li et al., 2024; Bettini et al., 2024), or sequential updates (Zhong et al., 2024). However, these methods often alter the learning objective, require prior knowledge of optimal diversity levels, involve delicate hyperparameter tuning, or necessitate maintaining agent-specific parameters and sequential updates (see Section 7).
These challenges lead us to ask: Can we design a shared MARL architecture that flexibly supports both specialised and homogeneous behaviours—without modifying the learning objective or relying on prior knowledge of the optimal diversity level? A key difficulty with FuPS (particularly for diverse behaviours) is the gradient interference among agents, whereby their updates negatively impact each other’s learning (Christianos et al., 2021; Zhong et al., 2024).
To address this, we propose HyperMARL, a novel parameter sharing method that leverages hypernetworks (Ha et al., 2016) to generate agent-specific actor and critic parameters conditioned on agent IDs or embeddings (Figure 1). Hypernetworks are neural networks that generate the weights of another network conditioned on an input context vector. Crucially, they have already proven effective at resolving conflicting gradients in multi-task RL (Navon et al., 2020) and mitigating catastrophic forgetting in continual learning (von Oswald et al., 2020). In HyperMARL, they adaptively generate specialised or homogeneous policies and decouple observation- from agent-conditioned gradients (Section 4.4), a mechanism that empirically reduces gradient variance and improves policy updates (Sarafian et al., 2021). We further show that HyperMARL attains lower policy gradient variance than FuPS and that this decoupling is critical for specialisation (Section 6), suggesting it helps mitigate cross-agent gradient interference in FuPS.
We validate HyperMARL on diverse MARL benchmarks—including Dispersion and Navigation from the Vectorized Multi-Agent Simulator (VMAS) (Bettini et al., 2022), Multi-Agent MuJoCo (MAMuJoCo) (Peng et al., 2021), and SMAX (Rutherford et al., 2024)—across environments with two to twenty agents that require homogeneous, heterogeneous, or mixed behaviours. HyperMARL consistently matches or outperforms NoPS, FuPS, and diversity-promoting methods such as DiCo (Bettini et al., 2024) and HAPPO (Zhong et al., 2024). Furthermore, HyperMARL achieves NoPS-level behavioural diversity while using a shared architecture – a challenge for conventional FuPS methods.
Our contributions are as follows:
  • We propose HyperMARL (Section 4), a method that leverages agent-conditioned hypernetworks to adaptively learn diverse, homogeneous or mixed behaviours, without altering the learning objective or requiring preset diversity levels.
  • We empirically evaluate HyperMARL (Section 5) across diverse MARL environments with up to 20 agents. HyperMARL consistently achieves competitive mean episode returns compared to NoPS, FuPS, and diversity-focused methods such as DiCo and HAPPO, while achieving NoPS-level behavioural diversity.
  • We demonstrate that HyperMARL empirically reduces policy gradient variance compared to FuPS and facilitates specialisation, suggesting the importance of gradient decoupling for mitigating cross-agent interference (Section 5.2; Section 6). Additionally, HyperMARL scales efficiently by maintaining a near-constant number of parameters as the number of agents increases (Section C.2).
2 Background
We formulate the fully cooperative multi-agent systems addressed in our work as a Dec-POMDP (Oliehoek & Amato, 2016). A Dec-POMDP is a tuple, $\langle\mathbb{I},\mathbb{S},\{\mathbb{A}^{i}\}_{i\in\mathbb{I}},R,\{\mathbb{O}^{i}\}_{i\in\mathbb{I}},O,T,\rho_{0},\gamma\rangle$, where $\mathbb{I}$ is the set of agents of size $n=|\mathbb{I}|$, $\mathbb{S}$ is the set of global states with an initial state distribution $\rho_{0}$, $\mathbb{A}^{i}$ is the action space for agent $i$ where $\mathbb{A}=\times_{i}\mathbb{A}^{i}$ is the joint action space, $R:\mathbb{S}\times\mathbb{A}\to\mathbb{R}$ is the shared reward function, $\mathbb{O}^{i}$ is the observation space for agent $i$ with the joint observation space $\mathbb{O}=\times_{i}\mathbb{O}^{i}$, $O:\mathbb{O}\times\mathbb{A}\times\mathbb{S}\to[0,1]$ is the probability of joint observation $\mathbf{o}\in\mathbb{O}$, i.e. $O(\mathbf{o},\mathbf{a},s)=\Pr(\mathbf{o_{t}}|s_{t},\mathbf{a_{t-1}})$, $T:\mathbb{S}\times\mathbb{A}\times\mathbb{S}\to[0,1]$ is the state transition function i.e. $T\left(s,\mathbf{a},s^{\prime}\right)=\Pr\left(s_{t+1}\mid s_{t},\mathbf{a}_{t}\right)$ and $\gamma$ is the discount factor.
In this setting, each agent $i$ receives a partial observation $o_{t}^{i}\in\mathbb{O}^{i}$. These observations are accumulated into an action-observation history $h_{t}^{i}=\left(o_{0}^{i},a_{0}^{i},\ldots,o_{t-1}^{i},a_{t-1}^{i},o_{t}^{i}\right)$. Each agent $i$ acts based on their decentralized policies $\pi^{i}(a^{i}|h^{i})$. The joint history and joint action are defined as follows $\mathbf{h_{t}}=(h_{t}^{1},\ldots,h_{t}^{n})$ and $\mathbf{a_{t}}=(a_{t}^{1},\ldots,a_{t}^{n})$. The goal is to learn an optimal joint policy $\boldsymbol{\pi^{*}}=(\pi^{1*},...,\pi^{n*})$ that maximizes the expected discounted return as follows [^2] We use simplified notion here omitting explicit dependence on state transitions and distributions for brevity.,
$\boldsymbol{\pi}^{*}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{s_{0}\sim\rho_{0},\;\mathbf{h}\sim\boldsymbol{\pi}}\left[G(\mathbf{h})\right]$ (1)
where $G(\mathbf{h})=\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})$.
The action-value and state-value functions of the joint policy $\boldsymbol{\pi}$ can be defined as:
$V(\mathbf{h})=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})\,\Big{|}\,\mathbf{h}\right],$ (2)
$Q(\mathbf{h},\mathbf{a})=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})\,\Big{|}\,\mathbf{h},\,\mathbf{a}_{0}=\mathbf{a}\right].$ (3)
2.1 Specialised Policies and Environments
Specialisation plays a key role in MARL, yet remains under-defined, so we define specialised environments and specialised policies.
Definition 2.1.
An environment is specialised if:
  1. 1.
    Distinct Agent Policies. The optimal joint policy $\boldsymbol{\pi}^{*}$ consists of at least two distinct agent policies, i.e., $\exists i,j\in\mathbb{I}$ such that $\pi^{i}\neq\pi^{j}$.
  2. 2.
    Non-Interchangeability. Any permutation $\sigma$ of the policies in $\boldsymbol{\pi}^{*}$, denoted as $\boldsymbol{\pi}^{\sigma}$, results in a weakly lower expected return:
    $\mathbb{E}_{\mathbf{h}\sim\boldsymbol{\pi}^{\sigma}}[G(\mathbf{h})]\leq\mathbb{E}_{\mathbf{h}\sim\boldsymbol{\pi}^{*}}[G(\mathbf{h})],$
    with strict inequality if the joint policies are non-symmetric (i.e., swapping any individual policy degrades performance).
For example, consider a specialised environment such as a football game, agents adopt distinct roles — such as ”attackers” and ”defenders” — to optimise the team’s performance. Permuting these roles (i.e., exchanging their policies) would typically lead to suboptimal results. Here, agents develop specialised policies by learning distinct, complementary behaviours essential for an optimal joint policy. Heterogeneous agents — with different action or observation spaces — are naturally specialised, while homogeneous agents can specialise by learning distinct policies critical to the team’s success. We analyse such specialised environments in Sections 3.1 and 5.2.
3 Are Independent or Fully Shared Policies Sufficient?
In this section, we introduce two simple matrix games and show that even in these basic settings, both fully independent and fully parameter shared policies have inherent limitations, highlighting the need for more nuanced approaches.
3.1 Specialisation and Synchronisation Game
Specialisation Game
Player 2
A B
Player 1 A (0.5, 0.5) (1, 1)
B (1, 1) (0.5, 0.5)
Table 1: 2-player specialisation game. Nash Equilibria in blue.
The Specialisation Game, inspired by the XOR game (Fu et al., 2022) and VMAS’s Dispersion (Bettini et al., 2022), is a two-player matrix game where each agent selects one of two food items. Payoffs are $0.5$ for matching choices and $1$ for different choices (payoff matrix in Table 1).
This game has two symmetric pure Nash equilibria on the anti-diagonal of the payoff matrix and can be extended to $n$ agents, with payoffs of 1 for unique actions and $1/n$ for matching actions. This satisfies the conditions of a Specialised Environment (Definition 2.1).
Synchronisation Game This is the inverse of the Specialisation Game, where players aim to choose the same food item. Payoffs are 1 for matching choices and 0.5 for different choices, resulting in pure Nash equilibria along the diagonal. For $n$ agents, matching actions receive a payoff of $1$, while differing choices receive $1/n$.
3.2 Limitations of Fully Shared and Independent Policies
Table 2: Final evaluation rewards (mean $\pm$ std) for REINFORCE variants in Specialisation and Synchronisation Games with 2, 4, and 8 agents. Bold values are within 1 std of the maximum. Fully shared or independent policies do not solve both games.
Specialisation Game
Method 2 4 8
NoPS 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
FuPS 0.50 $\pm$ 0.00 0.25 $\pm$ 0.00 0.13 $\pm$ 0.00
FuPS+ID 1.00 $\pm$ 0.00 0.95 $\pm$ 0.10 0.53 $\pm$ 0.09
Synchronisation Game
Method 2 4 8
NoPS 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.13 $\pm$ 0.00
FuPS 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
FuPS+ID 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.30 $\pm$ 0.35
In Appendix B (Proof 14) we show that a fully shared policy (without conditioning on agent IDs) cannot learn the optimal joint behaviour in the 2-player Specialisation Game. To assess this limitation in practice, we compare three REINFORCE (Williams, 1992) variants in both games with $n=2,4,8$ agents: NoPS (No Parameter Sharing), FuPS (Fully Parameter Sharing), and FuPS+ID (FuPS with one-hot agent IDs). All policies use single-layer neural networks with controlled parameter counts (see Appendix F for details).
Table 2 shows that NoPS consistently achieves optimal performance in the Specialisation Game, while FuPS performance declines as the number of agents increases. Conversely, in the Synchronisation Game, FuPS performs optimally, whereas NoPS fails with $n=8$. FuPS+ID initially performs well in both games but deteriorates at larger scales, indicating that ID conditioning alone is insufficient for fully distinct or uniform behaviours.
Overall, these experiments demonstrate that neither fully shared nor fully independent policies provide a universal solution—even in simple matrix games. This highlights the need for more flexible MARL strategies that handle diverse task demands. In Section 5, we show similar findings for modern MARL algorithms and more sophisticated environments.
4 HyperMARL
In this section, we introduce HyperMARL, an approach that uses agent-conditioned hypernetworks to learn diverse or homogeneous policies end-to-end, without modifying the standard RL objective or requiring prior knowledge of the task’s optimal diversity. By operating within a fully shared paradigm, HyperMARL leverages shared gradient information while enabling specialisation through the decoupling of observation- and agent-conditioned gradients. We present the pseudocode in Section C.1.
4.1 Hypernetworks for MARL
Hypernetworks (Ha et al., 2016) are neural networks that generate the weights of another network conditioned on an input context vector. In HyperMARL, we use them to adaptively produce per-agent policy and critic parameters.
As illustrated in Figure 1, for any agent $i$ with context $e^{i}$ (i.e., either a one-hot encoded ID or a learned embedding), the hypernetworks generate the agent-specific parameters:
$\theta^{i}=h_{\psi}^{\pi}(e^{i}),\quad\phi^{i}=h_{\varphi}^{V}(e^{i}),$ (4)
where $h_{\psi}^{\pi}$ and $h_{\varphi}^{V}$ are the hypernetworks for the policy and critic, respectively. The parameters $\theta^{i}$ and $\phi^{i}$ define each agent’s policy $\pi_{\theta^{i}}$ and critic $V_{\phi^{i}}$, dynamically enabling either specialised or homogeneous behaviours as required by the task.
We explore two types of hypernetworks in HyperMARL: Linear and Multilayer Perceptron (MLP) Hypernetworks, balancing expressiveness and computational efficiency.
Linear Hypernetworks Linear hypernetworks are a straightforward instantiation of HyperMARL. Given a one-hot agent ID, $\mathds{1}^{i}\in\mathbb{R}^{1\times n}$, a linear hypernetwork $h_{\psi}^{\pi}$ generates agent-specific parameters $\theta^{i}$ as  [^3] For conciseness we only show the policy parameters in this section.:
$\theta^{i}=h_{\psi}^{\pi}(\mathds{1}^{i})=\mathds{1}^{i}\cdot W+b$ (5)
where $W\in\mathbb{R}^{n\times m}$ is the weight matrix (with $m$ the per-agent parameter dimensionality and $n$ is the number of agents), and $b\in\mathbb{R}^{1\times m}$ is the bias vector. Since $\mathds{1}^{i}$ is one-hot encoded, each $\theta^{i}$ corresponds to a specific row of $W$ plus the shared bias $b$. If there is no shared bias term, this effectively replicates training of separate policies for each task (in our case, for each agent) (Beck et al., 2023), since there are no shared parameters and gradient updates are independent.
MLP Hypernetworks for Expressiveness While linear hypernetworks can be effective (Beck et al., 2023; Chang et al., 2020), they may lack the capacity to capture complex relationships between agents’ roles and their corresponding weights. To enhance expressiveness, we incorporate hidden layers and non-linear activations (e.g., ReLU) into the hypernetworks as follows:
$\theta^{i}=h_{\psi}^{\pi}(e^{i})=f_{\psi_{1}}^{\pi}\bigl{(}g_{\psi_{2}}^{\pi}(e^{i})\bigr{)}$ (6)
where $e^{i}$ is the agent ID or embedding for agent $i$, $g_{\psi_{2}}^{\pi}$ is the multilayer perceptrons (MLPs) that produce intermediate representations with non-linear activations, and $f_{\psi_{1}}^{\pi}$ and is the final linear transformations that generate the policy parameters.
This additional expressiveness may be critical in domains where roles overlap or where nuanced specialisations emerge dynamically. Unlike linear hypernetworks with one-hot agent IDs, MLP hypernetworks do not guarantee distinct weights for each agent. Additionally, they increase the total number of trainable parameters, necessitating a careful balance between expressiveness and computational overhead.
4.2 Agent Embeddings
The agent embedding $e^{i}$ represents the input to the hypernetwork. For our linear hypernetworks, we use one-hot encoded agent IDs as embeddings, resulting in a mix of agent-independent weights and shared weights as discussed in Section 4.1. For our MLP hypernetworks, we use learned agent embeddings, initialised orthogonally. These embeddings are optimised end-to-end alongside the hypernetwork, allowing agents to learn distinct or similar embeddings based on whether homogeneous or specialised policies are required.
4.3 Hypernetwork Initialisation
HyperMARL initialises its hypernetworks to ensure that the initial agent-specific parameters $\theta^{i}$ and $\phi^{i}$ adhere to the same distribution as standard networks without hypernetworks. For example, PPO commonly uses orthogonal initialisation (Engstrom et al., 2019), so HyperMARL generates orthogonally initialised weights for each agent, preserving the same fan in/out as the conventional network. This alignment with standard deep RL initialisation schemes facilitates stable learning and can be applied using any initialisation strategy.
4.4 Gradient Decoupling in HyperMARL
A core difficulty in FuPS is cross-agent gradient interference (Christianos et al., 2021; Zhong et al., 2024). In HyperMARL, gradient decoupling helps mitigate this issue.
Consider a fully cooperative MARL setting with a centralised critic, we can formulate the policy gradient for agent $i$ as follows (Albrecht et al., 2024; Kuba et al., 2021):
$\nabla_{\theta^{i}}J(\theta^{i})\;=\;\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\Bigl{[}A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\Bigr{]},$
where $\mathbf{h_{t}}$ and $\mathbf{a_{t}}$ are the joint histories and joint actions for all agents, $\theta^{i}$ denotes the parameters of agent $i$, and $A(\mathbf{h_{t}},\mathbf{a_{t}})=Q(\mathbf{h_{t}},\mathbf{a_{t}})-V(\mathbf{h_{t}})$ is the advantage function.
When using shared parameters $\theta$, this can reframed as:
$\nabla_{\theta}J(\theta)=\sum_{i=1}^{I}\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\left[A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\theta}\log\pi_{\theta}\left(a_{t}^{i}\mid h_{t}^{i},id\right)\right],$ (7)
where $id$ is our agent ID.
In HyperMARL, each agent’s policy parameters $\theta^{i}$ are generated by a hypernetwork $h_{\psi}^{\pi}$ conditioned on an agent embedding $e^{i}$:
$\theta^{i}\;=\;h_{\psi}^{\pi}\bigl{(}e^{i}\bigr{)}.$
Hence, the gradient with respect to $\psi$ becomes:
$\nabla_{\psi}J(\psi)\;=\;\sum_{i=1}^{I}\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\Bigl{[}A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\psi}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\Bigr{]}.$
Applying the chain rule:
$\nabla_{\psi}\,\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\;=\;\underbrace{\nabla_{\theta^{i}}\,\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})}_{\text{Observation-Conditioned}}\;\cdot\;\underbrace{\nabla_{\psi}\,h_{\psi}^{\pi}(e^{i})}_{\text{Agent-Conditioned}},$
leading to
$\nabla_{\psi}J(\psi)$ $=\sum_{i=1}^{I}\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\Bigl{[}A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})$
$\quad\cdot\nabla_{\psi}h_{\psi}^{\pi}(e^{i})\Bigr{]}.$ (8)
This factorisation separates:
Observation-Conditioned Gradients ($\nabla_{\theta^{i}}\!\log\pi_{\theta^{i}}$), which reflect how each policy updates in response to local observations.
Agent-Conditioned Gradients ($\nabla_{\psi}\,h_{\psi}^{\pi}(e^{i})$), which reflect how the hypernetwork generates each agent’s parameters from each agent’s embedding $e^{i}$.
These gradients thus remain partially decoupled, reducing cross-agent interference. Related work in Meta-RL shows that similar decoupling lowers policy gradient variance and stabilises learning (Sarafian et al., 2021). In Section 5.2, we observe a comparable reduction in variance with HyperMARL. Additionally, ablation studies in Section 6 show that removing gradient decoupling significantly degrades performance, highlighting its essential role in HyperMARL’s effectiveness.
5 Experiments
We evaluate HyperMARL on two primary questions:
Q1: Specialised Policy Learning: Can HyperMARL effectively learn specialised policies (see Section 2.1) using a shared hypernetwork, without altering the learning objective or relying on preset diversity levels?
Q2: Effectiveness in Homogeneous Tasks: Is HyperMARL competitive in environments that necessitate homogeneous behaviours?
5.1 Experimental Setup
Table 3: Summary of MARL environments used to evaluate HyperMARL.
Environment Agents Action Space Behaviour
Dispersion 4 Discrete Heterogeneous
Navigation 2, 4, 8 Continuous Homogeneous, Heterogeneous, Mixed
MAMuJoCo 2–17 Continuous Heterogeneous
SMAX 2–20 Discrete Homogeneous
Environments We use diverse cooperative MARL tasks (Table 3), varying in complexity, agent count (2 to 20), and required behaviours (homogeneous, heterogeneous, or mixed). Full details are in Appendix D.1.
MARL Baselines We focus on PPO variants, given their success in MARL (De Witt et al., 2020; Yu et al., 2022). Comparisons include IPPO (De Witt et al., 2020), which learns an actor and critic conditioned on individual observations, and MAPPO (Yu et al., 2022), which trains a centralised critic conditioned on the global state or concatenated agent observations. We use NoPS and FuPS variants of these methods, with all FuPS variants being conditioned on one-hot encoded agent IDs.
(a) IPPO
(b) MAPPO
Figure 2: Performance of IPPO and MAPPO on Dispersion after 20 million timesteps. We show the Interquartile Mean (IQM) of the Mean Episode Return and the 95% Stratified Bootstrap Confidence Intervals (CI) using (Agarwal et al., 2021). Hypernetworks achieve comparable performance to NoPS, while FuPS struggle with specialisation. Sample efficiency plots are provided in Appendix E.1.1.
(a) Policy Diversity using SND with Jenson-Shannon distance. Hypernetworks match NoPS-level diversity while sharing parameters. Bold indicates highest SND, asterisks show overlaps in std.
Method Algorithm Mean ± Std Dev
IPPO MLP Hypernet $6.3\times 10^{-4}\pm 2.0\times 10^{-5}$
Linear Hypernet $6.4\times 10^{-4}\pm 7.0\times 10^{-6}$
NoPS $\mathbf{6.9\times 10^{-4}\pm 7.0\times 10^{-6}}$
FuPS $2.4\times 10^{-4}\pm 1.1\times 10^{-4}$
MAPPO MLP Hypernet $\mathbf{6.7\times 10^{-4}\pm 5.0\times 10^{-6}}$
Linear Hypernet $6.6\times 10^{-4}\pm 1.2\times 10^{-5*}$
NoPS $6.6\times 10^{-4}\pm 1.5\times 10^{-5*}$
FuPS $3.7\times 10^{-4}\pm 8.1\times 10^{-5}$
(b) Average policy gradient variance across IPPO (left) and MAPPO (right). Our hypernetworks have lower variance than FuPS.
Privileged Baselines Beyond NoPS, we include two privileged baselines designed for heterogeneous policies. Diversity Control (DiCo) (Bettini et al., 2024) partitions each policy into shared and agent-specific components and enforces a target diversity level (via Wasserstein distances). Heterogeneous-Agent PPO (HAPPO) (Zhong et al., 2024) uses a shared centralised critic and updates individual actors sequentially, incorporating prior actor updates to allow agent heterogeneity. In contrast, HyperMARL uses fully shared weights, parallel updates, and no preset diversity levels. Table 7 summarises baseline selection criteria. Implementation and hyperparameters follow authors’ defaults unless otherwise stated (see Appendix F), and HyperMARL uses the same observations and generates architectures with equivalent capacity to the baselines. Training and evaluation protocols are consistent across all experiments (Appendix D.1.1).
5.1.1 Measuring Policy Diversity
To measure the diversity of the policies we System Neural Diversity (SND) (Bettini et al., 2023) (Equation 9) with Jensen-Shannon distance, where 0 indicates identical policies and 1 indicates maximal diversity. Details in Appendix D.1.2.
5.2 Q1: Specialised Policy Learning
Learning Diverse Behaviour (Dispersion) Figures 2(a) and 2(b) show that parameter sharing variants (IPPO-FuPS, MAPPO-FuPS – ($\bullet$)) fail to learn the diverse policies required by Dispersion, while their non-parameter sharing counterparts (IPPO-NoPS, MAPPO-NoPS–($\bullet$)) converge to the optimal policy. These results corroborate that parameter sharing methods struggle in environments requiring agent specialisation. In contrast, our shared linear and MLP hypernetworks ($\bullet$, $\bullet$) match NoPS performance, suggesting that a shared hypernetwork can effectively support agent specialisation.
Policy Diversity SND measurements (Table 3(a)) show that IPPO-FuPS and MAPPO-FuPS exhibit lower behavioural diversity than their NoPS counterparts, aligning with their weaker performance results. Notably, both linear and MLP hypernetworks achieve diversity levels comparable to NoPS despite using a shared architecture. Linear hypernetworks with one-hot agent IDs enforce weight seperation, while MLP hypernetworks have to dynamically learn diverse policies without explicit separation of weights or gradients.
Gradient Variance To investigate HyperMARL’s effect on training stability, we compute the average policy gradient variance by calculating the variance across all actor parameters at each update and averaging it over training. As shown in Figure 3(b), hypernetworks in both IPPO and MAPPO exhibit lower mean policy gradient variance than FuPS. This aligns with their ability to learn diverse behaviours and may stem from decoupling observation- and agent-condition gradients (Section 4.4), similar to Meta-RL findings (Sarafian et al., 2021).
Table 4: Mean episode return in MAMuJoCo for MAPPO variants. IQM of the mean episode returns with 95% stratified bootstrap CI. Bold denotes the highest score per scenario; * indicates overlap with the highest score within one CI. HyperMARL (MLP variant) is competitive with methods that use independent actors.
Scenario HAPPO FuPS Ind. Actors HyperMARL
Humanoid-v2 17x1 6501.15* (3015.88, 7229.79) 566.12 (536.36, 603.01) 6188.46* (5006.13, 6851.74) 6544.10 (3868.00, 6664.89)
Walker2d-v2 2x3 4748.06* (4366.94, 6230.81) 4574.39* (4254.21, 5068.32) 4747.05* (3974.76, 6249.58) 5064.86 (4635.10, 5423.42)
HalfCheetah-v2 2x3 6752.40* (6130.42, 7172.98) 6771.21* (6424.94, 7228.65) 6650.31* (5714.68, 7229.61) 7063.72 (6696.30, 7325.36)
Ant-v2 4x2 6031.92* (5924.32, 6149.22) 6148.58 (5988.63, 6223.88) 6046.23* (5924.62, 6216.57) 5940.16* (5485.77, 6280.59)
(a) All Agents with Same Goals
(b) All Agents with Different Goals
(c) Mixed Goals (Half Agents Share Goals)
Figure 4: Average Reward in Navigation for IPPO Variants. IQM and 95% CI of mean reward for IPPO Baselines, HyperMARL (linear and MLP), and DiCo in the Navigation environment across goal configurations. HyperMARL maintains strong IQM performance, particularly with more agents. Sample efficiency plots in Appendix E.3.
Diversity at Complexity and Scale (MAMuJoCo) Dispersion is designed to test diverse behaviours (Bettini et al., 2022), but we further evaluate HyperMARL on the more challenging heterogeneous MAMuJoCo environment. Following Zhong et al. (2024), we use a centralised but non-shared critic, applying HyperMARL only to the actor.
Table 4 shows that HyperMARL (MLP variant) is competitive with all baselines across various MAMuJoCo scenarios. Unlike HAPPO and MAPPO (independent actors), HyperMARL uses a shared actor and parallel updates, and yet manages strong performance. This is particularly notable in Humanoid-v2 17x1(Figure 5), where 17 agents to coordinate in a Humanoid control task, a notoriously difficult heterogeneous task(Zhong et al., 2024). This demonstrates HyperMARL’s ability to handle large-scale specialisation with a shared architecture.
Figure 5: Performance on the 17-agent Humanoid task. IQM of the mean episode returns with 95% stratified bootstrap CI. HyperMARL is able to match the performance of methods that leverage independent actors, while using a shared actor in this challenging heterogeneous task.
Adaptability (Navigation) We evaluate HyperMARL against DiCo (Bettini et al., 2024) in the Navigation environment (Bettini et al., 2022). For two agents ($n=2$), we adopt DiCo’s optimal diversity levels and replicate its hyperparameters. For larger teams ($n>2$), we perform hyperparameter sweeps for diversity levels and learning parameters across all methods (see Tables 14 and 15). Although DiCo defines optimal diversity for two agents, identifying suitable levels for larger team sizes proved challenging.
Figure 4 presents the IQM of the mean reward across different goal setups. HyperMARL consistently demonstrates higher average returns across all configurations, outperforming DiCo and matching or exceeding NoPS and FuPS performance for most settings. Unlike in Dispersion, IPPO-FuPS is competitive with NoPS and HyperMARL in the Navigation environment for tasks requiring diverse behaviours (mixed or different goals – Figures 4(b) and 4(c)) with two and four agents. We hypothesise that this arises from the dense rewards in Navigation, in contrast to the sparse rewards in Dispersion. However, when scaling to eight agents, HyperMARL outperforms all baselines, demonstrating its effectiveness in handling more complex coordination challenges.
Figure 6: IQM and 95% CI of mean return in SMAX. Performance of FuPS Recurrent IPPO and MAPPO, HyperMARL (MLP) on SMAX. HyperMARL performs comparably to these baselines.
5.3 Q2: Effectiveness in Homogeneous Tasks
SMAX  Finally, we test HyperMARL (MLP variant) on SMAX, comparing against the FuPS baseline from JaxMARL. Prior work (Yu et al., 2022; Fu et al., 2022) has found FuPS optimal in such environments. Figure 6 shows that across simple (2s3z), hard (3s5z), and SMACv2-style maps (smacv2_10, smacv2_20), HyperMARL is comparable to the FuPS baseline. Alongside results in Navigation (Figure 4(a)), this indicates that HyperMARL is also effective in tasks requiring homogeneous behaviours, large observation spaces, and coordination among many agents, reinforcing its effectiveness across diverse MARL scenarios.
6 Ablations
To better understand the importance of the different components in HyperMARL, we conduct an ablation study on four key variations of HyperMARL (MAPPO-MLP) in the Humanoid-v2 17x1 (Section 5.2). First, HyperMARL-S reduces the hypernetwork hidden layer size from 64 to 32, investigating the impact of network capacity. Second, HyperMARL w/ One-Hot replaces learned agent embeddings with one-hot IDs (Section 4.2). Third, HyperMARL w/o RF omits the initialisation scaling described in Section 4.3. Finally, HyperMARL w/o GD removes gradient decoupling (Section  4.4) by conditioning the hypernetwork on both observations and agent embeddings.
Figure 7 shows that HyperMARL-S converges more slowly but ultimately reaches an IQM close to the baseline HyperMARL, while learned agent embeddings offer better final performance than one-hot IDs (HyperMARL w/ One-Hot). Crucially, removing either gradient decoupling (HyperMARL w/o GD) or initialisation scaling (HyperMARL w/o RF) severely impairs learning, confirming their essential roles in HyperMARL ability to learn specialised policies. We see similar findings in ablations in Dispersion (Section 18).
Figure 7: Ablation results comparing HyperMARL and its variants in Humanoid-v2 17x1. Gradient decoupling and initialization are crucial for maintaining performance.
7 Related Work
Hypernetworks in RL and MARL. Hypernetworks have proven effective in meta-learning, multi-task learning, and continual learning in single-agent settings (Beck et al., 2023, 2024; Sarafian et al., 2021; Huang et al., 2021). In MARL, QMIX (Rashid et al., 2020) used hypernetworks (conditioned on a global state) to mix per-agent Q-values; however, each agent’s own network remained a standard MLP or GRU. Parallel work, CASH (Fu et al., 2025), conditions hypernetworks on local observations and team capabilities for zero-shot generalization with heterogeneous action spaces. In contrast, we focus on agent-conditioned hypernetworks for adaptive specialisation within a fixed state-action setting, leveraging gradient decoupling (absent in CASH) that we find critical for specialised behaviours.
Variants of Parameter Sharing. Selective Parameter Sharing (SePS) (Christianos et al., 2021) shares weights between similar groups of agents, identified via trajectory clustering. Pruning methods (Kim & Sung, 2023; Li et al., 2024) split a single network into agent-specific subnetworks.
Learning Diverse Policies. Shared parameters often limit policy diversity (Christianos et al., 2021; Kim & Sung, 2023; Fu et al., 2022; Li et al., 2021). Proposed solutions include: (1) maximizing mutual information between agent IDs and trajectories (Li et al., 2021), (2) role-based methods (Wang et al., 2020a, b), or (3) methods that use structural modifications or constraints to induce diversity in agent policies (Kim & Sung, 2023; Bettini et al., 2024; Li et al., 2024). Outside FuPS/NoPS, HAPPO (Zhong et al., 2024) uses a non-shared centralised critic with individual actors updated sequentially to learn heterogeneous behaviours.
8 Conclusion and Discussion
We introduced HyperMARL, an approach that uses agent-conditioned hypernetworks to generate per-agent actor and critic parameters without modifying the standard RL objective or requiring preset diversity levels. Our results show it can adaptively learn specialised, homogeneous, or mixed behaviours in settings with up to 20 agents. We also observe a link between HyperMARL’s performance and reduced policy gradient variance, underscoring the importance of decoupling observation- and agent-conditioned gradients. Overall, these findings establish HyperMARL as a versatile and promising architecture for diverse MARL tasks. Future work could explore more parameter-efficient hypernetworks (e.g., chunked variants (von Oswald et al., 2020; Chauhan et al., 2024)) to enable large-scale applications.
Impact Statement
This paper presents work whose goal is to advance the field of Multi-Agent Reinforcement Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Measuring Behavioural Diversity
A.1 Quantifying Team Diversity
We quantify policy diversity using System Neural Diversity (SND) (Bettini et al., 2023), which measures behavioural diversity based on differences in policy outputs:
$\operatorname{SND}\left(\left\{\pi^{i}\right\}_{i\in\mathbb{I}}\right)=\frac{2}{n(n-1)|\mathcal{O}|}\sum_{i=1}^{n}\sum_{j=i+1}^{n}\sum_{o\in\mathcal{O}}D\left(\pi^{i}(o),\pi^{j}(o)\right).$ (9)
where $n$ is the number of agents, $\mathcal{O}$ is a set of observations typically collected via policy rollouts, $\pi^{i}(o_{t})$ and $\pi^{j}(o_{t})$ are the outputs of policies $i$ and $j$ for observation $o_{t}$, and $D$ is our distance function between two probability distributions.
In contrast to  (Bettini et al., 2023), we use Jensen-Shannon Distance (JSD) (Endres & Schindelin, 2003; Lin, 1991) as $D$, rather than the Wasserstein metric (Vaserstein, 1969). As shown in Appendix A.2, JSD is a robust metric for both continuous and discrete cases, and provides a more reliable measure of policy distance.
A.2 Finding a Suitable Distance Function for Policy Diversity
The choice of distance function $D$ in Equation 9 is crucial for accurately measuring policy diversity. In MARL, policies are often represented as probability distributions over actions, making the choice of distance function non-trivial.
(Bettini et al., 2024) use the Wasserstein metric for continuous policies (Vaserstein, 1969) as distance function $D$, while (McKee et al., 2022) use the total variation distance for discrete policies. For discrete policies, Wasserstein distance would require a cost function representing the cost of changing from one action to another, which might not be trivial to come up with. On the other hand, although well-suited for discrete policies, TVD might miss changes in action probabilities because it measures the largest difference assigned to an event (i.e. action) between two probability distributions.
We consider a simple example to illustrate this point. Suppose we have two policies $\pi^{1}$ and $\pi^{2}$ with action probabilities as shown in Figure 8. $\pi^{1}$ stays constant, while $\pi^{2}$ changes gradually over timesteps. We see that even as $\pi^{2}$ changes over time, the $TVD(\pi^{1},\pi^{2})$ between $\pi^{1}$ and $\pi^{2}$ remains constant. This is because TVD only measures the largest difference between the two distributions, and does not consider the overall difference between them. On the other hand, the Jensen-Shannon distance (JSD) (Endres & Schindelin, 2003), which is the square root of the Jensen-Shannon divergence, does not have this problem as it is a smooth distance function. Furthermore, it satisfies the conditions for being a metric – it is non-negative, symmetry, and it satisfies the triangle inequality.
For continuous policies, as shown in Figure 9, JSD exhibits similar trends to the Wasserstein distance. Since JSD is a reasonable metric for both continuous and discrete probability distributions, we will use it as the distance metric for all experiments and propose it as a suitable distance function for measuring policy diversity in MARL.
We also summarise the properties of the various distance metrics in Table 5.
Figure 8: Gradual changes in $\pi^{2}$, result in gradual changes in the Jensen-Shannon distance (JSD), while the Total Variation Distance (TVD) can miss changes in action probabilities.
Figure 9: Jensen-Shannon distance (JSD) trends similarly to Wasserstein distance when we have continuous policies.
Method Kinds of Actions Metric Smooth Formula
[…]
Total Variation Distance Discrete Metric No $TV(p,q)=\frac{1}{2}\sum_{x}|p(x)-q(x)|$
Jensen-Shannon Divergence (Lin, 1991) Both Divergence Yes $JSD(p\parallel q)=\frac{1}{2}D_{KL}(p\parallel m)+\frac{1}{2}D_{KL}(q\parallel m),\,m=\frac{1}{2}(p+q)$
Jensen-Shannon Distance (Endres & Schindelin, 2003) Both Metric Yes $\sqrt{JSD(p\parallel q)}$
Table 5: Measure Policy Diversity
Appendix B Specialisation and Synchronisation Game
Theorem B.1.
A shared policy cannot learn the optimal behaviour for the two-player Specialisation Game.
Proof.
Let $\pi$ be a shared policy for both agents, and let $\alpha=\mathbb{P}(a_{i}=0)$ represent the probability of any agent choosing action 0. The expected return of $\pi$ for each agent is:
$E[R(\pi)]$ $=\mathbb{P}(a_{0}=0,a_{1}=0)\cdot 0.5+\mathbb{P}(a_{0}=0,a_{1}=1)\cdot 1$ (10)
$\quad+\mathbb{P}(a_{0}=1,a_{1}=0)\cdot 1+\mathbb{P}(a_{0}=1,a_{1}=1)\cdot 0.5$ (11)
$=0.5\alpha^{2}+2\alpha(1-\alpha)+0.5(1-\alpha)^{2}$ (12)
$=-\alpha^{2}+\alpha+0.5$ (13)
$=-(\alpha-0.5)^{2}+0.75$ (14)
Thus, $E[R(\pi)]\leq 0.75<1$ for all $\alpha\in[0,1]$, with the maximum at $\alpha=0.5$. Therefore, a shared policy cannot achieve the optimal return of 1, confirming the need for specialised behaviour to optimise rewards. ∎
In Table 6, we present the results for the Specialisation and Synchronisation Games when actions are sampled during evaluation rather than selecting the argmax, as was done in Table 2.
Table 6: Average Evaluation Reward at the end of training (mean $\pm$ std) for REINFORCE variants in the Specialisation and Synchronisation Games, across 2, 4, and 8 Agents. Bold indicates values within 1 std of the maximum. Actions are sampled during evaluation.
Specialisation Game
Method 2 4 8
NoPS .9996 $\pm$ .0004 .9988 $\pm$ .0007 .9989 $\pm$ .0008
FuPS .7493 $\pm$ .0062 .6826 $\pm$ .0039 .6545 $\pm$ .0025
FuPS+ID .$9$$\pm$ .0002 .9995 $\pm$ .0003 .7275 $\pm$ .0293
Synchronisation Game
Method 2 4 8
NoPS .9993 $\pm$ .0004 .9976 $\pm$ .0022 .1250 $\pm$ 0
FuPS .$9$$\pm$ .0002 1.0 $\pm$ 0 .3000 $\pm$ .3500
FuPS+ID .$9$$\pm$ 0 .$9$$\pm$ .0003 .1250 $\pm$ 0
Appendix C HyperMARL Details
C.1 HyperMARL Pseudocode
In Algorithm 1, we present the pseudocode for HyperMARL, with HyperMARL-specific steps highlighted in blue. HyperMARL leverages hypernetworks to dynamically generate the parameters of both actor and critic networks. The weights of the hypernetworks and the agent embeddings are automatically updated through automatic differentiation (autograd) based on the computed loss. Additionally, Figure 1 provides a visual representation of the HyperMARL architecture.
Algorithm 1 HyperMARL
1:  Input: Number of agents $n$, number of training iterations $K$, MARL algorithm parameters (e.g., MAPPO-specific hyperparameters)
2:  Initialize:
3:   Hypernetwork parameters $\psi,\varphi$Ensure $\theta^{i}$ and $\phi^{i}$ follow standard initialization schemes, e.g., orthogonal}
4:   Agent embeddings $\{e^{i}\}_{i=1}^{n}$One-hot or orthogonally initialized learnable parameters}
5:  Output: Optimized joint policy $\boldsymbol{\pi}$
6:  for each training iteration $k=0,1,\dots,K-1$ do
7:     for each agent $i=1,\dots,n$ do
8:        $\theta^{i}\leftarrow h_{\psi}^{\pi}(e^{i})$Policy parameters}
9:        $\phi^{i}\leftarrow h_{\varphi}^{V}(e^{i})$Critic parameters}
10:     end for
11:     Interact with environment using $\{\pi_{{\color[rgb]{0.2734375,0.51171875,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.51171875,0.70703125}\theta^{i}}}\}_{i=1}^{n}$ to collect trajectories $\mathcal{D}$
12:     Compute shared loss $\mathcal{L}$ from $\mathcal{D}$, using $\{V_{{\color[rgb]{0.2734375,0.51171875,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.51171875,0.70703125}\phi^{i}}}\}_{i=1}^{n}$Standard RL loss function}
13:     Update $\psi$, $\varphi$, and $\boldsymbol{e}$ by minimizing $\mathcal{L}$Optimise parameters.}
14:  end for
15:  Return $\boldsymbol{\pi}=(\pi^{1},\dots,\pi^{n})$
C.2 Scalability and Parameter Efficiency
Hypernetworks generate weights for the target network, which can lead to high-dimensional outputs and many parameters for deep target networks. This challenge is amplified in MLP-based hypernetworks, which include additional hidden layers. Figure 10 shows scaling trends:
  • NoPS and linear hypernetworks: Parameter count grows linearly with the number of agents.
  • FuPS: More efficient, as growth depends on one-hot vector size.
  • MLP hypernetworks: Scale better with larger populations, since they only require embeddings of fixed size for each new agent.
To reduce parameter count, techniques like shared hypernetworks, chunked hypernetworks (von Oswald et al., 2020; Chauhan et al., 2024), or producing low-rank weight approximations, can be used. While naive implementations are parameter-intensive, this might be less critical in RL and MARL which commonly have smaller actor-critic networks. Moreover, HyperMARL’s near-constant scaling with agents suggests strong potential for large-scale MARL applications.
Figure 10: Parameter scaling for IPPO variants with increasing agents (4 to 1024). MLP Hypernetworks scale nearly constantly, while NoPS, Linear Hypernetworks, and FuPS+One-Hot grow linearly. Log scale on both axes.
Figure 11: Dispersion performance with four agents. FuPS variants match HyperMARL in parameter count but still underperform.
To isolate the effects of parameter count, we scaled the FuPS networks (Figure 11) to match the number of trainable parameters in HyperMARL. Despite generating 10x smaller networks, HyperMARL consistently outperforms FuPS variants, showing its advantages extend beyond parameter count.
Appendix D Experiment Details
D.1 Environments
(a) Dispersion.
(b) Navigation.
(c) MAMuJoCo.
(d) SMAX.
Figure 12: Multi-Agent environments used in our experiments.
Dispersion (VMAS) (Bettini et al., 2022): A 2D environment where four agents collect unique food particles. This task requires specialised heterogeneous behaviours and resembles the Specialisation Game from Section 3.1.
Navigation (VMAS) (Bettini et al., 2022): Agents navigate in a 2D space to assigned goals, receiving dense rewards based on proximity. Configurations include shared goals (homogeneous), unique goals (heterogeneous), and mixed goals, where some agents share goals while others have unique ones.
Multi-Agent MuJoCo (MAMuJoCo) (Peng et al., 2021): A multi-agent extension of MuJoCo, where robot body parts (e.g., a cheetah’s legs) are modelled as different agents. Agents coordinate to perform efficient motion, requiring heterogeneous policies (Zhong et al., 2024).
SMAX (JaxMARL) (Rutherford et al., 2024): Discrete action tasks with 2 to 20 agents on SMACv1- and SMACv2-style maps. FuPS baselines have been shown optimal for these settings (Yu et al., 2022; Fu et al., 2022) indicating homogeneous behaviour is preferred here.
Table 7: Overview of Baseline Methods and Justifications. Selected baselines (✓) are chosen based on their relevance to our focus on parameter sharing and specialisation/generalisation in MARL. Non-selected methods (✗) are excluded due to implementation limitations or misalignment with our scope.
Baseline Method Category Chosen Justification
IPPO (NoPS, FuPS) (De Witt et al., 2020) NoPS/FuPS Serves as a strong MARL baseline for both independent (NoPS) and shared (FuPS) policies.
MAPPO (NoPS, FuPS) (Yu et al., 2022) NoPS/FuPS Provides a robust MARL baseline with centralized critics for both NoPS and FuPS configurations.
DiCo (Bettini et al., 2024) Architectural Enables comparison with a method that employs preset diversity levels and maintains both non-shared and shared parameters.
HAPPO (Zhong et al., 2024) Sequential Updates Facilitates comparison with a method designed for heterogeneous behaviours, utilizing sequential updates and maintaining non-shared parameters per agent.
SEAC (Christianos et al., 2020) Shared Experience Focuses on experience sharing rather than parameter sharing, which is outside our scope.
SePS (Christianos et al., 2021) SePS Requires pretraining, which is beyond the scope of our current study.
CDAS (Li et al., 2021) Intrinsic Reward Only implemented for off-policy methods which are outside our scope, and has been shown to underperform FuPS/NoPS (Fu et al., 2022).
ROMA/RODE (Wang et al., 2020b, a) Role-based Demonstrates limited practical performance (Christianos et al., 2020).
SNP-PS (Kim & Sung, 2023) Architectural/Pruning No available code implementation.
Kaleidoscope (Li et al., 2024) Architectural/Pruning Only implemented for off-policy methods, which is outside our scope.
D.1.1 Training and Evaluation
  • Training:
    • For Dispersion (5.2), we run 5 seeds and train for 20 million timesteps.
    • For Navigation (5.2), SMAX (5.3), and MaMuJoCo (5.2), we run 5 seeds and train for 10 million timesteps, consistent with the baselines.
  • Evaluation:
    • For Dispersion (5.2), evaluation is performed every 100k timesteps across 32 episodes.
    • For Navigation (5.2), following the baselines, evaluation is performed every 120k timesteps across 200 episodes.
    • For SMAX (5.3), evaluation is performed every 500k timesteps across 32 episodes.
    • For MaMuJoCo (5.2), following the baselines, evaluation is performed every 25 training episodes over 40 episodes.
D.1.2 Measuring Policy Diversity Details
We measure team diversity using the System Neural Diversity (SND) metric (Equation 9 (Bettini et al., 2023), details Section A) with Jensen-Shannon distance. SND ranges from 0 (identical policies across all agents) to 1 (maximum diversity). We collect a dataset of observations from IPPO-NoPS and IPPO-FuPS policies checkpointed at 5 and 20 million training steps. Each policy is rolled out for 10,000 episodes, generating 16 million observations. We then sample 1 million observations from this dataset to calculate the SND for each method tested.
D.2 HyperMARL Architecture
For the experiments in Sections 5.2 and 5.2, we use feedforward architectures, while for Section 5.3, we use recurrent GRU networks (Cho, 2014). For recurrent PPO variants, HyperMARL maintains the same architecture as for feedforward methods, generating only the actor and critic feedforward weights, not the GRU weights.
Appendix E Detailed Results
E.1 Dispersion Detailed Results
E.1.1 Sample Efficiency Plots
(a) IPPO Sample Efficiency
(b) MAPPO Sample Efficiency
Figure 13: Sample efficiency of IPPO and MAPPO variants on the Dispersion environment. While FuPS methods struggle to converge to optimal performance, Hypernetworks achieve efficiency comparable to NoPS. Even when running MAPPO-FuPS for longer (see Figure 14), it converges to suboptimal performance.
Figure 14: We see that even if we run MAPPO-FuPS on Dispersion for 40 million timesteps (double the timesteps of MLP Hypernetwork), it converges to suboptimal performance.
E.2 Detailed MAMujoco Plots
(a) Humanoid-v2 17x1
(b) Walker2d-v2 2x3
(c) HalfCheetah-v2 2x3
(d) Ant-v2 4x2
Figure 15: Performance of Recurrent IPPO and MAPPO on SMAX. HyperMARL performs comparably to these baselines.
E.3 Detailed Navigation Plots
(a) Same Goal (8 agents)
(b) Different Goals (8 agents)
(c) Four Goals (8 agents)
(d) Scenarios with 8 Agents
(e) Same Goal (4 agents)
(f) Different Goals (4 agents)
(g) Two Goals (4 agents)
(h) Scenarios with 4 Agents
(i) Same Goal (2 agents)
(j) Different Goals (2 agents)
(k) Scenarios with 2 Agents
Figure 17: Sample Efficiency of IPPO variants in the VMAS Navigation environment under different goal configurations and agent numbers.
E.4 Additional Ablations
Figure 18: Ablation results comparing HyperMARL with its variants in Dispersion. The results highlight that gradient decoupling is essential for maintaining HyperMARL’s performance.
Appendix F Hyperparameters
Table 8: Hyperparameters for Specialisation and Synchronisation Game
Hyperparameter Value
Number of agents and foods 2, 4, 8
Number of foods 3
Number of seeds 5
Number of trials 10,000
Evaluation trials 1,000
Batch size 10
Learning rate 0.01
Activation function ReLU
Output activation Softmax
Optimizer SGD
Total Number of Parameters
   2 Agents
NoPS: 36
FuPS: 34
FuPS+ID: 42
   4 Agents
NoPS: 112
FuPS: 100
FuPS+ID: 148
   8 Agents
NoPS: 384
FuPS: 328
FuPS+ID: 552
Table 9: IPPO and MAPPO Hyperparameters in Dispersion
Hyperparameter Value
LR 0.0005
GAMMA 0.99
VF_COEF 0.5
CLIP_EPS 0.2
ENT_COEF 0.01
NUM_ENVS 16
NUM_STEPS 128
GAE_LAMBDA 0.95
NUM_UPDATES 9765
EVAL_EPISODES 32
EVAL_INTERVAL 100000
MAX_GRAD_NORM 0.5
UPDATE_EPOCHS 4
NUM_MINIBATCHES 2
TOTAL_TIMESTEPS 20000000
ANNEAL_LR false
ACTOR_LAYERS [64, 64]
CRITIC_LAYERS [64, 64]
ACTIVATION relu
SEEDS 30,1,42,72858,2300658
ACTION_SPACE_TYPE discrete
Table 10: MLP Hypernet Hyperparameters in Dispersion
Parameter IPPO MAPPO
HYPERNET_EMBEDDING_DIM 4 8
EMBEDDING_DIM Sweep [4, 16, 64] [4, 8, 16, 64]
HYPERNET_HIDDEN_DIMS 64 64
Table 11: Dispersion Settings
Setting Value
n_food 4
n_agents 4
max_steps 100
food_radius 0.08
share_reward false
penalise_by_time true
continuous_actions false
Table 12: IPPO Hyperparameters for Navigation
Hyperparameters Value
LR 0.00005
NUM_ENVS 600
NUM_STEPS 100
TOTAL_TIMESTEPS $10^{6}$
UPDATE_EPOCHS 45
NUM_MINIBATCHES 30
GAMMA 0.9
GAE_LAMBDA 0.9
CLIP_EPS 0.2
ENT_COEF 0.0
VF_COEF 1.0
MAX_GRAD_NORM 5
ACTIVATION tanh
ANNEAL_LR False
ACTOR_LAYERS [256, 256]
CRITIC_LAYERS [256, 256]
ACTION_SPACE_TYPE continuous
Table 13: MLP Hypernet Hyperparameters in Navigation
Parameter IPPO MAPPO
HYPERNET_EMBEDDING_DIM 4 8
EMBEDDING_DIM Sweep [4, 16, 64] [4, 8, 16, 64]
HYPERNET_HIDDEN_DIMS 64 64
Table 14: DiCo Algorithm $SND\_des$ Hyperparameter
Goal Configuration Number of Agents SND_des
All agents same goal 2 0
4 0
8 0
All agents different goals 2 1.2 (From DiCo paper)
4 [-1,1.2,2.4] $\Rightarrow$ -1 (Best)
8 [-1,1.2,4.8] $\Rightarrow$ -1 (Best)
Some agents share goals 4 [-1,1.2] $\Rightarrow$ -1 (Best)
8 [-1,2.4,1.2] $\Rightarrow$ -1 (Best)
Table 15: Parameter Sweeps for IPPO Variants in Navigation Tasks with Four and Eight Agents
Parameter Sweeps
CLIP_EPS 0.2, 0.1
LR 5e-5, 5e-4, 2.5e-4
Algorithm Setting Selected Values
IPPO-FuPS 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.2, 5e-5
4 Agents (Different Goals) 0.2, 5e-5
4 Agents (Two Goals) 0.2, 5e-5
IPPO-Linear Hypernetwork 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.2, 5e-5
4 Agents (Different Goals) 0.1, 5e-5
4 Agents (Two Goals) 0.1, 5e-5
IPPO-MLP Hypernetwork 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.1, 5e-5
4 Agents (Different Goals) 0.1, 5e-5
4 Agents (Two Goals) 0.1, 5e-5
IPPO-NoPS 8 Agents (Same Goals) 0.1, 5e-5
8 Agents (Different Goals) 0.2, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.1, 5e-5
4 Agents (Different Goals) 0.2, 5e-5
4 Agents (Two Goals) 0.1, 5e-5
IPPO-Dico 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 2.5e-4
8 Agents (Four Goals) 0.1, 2.5e-4
4 Agents (Same Goals) 0.2, 5e-5
4 Agents (Different Goals) 0.1, 2.5e-4
4 Agents (Two Goals) 0.1, 5e-4
Table 16: Environment Settings for Navigation Task
Parameter Value
n_agents 2,4,8
agents_with_same_goal 1, n_agents/2, n_agents
max_steps 100
collisions False
split_goals False
observe_all_goals True
shared_rew False
lidar_range 0.35
agent_radius 0.1
continuous_actions True
Table 17: Default algorithm and model hyperparameters for the Ant-v2-4x2 environment (from  (Zhong et al., 2024)).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 1
clip_param 0.1
critic_epoch 5
critic_num_mini_batch 1
entropy_coef 0
fixed_order true
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 5
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.0005
data_chunk_length 10
gain 0.01
hidden_sizes [128, 128, 128]
initialization_method orthogonal_
lr 0.0005
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 18: Default algorithm and model hyperparameters for the Humanoid-v2-17x1 environment (from  (Zhong et al., 2024)).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 1
clip_param 0.1
critic_epoch 5
critic_num_mini_batch 1
entropy_coef 0
fixed_order false
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 5
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.0005
data_chunk_length 10
gain 0.01
hidden_sizes [128, 128, 128]
initialization_method orthogonal_
lr 0.0005
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 19: Default algorithm and model hyperparameters for the Walker2d-v2-2x3 environment (from  (Zhong et al., 2024)).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 2
clip_param 0.05
critic_epoch 5
critic_num_mini_batch 2
entropy_coef 0
fixed_order false
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 5
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.001
data_chunk_length 10
gain 0.01
hidden_sizes 128, 128, 128
initialization_method orthogonal_
lr 0.001
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 20: Default algorithm and model hyperparameters for the HalfCheetah-v2-2x3 environment (from  (Zhong et al., 2024)).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 1
clip_param 0.05
critic_epoch 15
critic_num_mini_batch 1
entropy_coef 0.01
fixed_order false
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 15
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.0005
data_chunk_length 10
gain 0.01
hidden_sizes 128, 128, 128
initialization_method orthogonal_
lr 0.0005
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 21: HyperMARL Hyperparameters Across MaMuJoCo Environments
Parameter Humanoid-v2-17x1 Walker2d-v2-2x3 HalfCheetah-v2-2x3 Ant-v2-4x2 Sweeps
AGENT_ID_EMBEDDING_DIM 64 64 64 64 None
HYPERNET_HIDDEN_DIMS 64 64 64 64 None
clip_param 0.075 0.0375 0.0375 0.075 [0.1,0.075,0.05,0.0375]
Table 22: Recurrent IPPO and MAPPO Hyperparameters in SMAX (from JaxMARL paper)
Hyperparameter IPPO Value MAPPO Value
LR 0.004 0.002
NUM_ENVS 128 128
NUM_STEPS 128 128
GRU_HIDDEN_DIM 128 128
FC_DIM_SIZE 128 128
TOTAL_TIMESTEPS 1e7 1e7
UPDATE_EPOCHS 4 4
NUM_MINIBATCHES 4 4
GAMMA 0.99 0.99
GAE_LAMBDA 0.95 0.95
CLIP_EPS 0.05 0.2
SCALE_CLIP_EPS False False
ENT_COEF 0.01 0.0
VF_COEF 0.5 0.5
MAX_GRAD_NORM 0.25 0.25
ACTIVATION relu relu
SEED 30,1,42,72858,2300658 30,1,42,72858,2300658
ANNEAL_LR True True
OBS_WITH_AGENT_ID - True
Table 23: Hyperparameter Sweeps and Final Values for Different Maps in SMAX. H- refers to HyperMARL MLP Hypernetworks.
Map Algorithm LR Range Chosen LR HNET Embedding Dim HNET Hidden Dims
2s3z IPPO 0.004 0.004 -
MAPPO 0.002 0.002 - -
H-IPPO 0.004 0.004 4 32
H-MAPPO 0.002 0.002 64 16
3s5z IPPO 0.004 0.004 - -
MAPPO 0.002,0.005,0.0003 0.002 - -
H-IPPO 0.004 0.004 64 16
H-MAPPO 0.002,0.005,0.0003 0.0003 64 16
smacv2_10_units IPPO 0.005,0.001,0.0003,0.004 0.001 - -
MAPPO 0.002,0.005,0.0003 0.0003 - -
H-IPPO 0.005,0.001,0.0003,0.004 0.005 4 64
H-MAPPO 0.002,0.005,0.0003 0.0003 64 16
smacv2_20_units IPPO 0.002,0.005,0.0003 0.005 - -
MAPPO 0.002,0.005,0.0003 0.0003 - -
H-IPPO 0.002,0.005,0.0003 0.005 64 64
H-MAPPO 0.002,0.005,0.0003 0.0003 4 64
Note: HNET Embedding Dim refers to the hypernetwork embedding dimension, chosen from the range {4, 16, 64}. HNET Hidden Dims refers to the hidden layer dimensions of the hypernetwork, chosen from the range {16, 32, 64}.

### v3

@@
 Abstract
-Adaptability is critical in cooperative multi-agent reinforcement learning (MARL), where agents must learn specialised or homogeneous behaviours for diverse tasks.
-While parameter sharing methods are sample-efficient, they often encounter gradient interference among agents, limiting their behavioural diversity.
-Conversely, non-parameter sharing approaches enable specialisation, but are computationally demanding and sample-inefficient.
-To address these issues, we propose HyperMARL, a parameter sharing approach that uses hypernetworks to dynamically generate agent-specific actor and critic parameters, without altering the learning objective or requiring preset diversity levels.
-By decoupling observation- and agent-conditioned gradients, HyperMARL empirically reduces policy gradient variance and facilitates specialisation within FuPS, suggesting it can mitigate cross-agent interference.
-Across multiple MARL benchmarks involving up to twenty agents—and requiring homogeneous, heterogeneous, or mixed behaviours—HyperMARL consistently performs competitively with fully shared, non-parameter-sharing, and diversity-promoting baselines, all while preserving a behavioural diversity level comparable to non-parameter sharing.
-These findings establish hypernetworks as a versatile approach for MARL across diverse environments.
-[^1] All code will be made available soon.
-Multi-Agent Reinforcement Learning, Parameter Sharing, Specialisation, Behavioural Diversity
+Adaptability to specialised or homogeneous behaviours is critical in cooperative multi-agent reinforcement learning (MARL).
+Parameter sharing (PS) techniques, common for efficient adaptation, often limit behavioural diversity due to cross-agent gradient interference, which we show can be exacerbated by the coupling of observations and agent IDs.
+Current remedies typically add complexity through altered objectives, manual preset diversity levels, or sequential updates.
+We ask: can shared policies adapt without these complexities?
+We propose HyperMARL, a PS approach using hypernetworks for dynamic agent-specific parameters, without altering the RL objective or requiring preset diversity levels.
+HyperMARL’s explicit decoupling of observation- and agent-conditioned gradients empirically reduces policy gradient variance, facilitates shared-policy adaptation (including specialisation), and helps mitigate cross-agent interference.
+Across diverse MARL benchmarks (up to 20 agents), requiring homogeneous, heterogeneous, or mixed behaviours, HyperMARL achieves competitive performance against key baselines – fully shared, non-parameter sharing, and three diversity-promoting methods – while preserving behavioural diversity comparable to non-parameter sharing.
+These findings establish HyperMARL as a versatile approach for adaptive MARL.
+The code is publicly available at https://github.com/KaleabTessera/HyperMARL.
 1 Introduction
+Specialist and generalist behaviours are critical to collective intelligence, enhancing performance and adaptability in both natural and artificial systems [47, 38, 39, 22, 45].
+In Multi-Agent Reinforcement Learning (MARL), this translates to a critical need for policies that can flexibly exhibit specialised, homogeneous, or mixed behaviours to meet diverse task demands [25, 6, 2].
 Figure 1: HyperMARL Policy Architecture.
 Common Agent-ID conditioned shared MARL policy (left) vs HyperMARL (right), which uses a hypernetwork to generate agent-specific policies and decouples observation- and agent-conditioned gradients.
-Specialist and generalist behaviours are critical to collective intelligence.
-In biological and social systems, individual specialisation enhances performance by allowing agents to excel at specific tasks (Woolley et al., 2015; Smith et al., 2008; Surowiecki, 2004), while generalist (homogeneous) behaviours promote adaptability and efficiency (Kassen, 2002; Williams & O’Reilly III, 1998).
-In Multi-Agent Reinforcement Learning (MARL), tasks similarly vary and require adaptability (Li et al., 2021; Bettini et al., 2024; Albrecht et al., 2024): some demand specialised agent behaviours, others benefit from homogeneous strategies, and some may require a blend of both (mixed behaviours).
-Thus optimal performance is tied to being able to represent the required behaviours.
-No parameter sharing (NoPS) (Lowe et al., 2017) enables specialisation by using separate policy and critic networks per agent, but introduces significant computational overhead and sample inefficiency (Christianos et al., 2021).
-Conversely, full parameter sharing (FuPS) (Tan, 1993; Gupta et al., 2017; Foerster et al., 2016) tackles these inefficiencies by training a single shared network, but struggles in scenarios requiring diverse or specialised behaviours (Kim & Sung, 2023; Fu et al., 2022; Li et al., 2021).
+Optimal MARL performance thus hinges on being able to represent the required behaviours.
+While No Parameter Sharing (NoPS) [28] enables specialisation by using distinct per-agent networks, it suffers from significant computational overhead and sample inefficiency [10].
+Conversely, Full Parameter Sharing (FuPS) [40, 17, 14], which trains a single shared network, improves efficiency but typically struggles to foster the behavioural diversity necessary for many complex tasks [23, 16, 25].
 Balancing FuPS efficiency with the capacity for diverse behaviours therefore remains a central open problem in MARL.
-Prior works have explored intrinsic-rewards (Li et al., 2021; Jiang & Lu, 2021), role-based allocations (Wang et al., 2020a, b), specialised architectures (Kim & Sung, 2023; Li et al., 2024; Bettini et al., 2024), or sequential updates (Zhong et al., 2024).
+Prior works have explored intrinsic-rewards [25, 21], role-based allocations [43, 44], specialised architectures [23, 26, 6], or sequential updates [49].
 However, these methods often alter the learning objective, require prior knowledge of optimal diversity levels, involve delicate hyperparameter tuning, or necessitate maintaining agent-specific parameters and sequential updates (see Section 7).
-These challenges lead us to ask: Can we design a shared MARL architecture that flexibly supports both specialised and homogeneous behaviours—without modifying the learning objective or relying on prior knowledge of the optimal diversity level?
-A key difficulty with FuPS (particularly for diverse behaviours) is the gradient interference among agents, whereby their updates negatively impact each other’s learning (Christianos et al., 2021; Zhong et al., 2024).
-To address this, we propose HyperMARL, a novel parameter sharing method that leverages hypernetworks (Ha et al., 2016) to generate agent-specific actor and critic parameters conditioned on agent IDs or embeddings (Figure 1).
-Hypernetworks are neural networks that generate the weights of another network conditioned on an input context vector.
-Crucially, they have already proven effective at resolving conflicting gradients in multi-task RL (Navon et al., 2020) and mitigating catastrophic forgetting in continual learning (von Oswald et al., 2020).
-In HyperMARL, they adaptively generate specialised or homogeneous policies and decouple observation- from agent-conditioned gradients (Section 4.4), a mechanism that empirically reduces gradient variance and improves policy updates (Sarafian et al., 2021).
-We further show that HyperMARL attains lower policy gradient variance than FuPS and that this decoupling is critical for specialisation (Section 6), suggesting it helps mitigate cross-agent gradient interference in FuPS.
-We validate HyperMARL on diverse MARL benchmarks—including Dispersion and Navigation from the Vectorized Multi-Agent Simulator (VMAS) (Bettini et al., 2022), Multi-Agent MuJoCo (MAMuJoCo) (Peng et al., 2021), and SMAX (Rutherford et al., 2024)—across environments with two to twenty agents that require homogeneous, heterogeneous, or mixed behaviours.
-HyperMARL consistently matches or outperforms NoPS, FuPS, and diversity-promoting methods such as DiCo (Bettini et al., 2024) and HAPPO (Zhong et al., 2024).
-Furthermore, HyperMARL achieves NoPS-level behavioural diversity while using a shared architecture – a challenge for conventional FuPS methods.
+These challenges lead us to ask: Can we design a shared MARL architecture that flexibly supports both specialised and homogeneous behaviours—without altered learning objectives, manual preset diversity levels, or sequential updates?
+A key difficulty with FuPS, particularly for diverse behaviours, was hypothesised to be gradient interference among agents, whereby their updates negatively impact each other’s learning [10, 49].
+We not only empirically validate this hypothesis but also demonstrate a critical insight: this conflict is significantly exacerbated by the common practice of coupling observations with agent IDs within a shared network (Fig.
+1 for coupling, Sec.
+3.2 for results).
+To counteract the challenges posed by this coupling, we propose HyperMARL, a novel parameter sharing method that explicitly decouples observation- and agent-conditioned gradients (Section 4.3) through hypernetworks [18].
+Specifically, agent-conditioned hypernetworks, which condition on an agent identifier (or embedding), and generate an agent’s actor and critic weights on the fly (Fig.
+1).
+Hypernetworks have already proved effective for resolving gradient conflicts in multi-task RL [30] and for mitigating catastrophic forgetting in continual learning [42], our work establishes their effectiveness in MARL.
+We further show that HyperMARL empirically attains lower policy gradient variance than FuPS and that this decoupling is critical for specialisation (Sec.
+5.26), suggesting it helps mitigate cross-agent gradient interference in shared architectures.
+We validate HyperMARL on diverse MARL benchmarks—including Dispersion and Navigation (VMAS) [5], Multi-Agent MuJoCo (MAMuJoCo) [34], and SMAX [36]—across environments with two to twenty agents that require homogeneous, heterogeneous, or mixed behaviours.
+HyperMARL consistently matches or outperforms NoPS, FuPS, and diversity-promoting methods such as DiCo [6],HAPPO [49] and Kaleidoscope [26], while achieving NoPS-level behavioural diversity while using a shared architecture.
 Our contributions are as follows:
-  • We propose HyperMARL (Section 4), a method that leverages agent-conditioned hypernetworks to adaptively learn diverse, homogeneous or mixed behaviours, without altering the learning objective or requiring preset diversity levels.
-  • We empirically evaluate HyperMARL (Section 5) across diverse MARL environments with up to 20 agents.
-HyperMARL consistently achieves competitive mean episode returns compared to NoPS, FuPS, and diversity-focused methods such as DiCo and HAPPO, while achieving NoPS-level behavioural diversity.
+  • Motivated by evidence that coupling agent IDs with observations can be detrimental, we propose HyperMARL (Sec.
+4), a method that leverages agent-conditioned hypernetworks to decouple observation- and agent-conditioned gradients.
+HyperMARL adaptively learns diverse, homogeneous, or mixed behaviours without altering the RL objective or requiring preset diversity levels.
+  • We thoroughly evaluate HyperMARL (Section 5) across diverse MARL environments with up to 20 agents.
+HyperMARL consistently achieves competitive mean episode returns compared to NoPS, FuPS, and diversity-focused methods such as DiCo, HAPPO and Kaleidoscope, while achieving NoPS-level behavioural diversity.
   • We demonstrate that HyperMARL empirically reduces policy gradient variance compared to FuPS and facilitates specialisation, suggesting the importance of gradient decoupling for mitigating cross-agent interference (Section 5.2; Section 6).
-Additionally, HyperMARL scales efficiently by maintaining a near-constant number of parameters as the number of agents increases (Section C.2).
 2 Background
-We formulate the fully cooperative multi-agent systems addressed in our work as a Dec-POMDP (Oliehoek & Amato, 2016).
+We formulate the fully cooperative multi-agent systems addressed in our work as a Dec-POMDP [31].
 A Dec-POMDP is a tuple, $\langle\mathbb{I},\mathbb{S},\{\mathbb{A}^{i}\}_{i\in\mathbb{I}},R,\{\mathbb{O}^{i}\}_{i\in\mathbb{I}},O,T,\rho_{0},\gamma\rangle$, where $\mathbb{I}$ is the set of agents of size $n=|\mathbb{I}|$, $\mathbb{S}$ is the set of global states with an initial state distribution $\rho_{0}$, $\mathbb{A}^{i}$ is the action space for agent $i$ where $\mathbb{A}=\times_{i}\mathbb{A}^{i}$ is the joint action space, $R:\mathbb{S}\times\mathbb{A}\to\mathbb{R}$ is the shared reward function, $\mathbb{O}^{i}$ is the observation space for agent $i$ with the joint observation space $\mathbb{O}=\times_{i}\mathbb{O}^{i}$, $O:\mathbb{O}\times\mathbb{A}\times\mathbb{S}\to[0,1]$ is the probability of joint observation $\mathbf{o}\in\mathbb{O}$, i.e.
@@
 The joint history and joint action are defined as follows $\mathbf{h_{t}}=(h_{t}^{1},\ldots,h_{t}^{n})$ and $\mathbf{a_{t}}=(a_{t}^{1},\ldots,a_{t}^{n})$.
-The goal is to learn an optimal joint policy $\boldsymbol{\pi^{*}}=(\pi^{1*},...,\pi^{n*})$ that maximizes the expected discounted return as follows [^2] We use simplified notion here omitting explicit dependence on state transitions and distributions for brevity.,
-$\boldsymbol{\pi}^{*}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{s_{0}\sim\rho_{0},\;\mathbf{h}\sim\boldsymbol{\pi}}\left[G(\mathbf{h})\right]$ (1)
-where $G(\mathbf{h})=\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})$.
-The action-value and state-value functions of the joint policy $\boldsymbol{\pi}$ can be defined as:
-$V(\mathbf{h})=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})\,\Big{|}\,\mathbf{h}\right],$ (2)
-$Q(\mathbf{h},\mathbf{a})=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})\,\Big{|}\,\mathbf{h},\,\mathbf{a}_{0}=\mathbf{a}\right].$ (3)
-2.1 Specialised Policies and Environments
+The goal is to learn an optimal joint policy $\boldsymbol{\pi^{*}}=(\pi^{1*},...,\pi^{n*})$ that maximizes the expected discounted return as follows, [^1] We use simplified notion here omitting explicit dependence on state transitions and distributions for brevity.
+$\boldsymbol{\pi}^{*}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{s_{0}\sim\rho_{0},\;\mathbf{h}\sim\boldsymbol{\pi}}\left[G(\mathbf{h})\right]$, where $G(\mathbf{h})=\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})$.
+Specialised Policies and Environments.
+We say an environment is specialised if its optimal joint policy contains at least two distinct, non-interchangeable agent policies ( Def.
+1 in App.
+C).
+Under this mild condition, tasks such as Dispersion (5.2) or our Specialisation Game (E.1) require agents to learn complementary roles rather than identical behaviours.
+3 Are Independent or Fully Shared Policies Enough?
+Standard independent (NoPS) and fully parameter-shared (FuPS) policies face inherent trade-offs in MARL.
+NoPS allows for uninhibited agent specialisation but can be sample inefficient and computationally expensive.
+FuPS, often conditioned with an agent ID (FuPS+ID), is more efficient but can struggle when agents must learn diverse behaviours.
+This section investigates the limitations of these common policy architectures.
+P2
+A B
+P1 A (0.5,0.5) (1,1)
+B (1,1) (0.5,0.5)
+Two-Player Payoff matrix
+$N$-player Interaction
+(a) Specialisation Game
+P2
+A B
+P1 A (1,1) (0.5,0.5)
+B (0.5,0.5) (1,1)
+Two-Player Payoff matrix
+$N$-player Interaction
+(b) Synchronisation Game
+Figure 2: Specialisation and Synchronisation Games.
+The Specialisation game (left), which encourages distinct actions, and the Synchronisation game (right), where rewards encourage identical actions.
+Depicted are their two-player payoff matrices (pure Nash equilibria in blue) and $N$-player interaction schematics.
+While simple in form, these games are challenging MARL benchmarks due to non-stationarity and exponentially scaling observation spaces (temporal version).
+To probe these limitations, we introduce two illustrative environments: the Specialisation Game, rewarding distinct actions, and the Synchronisation Game, rewarding identical actions.
+Both are inspired by prior work [16, 5, 32] and extended here to $N$-agent and temporal settings where agents observe prior joint actions (see Appendix E for full definitions).
+3.1 Limitations of Fully Shared and Independent Policies
+Figure 3: Average evaluation reward (mean $\pm$ 95% CI) for temporal Specialisation vs.
+Synchronisation using REINFORCE (10 seeds).
+Bold: highest mean, no CI overlap.
+Neither fully shared nor independent policies consistently achieve the highest mean reward.
+Specialisation Synchronisation
+#Ag NoPS FuPS FuPS+ID NoPS FuPS FuPS+ID
+2 0.88$\pm$0.09 0.50$\pm$0.00 0.64$\pm$0.10 0.83$\pm$0.12 1.00$\pm$0.00 0.91$\pm$0.09
+4 0.74$\pm$0.08 0.25$\pm$0.00 0.40$\pm$0.07 0.32$\pm$0.03 1.00$\pm$0.00 0.67$\pm$0.15
+8 0.68$\pm$0.02 0.12$\pm$0.00 0.25$\pm$0.03 0.14$\pm$0.00 1.00$\pm$0.00 0.54$\pm$0.10
+16 0.64$\pm$0.01 0.06$\pm$0.00 0.13$\pm$0.02 0.07$\pm$0.00 1.00$\pm$0.00 0.55$\pm$0.14
+(a) Avg. evaluation reward
+(b) Avg. gradient conflict
+Figure 4: Multi-agent policy gradient methods in the Specialisation environment.
+The FuPS+ID (No State) ablation outperforms FuPS+ID, showing near-orthogonal gradients (4(b)), indicating that observation–ID decoupling is important.
+HyperMARL (MLP) enables this decoupling while leveraging state information, and achieves better performance and reduced gradient conflict than FuPS+ID.
+FuPS without agent IDs provably cannot recover optimal pure Nash equilibria in the non-temporal 2-player Specialisation Game (Proof E.3, App.
+E).
+In practice, however, FuPS is often conditioned with agent IDs, and MARL policies must handle complexities beyond static, two-player interactions.
+We therefore evaluate standard architectures in the temporal $n$-player versions of these games [^2] Results for non-temporal (normal-form) variants are in App.
+E.5.
+We compare three standard architectures trained with REINFORCE [46]: 1) NoPS: independent policies ($\pi_{\theta^{i}}(a^{i}|o^{i})$); 2) FuPS: a single shared policy ($\pi_{\theta}(a^{i}|o^{i})$); and 3) FuPS+ID: a shared policy incorporating a one-hot agent ID ($\pi_{\theta}(a^{i}|o^{i},id^{i})$).
+All use single-layer networks, 10-step episodes, and $10,000$ training steps (further details in App.
+6).
+Empirical Performance.
+Table 3 shows that neither NoPS nor FuPS consistently achieves the highest mean evaluation rewards.
+NoPS excels in the Specialisation Game but is outperformed by FuPS (optimal) and FuPS+ID in the Synchronisation Game.
+Furthermore, the performance gaps widen as the number of agents increases (notably at $n=8$ and $n=16$), highlighting the scalability challenges of both fully independent and fully shared policies.
+3.2 Why FuPS+ID Fails to Specialise: The Problem of Gradient Conflict
+Despite being a universal approximator [19], FuPS+ID often struggles to learn diverse policies in practice (Table 3, [10, 49]).
+A key reason is gradient conflict: when a single network processes both observation $o$ and agent ID $id^{i}$, updates intended to specialise agent $i$ (based on $id^{i}$) can conflict with updates for agent $j$ (based on $id^{j}$), particularly if they share similar observations but require different actions.
+This obstructs the emergence of specialised behaviours (conflict measured via inter-agent gradient cosine similarity, App.
+E.4).
+Importance of Observation and ID Decoupling To investigate the effect of entangled observation and ID inputs, we introduce an ablation: FuPS+ID (No State), where the policy $\pi_{\theta}(a^{i}\mid id^{i})$ conditions only on the agent ID, ignoring observations.
+Surprisingly, FuPS+ID (No State) outperforms standard FuPS+ID in the Specialisation Game for all tested $N$ (Figure 4(a)), even when $N\leq 4$ (where observation spaces are small, suggesting the issue is not merely observation size).
+Figure 4(b) reveals why: FuPS+ID (No State) shows near-zero gradient conflict (nearly orthogonal gradients), whereas standard FuPS+ID exhibits negative cosine similarities (conflicting gradients).
+These results show that naively coupling observation and ID inputs in shared networks can lead to destructive interference, hindering specialisation.
+While discarding observations is not a general solution (most tasks require state information), this finding motivates designing architectures that can leverage both state and agent IDs, while minimising interference.
+Section 4 introduces HyperMARL (Figure 1), which explicitly decouples observation- and agent-conditioned gradients through agent-conditioned hypernetworks, leading to improved performance over FuPS variants and reduced gradient conflict compared to standard FuPS+ID (Figure 4).
+4 HyperMARL
+We introduce HyperMARL, an approach that uses agent-conditioned hypernetworks to learn diverse or homogeneous policies end-to-end, without modifying the standard RL objective or requiring manual preset diversity levels.
+By operating within a fully shared paradigm, HyperMARL leverages shared gradient information while enabling specialisation through the decoupling of observation- and agent-conditioned gradients.
+We present the pseudocode in Sec.
+F.1, with additional scaling (F.3) and runtime (F.4) details.
+4.1 Hypernetworks for MARL
+As illustrated in Figure 1, for any agent $i$ with context $e^{i}$ (i.e., either a one-hot encoded ID or a learned embedding), the hypernetworks generate the agent-specific parameters:
+$\theta^{i}=h_{\psi}^{\pi}(e^{i}),\quad\phi^{i}=h_{\varphi}^{V}(e^{i}),$ (1)
+where $h_{\psi}^{\pi}$ and $h_{\varphi}^{V}$ are the hypernetworks for the policy and critic, respectively.
+The parameters $\theta^{i}$ and $\phi^{i}$ define each agent’s policy $\pi_{\theta^{i}}$ and critic $V_{\phi^{i}}$, dynamically enabling either specialised or homogeneous behaviours as required by the task.
+Linear Hypernetworks Given a one-hot agent ID, $\mathds{1}^{i}\in\mathbb{R}^{1\times n}$, a linear hypernetwork $h_{\psi}^{\pi}$ generates agent-specific parameters $\theta^{i}$ as :
+$\theta^{i}=h_{\psi}^{\pi}(\mathds{1}^{i})=\mathds{1}^{i}\cdot W+b$ (2)
+where $W\in\mathbb{R}^{n\times m}$ is the weight matrix (with $m$ the per-agent parameter dimensionality and $n$ is the number of agents), and $b\in\mathbb{R}^{1\times m}$ is the bias vector.
+Since $\mathds{1}^{i}$ is one-hot encoded, each $\theta^{i}$ corresponds to a specific row of $W$ plus the shared bias $b$.
+If there is no shared bias term, this effectively replicates training of separate policies for each task (in our case, for each agent) [3], since there are no shared parameters and gradient updates are independent.
+MLP Hypernetworks for Expressiveness To enhance expressiveness, MLP Hypernetworks incorporate hidden layers and non-linearities:
+$\theta^{i}=h_{\psi}^{\pi}(e^{i})=f_{\psi_{1}}^{\pi}\bigl{(}g_{\psi_{2}}^{\pi}(e^{i})\bigr{)}$ (3)
+where $g_{\psi_{2}}^{\pi}$ is an MLP processing the agent context $e^{i}$, and $f_{\psi_{1}}^{\pi}$ is a final linear output layer.
+Unlike linear hypernetworks with one-hot agent IDs, MLP hypernetworks do not guarantee distinct weights for each agent.
+Additionally, they increase the total number of trainable parameters, necessitating a careful balance between expressiveness and computational overhead.
+4.2 Agent Embeddings and Initialisation
+The agent embedding $e^{i}$ is a one-hot encoded ID for Linear Hypernetworks.
+For MLP Hypernetworks, we use learned agent embeddings, orthogonally initialised and optimised end-to-end with the hypernetwork.
+HyperMARL’s hypernetworks are themselves initialised such that the generated agent-specific parameters ($\theta^{i},\phi^{i}$) initially match the distribution of standard direct initialisation schemes (e.g., orthogonal for PPO, preserving fan in/out), promoting stable learning.
+4.3 Gradient Decoupling in HyperMARL
+A core difficulty in FuPS is cross-agent gradient interference [10, 49].
+HyperMARL mitigates this by generating each agent’s parameters through a shared hypernetwork, thereby decoupling agent-conditioned and observation-conditioned components of the gradient.
+Hypernetwork gradients. Consider a fully cooperative MARL setting with a centralised critic, we can formulate the policy gradient for agent $i$ as follows [2, 24]:
+$\nabla_{\theta^{i}}J(\theta^{i})\;=\;\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\Bigl{[}A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\Bigr{]},$
+where $\mathbf{h_{t}}$ and $\mathbf{a_{t}}$ are the joint histories and joint actions for all agents, $\theta^{i}$ denotes the parameters of agent $i$, and $A(\mathbf{h_{t}},\mathbf{a_{t}})=Q(\mathbf{h_{t}},\mathbf{a_{t}})-V(\mathbf{h_{t}})$ is the advantage function.
+Decoupling.
+In HyperMARL each agent’s policy weights are produced by the hypernetwork $h_{\psi}^{\pi}$: $\theta^{i}=h_{\psi}^{\pi}(e^{i})$, so we optimise a single parameter vector $\psi$.
+Applying the chain rule and re-ordering the expectations:
+$\nabla_{\!\psi}J(\psi)=\sum_{i=1}^{I}\underbrace{\nabla_{\!\psi}h_{\psi}^{\pi}(e^{i})}_{\mathbf{J}_{i}\;\text{(agent-conditioned)}}\;\underbrace{\mathbb{E}_{\mathbf{h}_{t},\mathbf{a}_{t}\sim\boldsymbol{\pi}}\bigl{[}A(\mathbf{h}_{t},\mathbf{a}_{t})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\bigr{]}}_{Z_{i}\;\text{(observation-conditioned)}}.$ (4)
+  • Agent-conditioned factor $\mathbf{J}_{i}$.
+This Jacobian depends only on the fixed embedding $e^{i}$ and the hypernetwork weights $\psi$, therefore, it is deterministic with respect to mini-batch samples (as $e^{i}$ and $psi$ are fixed per gradient step), separating agent identity from trajectory noise.
+  • Observation-conditioned factor $Z_{i}$.
+The expectation averages trajectory noise per agent $i$ for its policy component $\pi_{\theta^{i}}$, prior to transformation by $\mathbf{J}_{i}$ and aggregation.
+The crucial structural decoupling in Equation (4) ensures HyperMARL first averages noise per agent (via factor $Z_{i}$) before applying the deterministic agent-conditioned transformation $\mathbf{J}_{i}$.
+This mitigates gradient interference common in FuPS+ID  [10, 49], where observation noise and agent identity become entangled (see Equation (12) in App.
+F.2).
+This is the MARL analogue of the task/state decomposition studied by [37, Eq.
+18] in Meta-RL.
+Section 5.2 empirically verifies the predicted variance drop, and ablations in Section 6 demonstrate that disabling decoupling degrades performance, underscoring its critical role.
+5 Experiments
+Our empirical evaluation of HyperMARL assesses whether agent-conditioned hypernetworks can enable adaptive (specialised or homogeneous) policies without altered RL objectives, preset diversity levels, or sequential updates.
+We structure our experiments to directly answer two key research questions: Q1: Specialised Policy Learning: Can HyperMARL effectively learn specialised policies via a shared hypernetwork?
+Q2: Effectiveness in Homogeneous Tasks: Is HyperMARL competitive in environments that necessitate homogeneous behaviours?
+To address these questions, HyperMARL is evaluated against five representative modern baselines across a carefully selected suite of MARL benchmarks (18 scenarios, up to 20 agents).
+All experiments use at least 5 seeds (details in App.
+G.2.1).
+5.1 Experimental Setup
+Table 1: MARL environments for evaluating HyperMARL.
+Env. Agents Action Behaviour
+Dispersion 4 Discrete Hetero.
+Navigation 2,4,8 Continuous Homo., Hetero., Mixed
+MAMuJoCo 2–17 Continuous Hetero.
+SMAX 2–20 Discrete Homo.
+Environments.
+HyperMARL is evaluated across 18 scenarios from four diverse MARL environments (Dispersion [5], Navigation [5], MAMuJoCo [34], and SMAX [36]) (Table 1).
+These were specifically chosen to rigorously test performance across varying complexities, agent counts (2 to 20), and distinct behaviours (heterogeneous, homogeneous, or mixed).
+Full details in Appendix G.1.
+Baselines.
+We evaluate HyperMARL against modern parameter sharing (PS) and diversity-promoting baselines.
+Core PS comparisons use FuPS+ID and NoPS.
+For specialisation, we include privileged baselines: DiCo [6] (shared and non-shared weights, preset diversity levels); HAPPO [49] (shared critic, sequential actor updates); and Kaleidoscope [26] (learnable masks, critic ensembles, diversity loss).
+We use IPPO/MAPPO [12, 48] as the underlying algorithm for all methods except Kaleidoscope (MATD3 backbone due to its intricate off-policy design; see App.
+H.1 for these results).
+Adhering to best evaluation practices [33], we  use original codebases/hyperparameters and environments for which baselines were tuned.
+HyperMARL uses identical observations and generates architectures of equivalent capacity to baselines.
+Training and evaluation (App.
+G.2.1) and hyperparameters (App.
+I) follow each baseline’s original setup.
+We detail our baseline and environment selection criteria in Table 4, with architecture details in App.
+G.2.
+(a) IPPO
+(b) MAPPO
+(c) Actor Gradient Var.
+(d) Policy Diversity
+Figure 5: Performance and gradient analysis.
+(a,b) IPPO and MAPPO on Dispersion (20M timesteps) - IQM of Mean Episode Return with 95% bootstrap CIs: Hypernetworks match NoPS performance while FuPS struggle with specialisation.
+Interval estimates in App.
+H.2.1.
+(c) Actor gradient variance: Hypernetworks achieve lower gradient variance than FuPS+ID.
+(d) Policy diversity (SND with Jensen–Shannon distance): Hypernetworks achieve NoPS-level diversity while sharing parameters.
+Measuring Policy Diversity. To measure the diversity of the policies we System Neural Diversity (SND) [7] (Equation 5) with Jensen-Shannon distance (details in App. G.2.2).
+5.2 Q1: Specialised Policy Learning
+Learning Diverse Behaviour (Dispersion) Figures 5(a) and 5(b) show that FuPS variants (IPPO-FuPS, MAPPO-FuPS – ($\bullet$)) can struggle to learn the diverse policies required by Dispersion (even when running for longer - Fig.
+19), while their NoPS counterparts (IPPO-NoPS, MAPPO-NoPS–($\bullet$)) converge to the optimal policy, corroborating standard FuPS limitations to learn diverse behaviour.
+In contrast, HyperMARL (both linear and MLP variants) ($\bullet$, $\bullet$) match NoPS performance, suggesting that a shared hypernetwork can effectively enable agent specialisation.
+SND policy diversity measurements (Fig.
+5(d)) confirm FuPS variants achieve lower behavioural diversity than NoPS, while HyperMARL notably matches NoPS-level diversity.
+Figure 6: 17-agent Humanoid learning dynamics (IQM, 95% CI).
+HyperMARL, utilising a shared actor architecture, outperforms MAPPO-FuPS (non-overlapping CIs) and matches the performance of methods employing non-shared or sequential actors.
+This challenging environment is recognised for its high variance in outcomes across different methods [49].
+Gradient Variance.
+HyperMARL (IPPO and MAPPO variants) also exhibits lower mean policy gradient variance than FuPS+ID across actor parameters (Fig.
+5(c)).
+This aligns with their ability to learn diverse behaviours and supports the hypothesis that its gradient decoupling mechanism (Sec.
+4.3) enhances training stability.
+Diversity at Complexity and Scale (MAMuJoCo).
+In the challenging MAMuJoCo heterogeneous control tasks (Table 2), HyperMARL (MLP variant) is broadly competitive.
+Notably, unlike HAPPO and MAPPO (independent actors), HyperMARL uses a shared actor and parallel updates, and yet manages strong performance, even in the 17-agent Humanoid-v2 notoriously difficult heterogeneous task[49] (Fig.
+6), matching methods that employ independent actors and sequential updates.
+Table 2: Mean episode return in MAMuJoCo for MAPPO variants(IQM, 95% CI).
+HyperMARL achieves the highest IQM in 3/4 scenarios (bold), and is the only method with shared actors to demonstrate stable learning in the notoriously difficult 17-agent Humanoid environment (see Figure 6 for learning dynamics).
+* indicates CI overlap with the top score.
+Scenario HAPPO FuPS+ID Ind. Actors HyperMARL (Ours)
+Humanoid-v2 17x1 6501.15* (3015.88, 7229.79) 566.12 (536.36, 603.01) 6188.46* (5006.13, 6851.74) 6544.10 (3868.00, 6664.89)
+Walker2d-v2 2x3 4748.06* (4366.94, 6230.81) 4574.39* (4254.21, 5068.32) 4747.05* (3974.76, 6249.58) 5064.86 (4635.10, 5423.42)
+HalfCheetah-v2 2x3 6752.40* (6130.42, 7172.98) 6771.21* (6424.94, 7228.65) 6650.31* (5714.68, 7229.61) 7063.72 (6696.30, 7325.36)
+Ant-v2 4x2 6031.92* (5924.32, 6149.22) 6148.58 (5988.63, 6223.88) 6046.23* (5924.62, 6216.57) 5940.16* (5485.77, 6280.59)
+Adaptability (Navigation).
+Navigation tasks [5] assess adaptability to homogeneous, heterogeneous, and mixed goals (some agents have the same goals, others different).
+We compare HyperMARL with baselines including DiCo [6].
+While using DiCo’s optimal preset diversity for n=2 agents, we note that identifying appropriate diversity levels for DiCo with larger teams ($n>2$) via hyperparameter sweeps proved challenging (see Tables 12 and 13).
+(a) Shared Goals
+(b) Unique Goals
+(c) Mixed Goals (Half Agents Share Goals)
+Figure 7: Average Reward (IQM, 95% CI) in Navigation for IPPO Variants.
+HyperMARL adapts robustly across goal configurations—(a) shared, (b) unique, and (c) mixed.
+Both linear and MLP versions consistently match or outperform IPPO baselines and DiCo, with the margin widening as the number of agents grows.
+Sample-efficiency curves appear in App.
+H.4.
+Across all tested goal configurations (shared, unique, and mixed), HyperMARL consistently achieves strong performance (Figure 7).
+It generally matches or outperforms NoPS and FuPS, and outperforming DiCo.
+Interestingly, unlike in sparse-reward tasks like Dispersion, FuPS remains competitive with NoPS and HyperMARL in Navigation scenarios requiring diverse behaviours for smaller teams ($n\in\{2,4\}$), likely due to Navigation’s dense rewards.
+However, HyperMARL distinguishes itself as the strongest method for n=8 agents, highlighting its effectiveness in handling more complex coordination challenges.
+(a) 2s3z
+(b) 3s5z
+(c) SMACv2 10 Units
+(d) SMACv2 20 Units
+Figure 8: IQM and 95% CI of mean win rate in SMAX.
+Performance of FuPS Recurrent IPPO and MAPPO and HyperMARL (MLP) on SMAX.
+HyperMARL performs comparably to FuPS baselines across all environments, demonstrating its effectiveness in tasks requiring homogeneous behaviours and using recurrency.
+Inverval estimates in App.
+22.
+5.3 Q2: Effectiveness in Homogeneous Tasks
+SMAX.
+Finally, we evaluate HyperMARL (MLP) on SMAX, where recurrent FuPS is the established baseline [36, 48, 16].
+Figure 8 shows while some FuPS variants might exhibit marginally faster initial convergence on simpler maps, HyperMARL achieves comparable final performance on all maps, using the same GRU backbone for partial observability.
+These results highlights two points: (i) HyperMARL is fully compatible with recurrent architectures essential under partial observability, and (ii) it has no intrinsic bias toward specialisation and can converge homogenous behaviour when it is optimal (also shown with strong same-goal Navigation performance (Fig.
+7(a))), even with large observation spaces and many agents.
+Summary.
+Our empirical results confirm HyperMARL effectively addresses both research questions.
+For Q1 (Specialisation), across Dispersion, MAMuJoCo, and Navigation, HyperMARL learned specialised policies, matched NoPS-level diversity and performance where FuPS+ID struggled, and scaled to complex, high-agent-count heterogeneous tasks.
+For Q2 (Homogeneity), HyperMARL demonstrated competitive performance against strong FuPS baselines in SMAX and shared-goal Navigation, confirming its versatility.
+6 Ablations
+We ablate two critical components of HyperMARL, gradient decoupling(Sec.
+4.3) and initialisation scaling (Sec.
+4.2) on the complex Humanoid-v2 and simpler Dispersion tasks.
+Removing gradient decoupling (HyperMARL w/o GD), forcing joint observation/embedding processing, universally degrades performance (Figure 9).
+Omitting initialisation scaling (HyperMARL w/o RF (reset fan in/out), which aligns initial generated weights with conventional networks, reveals task-dependent effects: it is crucial for the 17-agent Humanoid, echoing prior findings on hypernetwork initialisation [8], but has minor impact in Dispersion.
+Thus, while principled initialisation becomes vital with increasing complexity, gradient decoupling is fundamentally essential across tested scenarios.
+(a) Humanoid w/o GD
+(b) Dispersion w/o GD
+(c) Humanoid w/o RF
+(d) Dispersion w/o RF
+Figure 9: Ablation results comparing HyperMARL to variants without gradient decoupling (w/o GD) and without reset fan in/out initialization (w/o RF) across environments.
+Gradient decoupling (a,b) is consistently critical across both environments, while initialisation scaling (c,d) shows greater importance in the complex Humanoid task but less impact in the simpler Dispersion environment.
+7 Related Work
+Hypernetworks in RL and MARL.
+Hypernetworks are effective in single-agent RL for meta-learning, multi-task learning, and continual learning [3, 4, 37, 20].
+In MARL, QMIX [35] used hypernetworks (conditioned on a global state) to mix per-agent Q-values; however, each agent’s own network remained a standard GRU.
+Parallel work, CASH [15], conditions hypernetworks on local observations and team capabilities for zero-shot generalization with heterogeneous action spaces.
+In contrast, we focus on agent-conditioned hypernetworks for adaptive specialisation within a fixed state-action setting, leveraging gradient decoupling (absent in CASH) that we find critical for specialised behaviours.
+Variants of Parameter Sharing.
+Selective Parameter Sharing (SePS) [10] shares weights between similar groups of agents, identified via trajectory clustering.
+Pruning methods [23, 26] split a single network into agent-specific subnetworks.
+Learning Diverse Policies.
+Shared parameters often limit policy diversity [10, 23, 16, 25].
+Proposed solutions include: (1) maximizing mutual information between agent IDs and trajectories [25], (2) role-based methods [43, 44], or (3) methods that use structural modifications or constraints to induce diversity in agent policies [23, 6, 26].
+Outside FuPS/NoPS, HAPPO [49] uses a non-shared centralised critic with individual actors updated sequentially to learn heterogeneous behaviours.
+8 Conclusion
+We introduced HyperMARL, an approach that uses agent-conditioned hypernetworks to generate per-agent parameters without modifying the standard RL objective or requiring preset diversity levels.
+Our results show HyperMARL can adaptively learn specialised, homogeneous, or mixed behaviours in settings with up to 20 agents.
+We also observe a link between HyperMARL’s performance and reduced policy gradient variance, underscoring the importance of decoupling observation- and agent-conditioned gradients.
+Overall, these findings establish HyperMARL as a promising architecture for diverse MARL tasks.
+We discuss limitations in App.
+A, most notably parameter count, which can be remedied by parameter-efficient hypernetworks (e.g., chunked variants [42, 9]).
+Appendix A Limitations
+Hypernetworks generate weights for target networks, which can lead to high-dimensional outputs and many parameters for deep target networks, particularly when using MLP-based hypernetworks.
+While HyperMARL uses more parameters than NoPS and FuPS for few agents, it scales almost constantly with agent count, an attractive property for large-scale MARL.
+Parameter efficiency could be improved through chunking techniques [42, 9], or low-rank weight approximations.
+This parameter overhead is often acceptable in RL/MARL given typically smaller actor-critic networks, and HyperMARL’s favorable agent scaling (see App.
+F.3).
+Appendix B Broader Impact
+This paper presents work whose goal is to advance the field of Multi-Agent Reinforcement Learning.
+There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix C Specialised Policies and Environments
 Specialisation plays a key role in MARL, yet remains under-defined, so we define specialised environments and specialised policies.
-Definition 2.1.
-An environment is specialised if:
+Definition 1.
+An environment is specialised if the following both hold:
   1. 1.
@@
     with strict inequality if the joint policies are non-symmetric (i.e., swapping any individual policy degrades performance).
-For example, consider a specialised environment such as a football game, agents adopt distinct roles — such as ”attackers” and ”defenders” — to optimise the team’s performance.
+For example, consider a specialised environment such as a football game, optimal team performance typically requires players in distinct roles (e.g., "attackers," "defenders").
 Permuting these roles (i.e., exchanging their policies) would typically lead to suboptimal results.
 Here, agents develop specialised policies by learning distinct, complementary behaviours essential for an optimal joint policy.
-Heterogeneous agents — with different action or observation spaces — are naturally specialised, while homogeneous agents can specialise by learning distinct policies critical to the team’s success.
-We analyse such specialised environments in Sections 3.1 and 5.2.
-3 Are Independent or Fully Shared Policies Sufficient?
-In this section, we introduce two simple matrix games and show that even in these basic settings, both fully independent and fully parameter shared policies have inherent limitations, highlighting the need for more nuanced approaches.
-3.1 Specialisation and Synchronisation Game
-Specialisation Game
-Player 2
-A B
-Player 1 A (0.5, 0.5) (1, 1)
-B (1, 1) (0.5, 0.5)
-Table 1: 2-player specialisation game. Nash Equilibria in blue.
-The Specialisation Game, inspired by the XOR game (Fu et al., 2022) and VMAS’s Dispersion (Bettini et al., 2022), is a two-player matrix game where each agent selects one of two food items.
-Payoffs are $0.5$ for matching choices and $1$ for different choices (payoff matrix in Table 1).
-This game has two symmetric pure Nash equilibria on the anti-diagonal of the payoff matrix and can be extended to $n$ agents, with payoffs of 1 for unique actions and $1/n$ for matching actions.
-This satisfies the conditions of a Specialised Environment (Definition 2.1).
-Synchronisation Game This is the inverse of the Specialisation Game, where players aim to choose the same food item.
-Payoffs are 1 for matching choices and 0.5 for different choices, resulting in pure Nash equilibria along the diagonal.
-For $n$ agents, matching actions receive a payoff of $1$, while differing choices receive $1/n$.
-3.2 Limitations of Fully Shared and Independent Policies
-Table 2: Final evaluation rewards (mean $\pm$ std) for REINFORCE variants in Specialisation and Synchronisation Games with 2, 4, and 8 agents.
-Bold values are within 1 std of the maximum.
-Fully shared or independent policies do not solve both games.
-Specialisation Game
-Method 2 4 8
-NoPS 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
-FuPS 0.50 $\pm$ 0.00 0.25 $\pm$ 0.00 0.13 $\pm$ 0.00
-FuPS+ID 1.00 $\pm$ 0.00 0.95 $\pm$ 0.10 0.53 $\pm$ 0.09
-Synchronisation Game
-Method 2 4 8
-NoPS 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.13 $\pm$ 0.00
-FuPS 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
-FuPS+ID 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.30 $\pm$ 0.35
-In Appendix B (Proof 14) we show that a fully shared policy (without conditioning on agent IDs) cannot learn the optimal joint behaviour in the 2-player Specialisation Game.
-To assess this limitation in practice, we compare three REINFORCE (Williams, 1992) variants in both games with $n=2,4,8$ agents: NoPS (No Parameter Sharing), FuPS (Fully Parameter Sharing), and FuPS+ID (FuPS with one-hot agent IDs).
-All policies use single-layer neural networks with controlled parameter counts (see Appendix F for details).
-Table 2 shows that NoPS consistently achieves optimal performance in the Specialisation Game, while FuPS performance declines as the number of agents increases.
-Conversely, in the Synchronisation Game, FuPS performs optimally, whereas NoPS fails with $n=8$.
-FuPS+ID initially performs well in both games but deteriorates at larger scales, indicating that ID conditioning alone is insufficient for fully distinct or uniform behaviours.
-Overall, these experiments demonstrate that neither fully shared nor fully independent policies provide a universal solution—even in simple matrix games.
-This highlights the need for more flexible MARL strategies that handle diverse task demands.
-In Section 5, we show similar findings for modern MARL algorithms and more sophisticated environments.
-4 HyperMARL
-In this section, we introduce HyperMARL, an approach that uses agent-conditioned hypernetworks to learn diverse or homogeneous policies end-to-end, without modifying the standard RL objective or requiring prior knowledge of the task’s optimal diversity.
-By operating within a fully shared paradigm, HyperMARL leverages shared gradient information while enabling specialisation through the decoupling of observation- and agent-conditioned gradients.
-We present the pseudocode in Section C.1.
-4.1 Hypernetworks for MARL
-Hypernetworks (Ha et al., 2016) are neural networks that generate the weights of another network conditioned on an input context vector.
-In HyperMARL, we use them to adaptively produce per-agent policy and critic parameters.
-As illustrated in Figure 1, for any agent $i$ with context $e^{i}$ (i.e., either a one-hot encoded ID or a learned embedding), the hypernetworks generate the agent-specific parameters:
-$\theta^{i}=h_{\psi}^{\pi}(e^{i}),\quad\phi^{i}=h_{\varphi}^{V}(e^{i}),$ (4)
-where $h_{\psi}^{\pi}$ and $h_{\varphi}^{V}$ are the hypernetworks for the policy and critic, respectively.
-The parameters $\theta^{i}$ and $\phi^{i}$ define each agent’s policy $\pi_{\theta^{i}}$ and critic $V_{\phi^{i}}$, dynamically enabling either specialised or homogeneous behaviours as required by the task.
-We explore two types of hypernetworks in HyperMARL: Linear and Multilayer Perceptron (MLP) Hypernetworks, balancing expressiveness and computational efficiency.
-Linear Hypernetworks Linear hypernetworks are a straightforward instantiation of HyperMARL.
-Given a one-hot agent ID, $\mathds{1}^{i}\in\mathbb{R}^{1\times n}$, a linear hypernetwork $h_{\psi}^{\pi}$ generates agent-specific parameters $\theta^{i}$ as  [^3] For conciseness we only show the policy parameters in this section.:
-$\theta^{i}=h_{\psi}^{\pi}(\mathds{1}^{i})=\mathds{1}^{i}\cdot W+b$ (5)
-where $W\in\mathbb{R}^{n\times m}$ is the weight matrix (with $m$ the per-agent parameter dimensionality and $n$ is the number of agents), and $b\in\mathbb{R}^{1\times m}$ is the bias vector.
-Since $\mathds{1}^{i}$ is one-hot encoded, each $\theta^{i}$ corresponds to a specific row of $W$ plus the shared bias $b$.
-If there is no shared bias term, this effectively replicates training of separate policies for each task (in our case, for each agent) (Beck et al., 2023), since there are no shared parameters and gradient updates are independent.
-MLP Hypernetworks for Expressiveness While linear hypernetworks can be effective (Beck et al., 2023; Chang et al., 2020), they may lack the capacity to capture complex relationships between agents’ roles and their corresponding weights.
-To enhance expressiveness, we incorporate hidden layers and non-linear activations (e.g., ReLU) into the hypernetworks as follows:
-$\theta^{i}=h_{\psi}^{\pi}(e^{i})=f_{\psi_{1}}^{\pi}\bigl{(}g_{\psi_{2}}^{\pi}(e^{i})\bigr{)}$ (6)
-where $e^{i}$ is the agent ID or embedding for agent $i$, $g_{\psi_{2}}^{\pi}$ is the multilayer perceptrons (MLPs) that produce intermediate representations with non-linear activations, and $f_{\psi_{1}}^{\pi}$ and is the final linear transformations that generate the policy parameters.
-This additional expressiveness may be critical in domains where roles overlap or where nuanced specialisations emerge dynamically.
-Unlike linear hypernetworks with one-hot agent IDs, MLP hypernetworks do not guarantee distinct weights for each agent.
-Additionally, they increase the total number of trainable parameters, necessitating a careful balance between expressiveness and computational overhead.
-4.2 Agent Embeddings
-The agent embedding $e^{i}$ represents the input to the hypernetwork.
-For our linear hypernetworks, we use one-hot encoded agent IDs as embeddings, resulting in a mix of agent-independent weights and shared weights as discussed in Section 4.1.
-For our MLP hypernetworks, we use learned agent embeddings, initialised orthogonally.
-These embeddings are optimised end-to-end alongside the hypernetwork, allowing agents to learn distinct or similar embeddings based on whether homogeneous or specialised policies are required.
-4.3 Hypernetwork Initialisation
-HyperMARL initialises its hypernetworks to ensure that the initial agent-specific parameters $\theta^{i}$ and $\phi^{i}$ adhere to the same distribution as standard networks without hypernetworks.
-For example, PPO commonly uses orthogonal initialisation (Engstrom et al., 2019), so HyperMARL generates orthogonally initialised weights for each agent, preserving the same fan in/out as the conventional network.
-This alignment with standard deep RL initialisation schemes facilitates stable learning and can be applied using any initialisation strategy.
-4.4 Gradient Decoupling in HyperMARL
-A core difficulty in FuPS is cross-agent gradient interference (Christianos et al., 2021; Zhong et al., 2024). In HyperMARL, gradient decoupling helps mitigate this issue.
-Consider a fully cooperative MARL setting with a centralised critic, we can formulate the policy gradient for agent $i$ as follows (Albrecht et al., 2024; Kuba et al., 2021):
-$\nabla_{\theta^{i}}J(\theta^{i})\;=\;\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\Bigl{[}A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\Bigr{]},$
-where $\mathbf{h_{t}}$ and $\mathbf{a_{t}}$ are the joint histories and joint actions for all agents, $\theta^{i}$ denotes the parameters of agent $i$, and $A(\mathbf{h_{t}},\mathbf{a_{t}})=Q(\mathbf{h_{t}},\mathbf{a_{t}})-V(\mathbf{h_{t}})$ is the advantage function.
-When using shared parameters $\theta$, this can reframed as:
-$\nabla_{\theta}J(\theta)=\sum_{i=1}^{I}\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\left[A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\theta}\log\pi_{\theta}\left(a_{t}^{i}\mid h_{t}^{i},id\right)\right],$ (7)
-where $id$ is our agent ID.
-In HyperMARL, each agent’s policy parameters $\theta^{i}$ are generated by a hypernetwork $h_{\psi}^{\pi}$ conditioned on an agent embedding $e^{i}$:
-$\theta^{i}\;=\;h_{\psi}^{\pi}\bigl{(}e^{i}\bigr{)}.$
-Hence, the gradient with respect to $\psi$ becomes:
-$\nabla_{\psi}J(\psi)\;=\;\sum_{i=1}^{I}\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\Bigl{[}A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\psi}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\Bigr{]}.$
-Applying the chain rule:
-$\nabla_{\psi}\,\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\;=\;\underbrace{\nabla_{\theta^{i}}\,\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})}_{\text{Observation-Conditioned}}\;\cdot\;\underbrace{\nabla_{\psi}\,h_{\psi}^{\pi}(e^{i})}_{\text{Agent-Conditioned}},$
-leading to
-$\nabla_{\psi}J(\psi)$ $=\sum_{i=1}^{I}\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\Bigl{[}A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})$
-$\quad\cdot\nabla_{\psi}h_{\psi}^{\pi}(e^{i})\Bigr{]}.$ (8)
-This factorisation separates:
-Observation-Conditioned Gradients ($\nabla_{\theta^{i}}\!\log\pi_{\theta^{i}}$), which reflect how each policy updates in response to local observations.
-Agent-Conditioned Gradients ($\nabla_{\psi}\,h_{\psi}^{\pi}(e^{i})$), which reflect how the hypernetwork generates each agent’s parameters from each agent’s embedding $e^{i}$.
-These gradients thus remain partially decoupled, reducing cross-agent interference.
-Related work in Meta-RL shows that similar decoupling lowers policy gradient variance and stabilises learning (Sarafian et al., 2021).
-In Section 5.2, we observe a comparable reduction in variance with HyperMARL.
-Additionally, ablation studies in Section 6 show that removing gradient decoupling significantly degrades performance, highlighting its essential role in HyperMARL’s effectiveness.
-5 Experiments
-We evaluate HyperMARL on two primary questions:
-Q1: Specialised Policy Learning: Can HyperMARL effectively learn specialised policies (see Section 2.1) using a shared hypernetwork, without altering the learning objective or relying on preset diversity levels?
-Q2: Effectiveness in Homogeneous Tasks: Is HyperMARL competitive in environments that necessitate homogeneous behaviours?
-5.1 Experimental Setup
-Table 3: Summary of MARL environments used to evaluate HyperMARL.
-Environment Agents Action Space Behaviour
-Dispersion 4 Discrete Heterogeneous
-Navigation 2, 4, 8 Continuous Homogeneous, Heterogeneous, Mixed
-MAMuJoCo 2–17 Continuous Heterogeneous
-SMAX 2–20 Discrete Homogeneous
-Environments We use diverse cooperative MARL tasks (Table 3), varying in complexity, agent count (2 to 20), and required behaviours (homogeneous, heterogeneous, or mixed).
-Full details are in Appendix D.1.
-MARL Baselines We focus on PPO variants, given their success in MARL (De Witt et al., 2020; Yu et al., 2022).
-Comparisons include IPPO (De Witt et al., 2020), which learns an actor and critic conditioned on individual observations, and MAPPO (Yu et al., 2022), which trains a centralised critic conditioned on the global state or concatenated agent observations.
-We use NoPS and FuPS variants of these methods, with all FuPS variants being conditioned on one-hot encoded agent IDs.
-(a) IPPO
-(b) MAPPO
-Figure 2: Performance of IPPO and MAPPO on Dispersion after 20 million timesteps.
-We show the Interquartile Mean (IQM) of the Mean Episode Return and the 95% Stratified Bootstrap Confidence Intervals (CI) using (Agarwal et al., 2021).
-Hypernetworks achieve comparable performance to NoPS, while FuPS struggle with specialisation.
-Sample efficiency plots are provided in Appendix E.1.1.
-(a) Policy Diversity using SND with Jenson-Shannon distance. Hypernetworks match NoPS-level diversity while sharing parameters. Bold indicates highest SND, asterisks show overlaps in std.
-Method Algorithm Mean ± Std Dev
-IPPO MLP Hypernet $6.3\times 10^{-4}\pm 2.0\times 10^{-5}$
-Linear Hypernet $6.4\times 10^{-4}\pm 7.0\times 10^{-6}$
-NoPS $\mathbf{6.9\times 10^{-4}\pm 7.0\times 10^{-6}}$
-FuPS $2.4\times 10^{-4}\pm 1.1\times 10^{-4}$
-MAPPO MLP Hypernet $\mathbf{6.7\times 10^{-4}\pm 5.0\times 10^{-6}}$
-Linear Hypernet $6.6\times 10^{-4}\pm 1.2\times 10^{-5*}$
-NoPS $6.6\times 10^{-4}\pm 1.5\times 10^{-5*}$
-FuPS $3.7\times 10^{-4}\pm 8.1\times 10^{-5}$
-(b) Average policy gradient variance across IPPO (left) and MAPPO (right). Our hypernetworks have lower variance than FuPS.
-Privileged Baselines Beyond NoPS, we include two privileged baselines designed for heterogeneous policies.
-Diversity Control (DiCo) (Bettini et al., 2024) partitions each policy into shared and agent-specific components and enforces a target diversity level (via Wasserstein distances).
-Heterogeneous-Agent PPO (HAPPO) (Zhong et al., 2024) uses a shared centralised critic and updates individual actors sequentially, incorporating prior actor updates to allow agent heterogeneity.
-In contrast, HyperMARL uses fully shared weights, parallel updates, and no preset diversity levels.
-Table 7 summarises baseline selection criteria.
-Implementation and hyperparameters follow authors’ defaults unless otherwise stated (see Appendix F), and HyperMARL uses the same observations and generates architectures with equivalent capacity to the baselines.
-Training and evaluation protocols are consistent across all experiments (Appendix D.1.1).
-5.1.1 Measuring Policy Diversity
-To measure the diversity of the policies we System Neural Diversity (SND) (Bettini et al., 2023) (Equation 9) with Jensen-Shannon distance, where 0 indicates identical policies and 1 indicates maximal diversity.
-Details in Appendix D.1.2.
-5.2 Q1: Specialised Policy Learning
-Learning Diverse Behaviour (Dispersion) Figures 2(a) and 2(b) show that parameter sharing variants (IPPO-FuPS, MAPPO-FuPS – ($\bullet$)) fail to learn the diverse policies required by Dispersion, while their non-parameter sharing counterparts (IPPO-NoPS, MAPPO-NoPS–($\bullet$)) converge to the optimal policy.
-These results corroborate that parameter sharing methods struggle in environments requiring agent specialisation.
-In contrast, our shared linear and MLP hypernetworks ($\bullet$, $\bullet$) match NoPS performance, suggesting that a shared hypernetwork can effectively support agent specialisation.
-Policy Diversity SND measurements (Table 3(a)) show that IPPO-FuPS and MAPPO-FuPS exhibit lower behavioural diversity than their NoPS counterparts, aligning with their weaker performance results.
-Notably, both linear and MLP hypernetworks achieve diversity levels comparable to NoPS despite using a shared architecture.
-Linear hypernetworks with one-hot agent IDs enforce weight seperation, while MLP hypernetworks have to dynamically learn diverse policies without explicit separation of weights or gradients.
-Gradient Variance To investigate HyperMARL’s effect on training stability, we compute the average policy gradient variance by calculating the variance across all actor parameters at each update and averaging it over training.
-As shown in Figure 3(b), hypernetworks in both IPPO and MAPPO exhibit lower mean policy gradient variance than FuPS.
-This aligns with their ability to learn diverse behaviours and may stem from decoupling observation- and agent-condition gradients (Section 4.4), similar to Meta-RL findings (Sarafian et al., 2021).
-Table 4: Mean episode return in MAMuJoCo for MAPPO variants.
-IQM of the mean episode returns with 95% stratified bootstrap CI.
-Bold denotes the highest score per scenario; * indicates overlap with the highest score within one CI.
-HyperMARL (MLP variant) is competitive with methods that use independent actors.
-Scenario HAPPO FuPS Ind. Actors HyperMARL
-Humanoid-v2 17x1 6501.15* (3015.88, 7229.79) 566.12 (536.36, 603.01) 6188.46* (5006.13, 6851.74) 6544.10 (3868.00, 6664.89)
-Walker2d-v2 2x3 4748.06* (4366.94, 6230.81) 4574.39* (4254.21, 5068.32) 4747.05* (3974.76, 6249.58) 5064.86 (4635.10, 5423.42)
-HalfCheetah-v2 2x3 6752.40* (6130.42, 7172.98) 6771.21* (6424.94, 7228.65) 6650.31* (5714.68, 7229.61) 7063.72 (6696.30, 7325.36)
-Ant-v2 4x2 6031.92* (5924.32, 6149.22) 6148.58 (5988.63, 6223.88) 6046.23* (5924.62, 6216.57) 5940.16* (5485.77, 6280.59)
-(a) All Agents with Same Goals
-(b) All Agents with Different Goals
-(c) Mixed Goals (Half Agents Share Goals)
-Figure 4: Average Reward in Navigation for IPPO Variants.
-IQM and 95% CI of mean reward for IPPO Baselines, HyperMARL (linear and MLP), and DiCo in the Navigation environment across goal configurations.
-HyperMARL maintains strong IQM performance, particularly with more agents.
-Sample efficiency plots in Appendix E.3.
-Diversity at Complexity and Scale (MAMuJoCo) Dispersion is designed to test diverse behaviours (Bettini et al., 2022), but we further evaluate HyperMARL on the more challenging heterogeneous MAMuJoCo environment.
-Following Zhong et al.
-(2024), we use a centralised but non-shared critic, applying HyperMARL only to the actor.
-Table 4 shows that HyperMARL (MLP variant) is competitive with all baselines across various MAMuJoCo scenarios.
-Unlike HAPPO and MAPPO (independent actors), HyperMARL uses a shared actor and parallel updates, and yet manages strong performance.
-This is particularly notable in Humanoid-v2 17x1(Figure 5), where 17 agents to coordinate in a Humanoid control task, a notoriously difficult heterogeneous task(Zhong et al., 2024).
-This demonstrates HyperMARL’s ability to handle large-scale specialisation with a shared architecture.
-Figure 5: Performance on the 17-agent Humanoid task.
-IQM of the mean episode returns with 95% stratified bootstrap CI.
-HyperMARL is able to match the performance of methods that leverage independent actors, while using a shared actor in this challenging heterogeneous task.
-Adaptability (Navigation) We evaluate HyperMARL against DiCo (Bettini et al., 2024) in the Navigation environment (Bettini et al., 2022).
-For two agents ($n=2$), we adopt DiCo’s optimal diversity levels and replicate its hyperparameters.
-For larger teams ($n>2$), we perform hyperparameter sweeps for diversity levels and learning parameters across all methods (see Tables 14 and 15).
-Although DiCo defines optimal diversity for two agents, identifying suitable levels for larger team sizes proved challenging.
-Figure 4 presents the IQM of the mean reward across different goal setups.
-HyperMARL consistently demonstrates higher average returns across all configurations, outperforming DiCo and matching or exceeding NoPS and FuPS performance for most settings.
-Unlike in Dispersion, IPPO-FuPS is competitive with NoPS and HyperMARL in the Navigation environment for tasks requiring diverse behaviours (mixed or different goals – Figures 4(b) and 4(c)) with two and four agents.
-We hypothesise that this arises from the dense rewards in Navigation, in contrast to the sparse rewards in Dispersion.
-However, when scaling to eight agents, HyperMARL outperforms all baselines, demonstrating its effectiveness in handling more complex coordination challenges.
-Figure 6: IQM and 95% CI of mean return in SMAX. Performance of FuPS Recurrent IPPO and MAPPO, HyperMARL (MLP) on SMAX. HyperMARL performs comparably to these baselines.
-5.3 Q2: Effectiveness in Homogeneous Tasks
-SMAX  Finally, we test HyperMARL (MLP variant) on SMAX, comparing against the FuPS baseline from JaxMARL.
-Prior work (Yu et al., 2022; Fu et al., 2022) has found FuPS optimal in such environments.
-Figure 6 shows that across simple (2s3z), hard (3s5z), and SMACv2-style maps (smacv2_10, smacv2_20), HyperMARL is comparable to the FuPS baseline.
-Alongside results in Navigation (Figure 4(a)), this indicates that HyperMARL is also effective in tasks requiring homogeneous behaviours, large observation spaces, and coordination among many agents, reinforcing its effectiveness across diverse MARL scenarios.
-6 Ablations
-To better understand the importance of the different components in HyperMARL, we conduct an ablation study on four key variations of HyperMARL (MAPPO-MLP) in the Humanoid-v2 17x1 (Section 5.2).
-First, HyperMARL-S reduces the hypernetwork hidden layer size from 64 to 32, investigating the impact of network capacity.
-Second, HyperMARL w/ One-Hot replaces learned agent embeddings with one-hot IDs (Section 4.2).
-Third, HyperMARL w/o RF omits the initialisation scaling described in Section 4.3.
-Finally, HyperMARL w/o GD removes gradient decoupling (Section  4.4) by conditioning the hypernetwork on both observations and agent embeddings.
-Figure 7 shows that HyperMARL-S converges more slowly but ultimately reaches an IQM close to the baseline HyperMARL, while learned agent embeddings offer better final performance than one-hot IDs (HyperMARL w/ One-Hot).
-Crucially, removing either gradient decoupling (HyperMARL w/o GD) or initialisation scaling (HyperMARL w/o RF) severely impairs learning, confirming their essential roles in HyperMARL ability to learn specialised policies.
-We see similar findings in ablations in Dispersion (Section 18).
-Figure 7: Ablation results comparing HyperMARL and its variants in Humanoid-v2 17x1. Gradient decoupling and initialization are crucial for maintaining performance.
-7 Related Work
-Hypernetworks in RL and MARL.
-Hypernetworks have proven effective in meta-learning, multi-task learning, and continual learning in single-agent settings (Beck et al., 2023, 2024; Sarafian et al., 2021; Huang et al., 2021).
-In MARL, QMIX (Rashid et al., 2020) used hypernetworks (conditioned on a global state) to mix per-agent Q-values; however, each agent’s own network remained a standard MLP or GRU.
-Parallel work, CASH (Fu et al., 2025), conditions hypernetworks on local observations and team capabilities for zero-shot generalization with heterogeneous action spaces.
-In contrast, we focus on agent-conditioned hypernetworks for adaptive specialisation within a fixed state-action setting, leveraging gradient decoupling (absent in CASH) that we find critical for specialised behaviours.
-Variants of Parameter Sharing.
-Selective Parameter Sharing (SePS) (Christianos et al., 2021) shares weights between similar groups of agents, identified via trajectory clustering.
-Pruning methods (Kim & Sung, 2023; Li et al., 2024) split a single network into agent-specific subnetworks.
-Learning Diverse Policies.
-Shared parameters often limit policy diversity (Christianos et al., 2021; Kim & Sung, 2023; Fu et al., 2022; Li et al., 2021).
-Proposed solutions include: (1) maximizing mutual information between agent IDs and trajectories (Li et al., 2021), (2) role-based methods (Wang et al., 2020a, b), or (3) methods that use structural modifications or constraints to induce diversity in agent policies (Kim & Sung, 2023; Bettini et al., 2024; Li et al., 2024).
-Outside FuPS/NoPS, HAPPO (Zhong et al., 2024) uses a non-shared centralised critic with individual actors updated sequentially to learn heterogeneous behaviours.
-8 Conclusion and Discussion
-We introduced HyperMARL, an approach that uses agent-conditioned hypernetworks to generate per-agent actor and critic parameters without modifying the standard RL objective or requiring preset diversity levels.
-Our results show it can adaptively learn specialised, homogeneous, or mixed behaviours in settings with up to 20 agents.
-We also observe a link between HyperMARL’s performance and reduced policy gradient variance, underscoring the importance of decoupling observation- and agent-conditioned gradients.
-Overall, these findings establish HyperMARL as a versatile and promising architecture for diverse MARL tasks.
-Future work could explore more parameter-efficient hypernetworks (e.g., chunked variants (von Oswald et al., 2020; Chauhan et al., 2024)) to enable large-scale applications.
-Impact Statement
-This paper presents work whose goal is to advance the field of Multi-Agent Reinforcement Learning.
-There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Measuring Behavioural Diversity
-A.1 Quantifying Team Diversity
-We quantify policy diversity using System Neural Diversity (SND) (Bettini et al., 2023), which measures behavioural diversity based on differences in policy outputs:
+While agents with heterogeneous capabilities (e.g., different action spaces) are inherently specialised, homogeneous agents can also learn distinct policies.
+Such environments are analysed in Sections E.1 and 5.2.
+Appendix D Measuring Behavioural Diversity
+D.1 Quantifying Team Diversity
+We quantify policy diversity using System Neural Diversity (SND) [7], which measures behavioural diversity based on differences in policy outputs:
 $\operatorname{SND}\left(\left\{\pi^{i}\right\}_{i\in\mathbb{I}}\right)=\frac{2}{n(n-1)|\mathcal{O}|}\sum_{i=1}^{n}\sum_{j=i+1}^{n}\sum_{o\in\mathcal{O}}D\left(\pi^{i}(o),\pi^{j}(o)\right).$ (5)
 where $n$ is the number of agents, $\mathcal{O}$ is a set of observations typically collected via policy rollouts, $\pi^{i}(o_{t})$ and $\pi^{j}(o_{t})$ are the outputs of policies $i$ and $j$ for observation $o_{t}$, and $D$ is our distance function between two probability distributions.
-In contrast to  (Bettini et al., 2023), we use Jensen-Shannon Distance (JSD) (Endres & Schindelin, 2003; Lin, 1991) as $D$, rather than the Wasserstein metric (Vaserstein, 1969).
-As shown in Appendix A.2, JSD is a robust metric for both continuous and discrete cases, and provides a more reliable measure of policy distance.
-A.2 Finding a Suitable Distance Function for Policy Diversity
+In contrast to  [7], we use Jensen-Shannon Distance (JSD) [13, 27] as $D$, rather than the Wasserstein metric [41].
+As shown in Appendix D.2, JSD is a robust metric for both continuous and discrete cases, and provides a more reliable measure of policy distance.
+D.2 Finding a Suitable Distance Function for Policy Diversity
 The choice of distance function $D$ in Equation 5 is crucial for accurately measuring policy diversity.
 In MARL, policies are often represented as probability distributions over actions, making the choice of distance function non-trivial.
-(Bettini et al., 2024) use the Wasserstein metric for continuous policies (Vaserstein, 1969) as distance function $D$, while (McKee et al., 2022) use the total variation distance for discrete policies.
+[6] use the Wasserstein metric for continuous policies [41] as distance function $D$, while [29] use the total variation distance for discrete policies.
 For discrete policies, Wasserstein distance would require a cost function representing the cost of changing from one action to another, which might not be trivial to come up with.
@@
 This is because TVD only measures the largest difference between the two distributions, and does not consider the overall difference between them.
-On the other hand, the Jensen-Shannon distance (JSD) (Endres & Schindelin, 2003), which is the square root of the Jensen-Shannon divergence, does not have this problem as it is a smooth distance function.
+On the other hand, the Jensen-Shannon distance (JSD) [13], which is the square root of the Jensen-Shannon divergence, does not have this problem as it is a smooth distance function.
 Furthermore, it satisfies the conditions for being a metric – it is non-negative, symmetry, and it satisfies the triangle inequality.
@@
 Total Variation Distance Discrete Metric No $TV(p,q)=\frac{1}{2}\sum_{x}|p(x)-q(x)|$
-Jensen-Shannon Divergence (Lin, 1991) Both Divergence Yes $JSD(p\parallel q)=\frac{1}{2}D_{KL}(p\parallel m)+\frac{1}{2}D_{KL}(q\parallel m),\,m=\frac{1}{2}(p+q)$
-Jensen-Shannon Distance (Endres & Schindelin, 2003) Both Metric Yes $\sqrt{JSD(p\parallel q)}$
+Jensen-Shannon Divergence [27] Both Divergence Yes $JSD(p\parallel q)=\frac{1}{2}D_{KL}(p\parallel m)+\frac{1}{2}D_{KL}(q\parallel m),\,m=\frac{1}{2}(p+q)$
+Jensen-Shannon Distance [13] Both Metric Yes $\sqrt{JSD(p\parallel q)}$
 Table 3: Measure Policy Diversity
-Appendix B Specialisation and Synchronisation Game
-Theorem B.1.
-A shared policy cannot learn the optimal behaviour for the two-player Specialisation Game.
+Appendix E Specialisation and Synchronisation Games
+To study the challenges of specialisation and coordination in an isolated setting, we introduce the Specialisation and Synchronisation Games, drawing inspiration from a version of the XOR game [16], VMAS’s Dispersion [5] and coordination and anti-coordination games in game theory [32].
+These environments encourage agents to take distinct actions (Specialisation) or take identical actions (Synchronisation).
+Despite their deceptively simple payoff structure, these games present substantial learning challenges – non-stationary reward distributions driven by others’ adapting behaviours and in their temporal extension, the joint observation spaces grows exponentially with the number of agents.
+E.1 Specialisation and Synchronisation Games Description
+Specialisation Game.
+Agents are encouraged to choose distinct actions.
+In the simplest setting, it is a two-player matrix game where each agent selects between two actions ($A$ or $B$).
+As shown in Figure 2(a), agents receive a payoff of $1.0$ when their actions differ (creating two pure Nash equilibria on the anti-diagonal) and $0.5$ when they match.
+This structure satisfies Definition 1, since optimal joint policies require complementary, non-identical strategies.
+There is also a symmetric mixed-strategy equilibrium in which each agent plays $A$ and $B$ with probability $0.5$.
+Synchronisation Game.
+Conversely, agents are encouraged to coordinate and choose identical actions.
+The payoff matrix inverts the Specialisation game’s structure, agents receive $1$ for matching actions and $0.5$ for differing ones.
+This creates two pure Nash equilibria along the diagonal of the payoff matrix (Figure 2(b)), and incentivises uniform behaviour across agents.
+$N$-Agent Extension.
+Both games naturally scale to $n$ agents and $n$ possible actions.
+In Specialisation, unique actions receive a payoff of $1.0$, while selecting the same action receives payoffs of $\frac{1}{k}$, where $k$ is the number of agents choosing that action.
+In contrast, in Synchronisation, agents receive maximum payoffs ($1.0$) only when all actions match.
+For partial coordination, rewards follow a hyperbolic scale, $\frac{1}{n-k+1}$, encouraging agents to align their choices.
+Visualisations in Figure 2 and detailed reward profiles appear in Figure 12.
+E.2 General-$n$ Payoff Definitions
+Both games generalise naturally to $n$ agents and $n$ possible actions. We show the reward profiles for $n=5$ agents in Figure 12.
+Let $\mathbf{a}=(a^{1},\dots,a^{n})\in\{1,\dots,n\}^{n}$ and $k_{a}=\bigl{|}\{\,j:a^{j}=a\}\bigr{|}$ be the joint action profile and the count of agents choosing action $a$, respectively.
+Temporal Extension.
+To model sequential decision-making, we extend each normal-form game into a repeated Markov game, where the state at time $t$ is the joint action at time $t-1$.
+At each step $t$ all agents observe $a^{t-1}$, choose $a_{i}^{t}$, and receive the original Specialisation or Synchronisation payoff.
+This repeated setup isolates how agents adapt based on past joint actions, exposing temporal patterns of specialisation and coordination.
+E.2.1 Specialisation Game
+The reward is formulated as follows:
+$r_{\mathrm{spec}}^{\,i}(\mathbf{a})=\begin{cases}1.0,&\text{if }k_{a^{\,i}}=1\quad\bigl{(}\text{unique action}\bigr{)};\\[6.0pt] \displaystyle\frac{1}{k_{a^{\,i}}},&\text{if }k_{a^{\,i}}>1\quad\bigl{(}\text{shared action}\bigr{)}.\end{cases}$
+E.2.2 Synchronisation Game
+The reward is formulated as follows:
+$r_{\mathrm{sync}}^{\,i}(\mathbf{a})=\frac{1}{\,n-k_{a^{\,i}}+1\,},$
+so that $r_{\mathrm{sync}}^{\,i}=1.0$ when $k_{a^{\,i}}=n$ (all agents select the same action), and otherwise follows a hyperbolic scale encouraging consensus.
+Figure 12: Reward profiles for the Specialisation (blue) and Synchronisation (orange) games with $n=5$ agents.
+In the Specialisation game, an agent’s payoff peaks when it selects a unique action, and then decays as when actions are shared.
+In the Synchronisation game, payoffs follow a hyperbolic scale $1/(n-k+1)$, reaching maximum only under full consensus, thereby incentivising alignment.
+E.3 Proof that FuPS cannot represent the optimal policy in the two-player Specialisation Game
+Theorem 1.
+A stochastic, shared policy without agent IDs cannot learn the optimal behaviour for the two-player Specialisation Game.
 Proof.
@@
 ∎
-In Table 6, we present the results for the Specialisation and Synchronisation Games when actions are sampled during evaluation rather than selecting the argmax, as was done in Table 2.
-Table 6: Average Evaluation Reward at the end of training (mean $\pm$ std) for REINFORCE variants in the Specialisation and Synchronisation Games, across 2, 4, and 8 Agents.
-Bold indicates values within 1 std of the maximum.
-Actions are sampled during evaluation.
-Specialisation Game
-Method 2 4 8
-NoPS .9996 $\pm$ .0004 .9988 $\pm$ .0007 .9989 $\pm$ .0008
-FuPS .7493 $\pm$ .0062 .6826 $\pm$ .0039 .6545 $\pm$ .0025
-FuPS+ID .$9$$\pm$ .0002 .9995 $\pm$ .0003 .7275 $\pm$ .0293
-Synchronisation Game
-Method 2 4 8
-NoPS .9993 $\pm$ .0004 .9976 $\pm$ .0022 .1250 $\pm$ 0
-FuPS .$9$$\pm$ .0002 1.0 $\pm$ 0 .3000 $\pm$ .3500
-FuPS+ID .$9$$\pm$ 0 .$9$$\pm$ .0003 .1250 $\pm$ 0
-Appendix C HyperMARL Details
-C.1 HyperMARL Pseudocode
+E.4 Measuring Agent Gradient Conflict
+We measure gradient conflicts, via the cosine similarity between agents’ gradients $\cos\bigl{(}g_{t}^{(i)},g_{t}^{(j)}\bigr{)}=\frac{\langle g_{t}^{(i)},g_{t}^{(j)}\rangle}{\|g_{t}^{(i)}\|\|g_{t}^{(j)}\|}$, where $g_{t}^{(i)}=\nabla_{\theta}\mathcal{L}^{(i)}(\theta_{t})$.
+E.5 Empirical Results in N-player Specialisation and Synchronisation Normal-Form Game
+To assess this limitations of FuPS and NoPS in practice, we compare three REINFORCE [46] variants in both games with $n=2,4,8,16,32$ agents: NoPS (No Parameter Sharing), FuPS (Fully Parameter Sharing), and FuPS+ID (FuPS with one-hot agent IDs).
+All policies use single-layer neural networks with controlled parameter counts (see Appendix I for details).
+Figure 13: Mean evaluation reward (mean ± standard error) as a function of the number of agents/actions in the Specialisation (left) and Synchronisation (right) games.
+In the Specialisation game, vanilla policy gradients (PG, i.e.
+REINFORCE) with FuPS collapse as the team grows, whereas our identity-aware variant (PG-FuPS+ID) retains near-optimal performance.
+In the Synchronisation game, PG-NoPS performs well at small scales but degrades with more agents, while both PG-FuPS and PG-FuPS+ID remain at optimal reward across all scales.
+Appendix F HyperMARL Details
+F.1 HyperMARL Pseudocode
 In Algorithm 1, we present the pseudocode for HyperMARL, with HyperMARL-specific steps highlighted in blue.
@@
 1:  Input: Number of agents $n$, number of training iterations $K$, MARL algorithm parameters (e.g., MAPPO-specific hyperparameters)
-2:  Initialize:
+2:  Initialise:
 3:   Hypernetwork parameters $\psi,\varphi$Ensure $\theta^{i}$ and $\phi^{i}$ follow standard initialization schemes, e.g., orthogonal}
-4:   Agent embeddings $\{e^{i}\}_{i=1}^{n}$One-hot or orthogonally initialized learnable parameters}
+4:   Agent embeddings $\{e^{i}\}_{i=1}^{n}$One-hot or orthogonally initialised learnable parameters}
 5:  Output: Optimized joint policy $\boldsymbol{\pi}$
@@
 15:  Return $\boldsymbol{\pi}=(\pi^{1},\dots,\pi^{n})$
-C.2 Scalability and Parameter Efficiency
+F.2 Variance of the HyperMARL Gradient Estimator
+Unbiased estimator.
+Following from Equation 4, we can write the unbiased estimator for HyperMARLs gradient as follows:
+$\widehat{g}_{\mathrm{HM}}=\sum_{i=1}^{I}\underbrace{\nabla_{\!\psi}h_{\psi}^{\pi}(e^{i})}_{\mathbf{J}_{i}}\;\cdot\;\underbrace{\Bigl{(}\frac{1}{B}\sum_{b=1}^{B}\sum_{t=0}^{T-1}A(\mathbf{h}_{t}^{(b)},\mathbf{a}_{t}^{(b)})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i,(b)}\!\mid h_{t}^{i,(b)})\Bigr{)}}_{\widehat{Z}_{i}}=\sum_{i=1}^{I}\mathbf{J}_{i}\,\widehat{Z}_{i},$
+where $B$ trajectories $\{\tau^{(b)}\}_{b=1}^{B}$ are sampled i.i.d. and $\widehat{Z}_{i}$ is the empirical analogue of the observation-conditioned factor.
+Assumptions.
+(A1) trajectories are i.i.d.; (A2) all second moments are finite; (A3) $\psi,\theta,e^{i}$ are fixed during the backward pass.
+Variance expansion.
+Since each $\mathbf{J}_{i}$ is deterministic under (A3), we may factor them outside the covariance:
+$\operatorname{Var}\bigl{(}\widehat{g}_{\text{\tiny HM}}\bigr{)}$ $=\operatorname{Cov}\!\Bigl{(}\sum_{i}\mathbf{J}_{i}\widehat{Z}_{i},\;\sum_{j}\mathbf{J}_{j}\widehat{Z}_{j}\Bigr{)}$ (by def.
+$\operatorname{Var}(X)=\operatorname{Cov}(X,X)$)
+$=\sum_{i,j}\operatorname{Cov}\bigl{(}\mathbf{J}_{i}\widehat{Z}_{i},\;\mathbf{J}_{j}\widehat{Z}_{j}\bigr{)}$ (bilinearity of $\operatorname{Cov}$)
+$=\sum_{i,j}\mathbf{J}_{i}\;\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})\;\mathbf{J}_{j}^{\!\top}$ (pull deterministic matrices out of $\operatorname{Cov}$)
+Equation (LABEL:eq:variance_split_detail) makes explicit that all trajectory-induced covariance is captured $\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})$, while the agent-conditioned Jacobians $\mathbf{J}_{i}$ remain trajectory noise-free.
+Mini-batch update and covariance.
+Let $\widehat{Z}_{i}$ be the unbiased mini-batch estimate of $Z_{i}$ and $\widehat{g}_{\text{\tiny HM}}=\sum_{i}\mathbf{J}_{i}\widehat{Z}_{i}$ the stochastic update.
+Because every $\mathbf{J}_{i}$ is deterministic (wrt.
+to mini-batch),
+$\operatorname{Var}\!\bigl{(}\widehat{g}_{\text{\tiny HM}}\bigr{)}=\sum_{i,j}\mathbf{J}_{i}\;\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})\;\mathbf{J}_{j}^{\!\top},$ (12)
+(derivation in Appendix F.2).
+Equation (12) shows that HyperMARL first averages noise within each agent ($\widehat{Z}_{i}$) and only then applies $\mathbf{J}_{i}$.
+FuPS+ID, by contrast, updates the shared weights $\theta$ with every raw sample $A\,\nabla_{\theta}\log\pi_{\theta}[h,\mathrm{id}]$, leaving observation noise and agent ID entangled and making it susceptible to gradient interference [10, 49].
+F.3 Scalability and Parameter Efficiency
 Hypernetworks generate weights for the target network, which can lead to high-dimensional outputs and many parameters for deep target networks.
@@
   • MLP hypernetworks: Scale better with larger populations, since they only require embeddings of fixed size for each new agent.
-To reduce parameter count, techniques like shared hypernetworks, chunked hypernetworks (von Oswald et al., 2020; Chauhan et al., 2024), or producing low-rank weight approximations, can be used.
+To reduce parameter count, techniques like shared hypernetworks, chunked hypernetworks [42, 9], or producing low-rank weight approximations, can be used.
 While naive implementations are parameter-intensive, this might be less critical in RL and MARL which commonly have smaller actor-critic networks.
@@
 Despite generating 10x smaller networks, HyperMARL consistently outperforms FuPS variants, showing its advantages extend beyond parameter count.
-Appendix D Experiment Details
-D.1 Environments
+F.4 Speed and Memory Usage
+We examine the computational efficiency of HyperMARL compared to NoPS and FuPS by measuring inference speed (Figure 16(a)) and GPU memory usage (Figure 16(b)) as we scale the number of agents.
+The benchmarks were conducted using JAX on a single NVIDIA GPU (T4) with a recurrent (GRU-based) policy architecture.
+All experiments used fixed network sizes (64-dimensional embeddings and hidden layers) with a batch size of 128 and 64 parallel environments, allowing us to isolate the effects of varying agent count.
+Each measurement represents the average of 100 forward passes per configuration, with operations repeated across 10 independent trials.
+(a) Forward pass inference time
+(b) GPU memory usage
+Figure 16: Computational efficiency scaling with number of agents.
+HyperMARL offers a balance between NoPS and FuPS.
+Notably, in real-world deployments, NoPS incurs additional data transfer and synchronisation costs not reflected here, further widening the efficiency gap.
+The results demonstrate that HyperMARL offers a balance between the extremes represented by NoPS and FuPS.
+In practice, NoPS incurs additional data transfer and update costs, widening the efficiency gap.
+Appendix G Experiment Details
+G.1 Environments
 (a) Dispersion.
@@
 Figure 17: Multi-Agent environments used in our experiments.
-Dispersion (VMAS) (Bettini et al., 2022): A 2D environment where four agents collect unique food particles.
-This task requires specialised heterogeneous behaviours and resembles the Specialisation Game from Section 3.1.
-Navigation (VMAS) (Bettini et al., 2022): Agents navigate in a 2D space to assigned goals, receiving dense rewards based on proximity.
+Dispersion (VMAS) [5]: A 2D environment where four agents collect unique food particles. This task requires specialised heterogeneous behaviours and resembles the Specialisation Game from Section E.1.
+Navigation (VMAS) [5]: Agents navigate in a 2D space to assigned goals, receiving dense rewards based on proximity.
 Configurations include shared goals (homogeneous), unique goals (heterogeneous), and mixed goals, where some agents share goals while others have unique ones.
-Multi-Agent MuJoCo (MAMuJoCo) (Peng et al., 2021): A multi-agent extension of MuJoCo, where robot body parts (e.g., a cheetah’s legs) are modelled as different agents.
-Agents coordinate to perform efficient motion, requiring heterogeneous policies (Zhong et al., 2024).
-SMAX (JaxMARL) (Rutherford et al., 2024): Discrete action tasks with 2 to 20 agents on SMACv1- and SMACv2-style maps.
-FuPS baselines have been shown optimal for these settings (Yu et al., 2022; Fu et al., 2022) indicating homogeneous behaviour is preferred here.
-Table 7: Overview of Baseline Methods and Justifications.
-Selected baselines (✓) are chosen based on their relevance to our focus on parameter sharing and specialisation/generalisation in MARL.
-Non-selected methods (✗) are excluded due to implementation limitations or misalignment with our scope.
-Baseline Method Category Chosen Justification
-IPPO (NoPS, FuPS) (De Witt et al., 2020) NoPS/FuPS Serves as a strong MARL baseline for both independent (NoPS) and shared (FuPS) policies.
-MAPPO (NoPS, FuPS) (Yu et al., 2022) NoPS/FuPS Provides a robust MARL baseline with centralized critics for both NoPS and FuPS configurations.
-DiCo (Bettini et al., 2024) Architectural Enables comparison with a method that employs preset diversity levels and maintains both non-shared and shared parameters.
-HAPPO (Zhong et al., 2024) Sequential Updates Facilitates comparison with a method designed for heterogeneous behaviours, utilizing sequential updates and maintaining non-shared parameters per agent.
-SEAC (Christianos et al., 2020) Shared Experience Focuses on experience sharing rather than parameter sharing, which is outside our scope.
-SePS (Christianos et al., 2021) SePS Requires pretraining, which is beyond the scope of our current study.
-CDAS (Li et al., 2021) Intrinsic Reward Only implemented for off-policy methods which are outside our scope, and has been shown to underperform FuPS/NoPS (Fu et al., 2022).
-ROMA/RODE (Wang et al., 2020b, a) Role-based Demonstrates limited practical performance (Christianos et al., 2020).
-SNP-PS (Kim & Sung, 2023) Architectural/Pruning No available code implementation.
-Kaleidoscope (Li et al., 2024) Architectural/Pruning Only implemented for off-policy methods, which is outside our scope.
-D.1.1 Training and Evaluation
+Multi-Agent MuJoCo (MAMuJoCo) [34]: A multi-agent extension of MuJoCo, where robot body parts (e.g., a cheetah’s legs) are modelled as different agents.
+Agents coordinate to perform efficient motion, requiring heterogeneous policies [49].
+SMAX (JaxMARL) [36]: Discrete action tasks with 2 to 20 agents on SMACv1- and SMACv2-style maps.
+FuPS baselines have been shown optimal for these settings [48, 16] indicating homogeneous behaviour is preferred here.
+Table 4: Baseline Methods Selection and Justification.
+Selected methods (✓) were chosen based on their relevance to parameter sharing and specialisation/generalisation in MARL, while excluded methods (✗) did not align with our research objectives or had implementation constraints.
+Our experimental design systematically addresses key questions on agent specialisation and homogeneity, therefore we selected baselines with demonstrated strong performance in their respective settings, ensuring fair and rigorous comparison.
+Method Category Selected Justification & Experimental Settings
+IPPO [12]
+(NoPS, FuPS+ID) NoPS/FuPS Established MARL baseline implementing both independent (NoPS) and fully shared (FuPS+ID) policy configurations.
+Tested in: Dispersion, Navigation, SMAX (two SMACv1 maps and two SMACv2 maps, with 10 and 20 agents).
+MAPPO [48]
+(NoPS, FuPS+ID) NoPS/FuPS Strong baseline with centralized critics for both NoPS and FuPS+ID architectures.
+Tested in: Dispersion, MAMuJoCo, SMAX (two SMACv1 maps and two SMACv2 maps, with 10 and 20 agents).
+DiCo [6] Architectural Diversity Provides comparison with a method employing preset diversity levels that balances shared and non-shared parameters.
+Tested in: Dispersion and Navigation (as per original paper).
+Original hyperparameters used for $n=2$ agents; parameter sweeps conducted for $n>2$ to identify optimal diversity levels.
+HAPPO [49] Sequential Updates Enables comparison with a method designed for heterogeneous behaviours using sequential policy updates with agent-specific parameters.
+Tested in: MAMuJoCo, selecting 4/6 scenarios from the original paper, including the challenging 17-agent humanoid task.
+Walker and Hopper variants were excluded as MAPPO and HAPPO performed similarly in these environments.
+Kaleidoscope [26] Architectural Pruning Implemented for off-policy evaluation using its MATD3 implementation with tuned MaMuJoCo hyperparameters.
+Tested in: MAMuJoCo environments Ant-v2, HalfCheetah-v2, Walker2d-v2 (overlapping with our IPPO experiments), and Swimmer-v2-10x2 (highest agent count variant).
+Included to evaluate HyperMARL’s competitiveness against a method with ensemble critics and diversity loss, in an off-policy setting.
+SEAC [11] Shared Experience Focuses primarily on experience sharing rather than parameter sharing architecture, falling outside our research scope.
+SePS [10] Selective Parameter Sharing Requires pretraining phase, which extends beyond the scope of our current study focused on end-to-end learning approaches.
+CDAS [25] Intrinsic Reward Only implemented for off-policy methods and has been demonstrated to underperform FuPS/NoPS architectures [16], making it less suitable for our primary on-policy comparisons.
+ROMA/RODE [43, 44] Role-based Shows limited practical performance advantages in comparative studies [11], suggesting other baselines provide more rigorous comparison points.
+SNP-PS [23] Architectural Pruning No publicly available implementation, preventing direct, reproducible comparison.
+G.2 HyperMARL Architecture Details
+For the Dispersion and Navigations results (Sec.
+5.2) we use feedforward architectures, where we use HyperMARL to generate both the actor and critic networks.
+For the MAPPO experiments in Section 5.2, for fairness in comparisons with HAPPO and MAPPO, we maintain the centralised critic conditioned on the global state and only use HyperMARL to generate the weights of the actors.
+For the recurrent IPPO experiments in Section 5.3, HyperMARL only generates the actor and critic feedforward weights, not the GRU weights.
+G.2.1 Training and Evaluation
   • Training:
-    • For Dispersion (5.2), we run 5 seeds and train for 20 million timesteps.
+    • For Dispersion (5.2), we run 10 seeds and train for 20 million timesteps.
     • For Navigation (5.2), SMAX (5.3), and MaMuJoCo (5.2), we run 5 seeds and train for 10 million timesteps, consistent with the baselines.
@@
     • For MaMuJoCo (5.2), following the baselines, evaluation is performed every 25 training episodes over 40 episodes.
-D.1.2 Measuring Policy Diversity Details
-We measure team diversity using the System Neural Diversity (SND) metric (Equation 9 (Bettini et al., 2023), details Section A) with Jensen-Shannon distance.
+G.2.2 Measuring Policy Diversity Details
+We measure team diversity using the System Neural Diversity (SND) metric (Equation 5 [7], details Section D) with Jensen-Shannon distance.
 SND ranges from 0 (identical policies across all agents) to 1 (maximum diversity).
@@
 We then sample 1 million observations from this dataset to calculate the SND for each method tested.
-D.2 HyperMARL Architecture
-For the experiments in Sections 5.2 and 5.2, we use feedforward architectures, while for Section 5.3, we use recurrent GRU networks (Cho, 2014).
-For recurrent PPO variants, HyperMARL maintains the same architecture as for feedforward methods, generating only the actor and critic feedforward weights, not the GRU weights.
-Appendix E Detailed Results
-E.1 Dispersion Detailed Results
-E.1.1 Sample Efficiency Plots
-(a) IPPO Sample Efficiency
-(b) MAPPO Sample Efficiency
-Figure 13: Sample efficiency of IPPO and MAPPO variants on the Dispersion environment.
-While FuPS methods struggle to converge to optimal performance, Hypernetworks achieve efficiency comparable to NoPS.
-Even when running MAPPO-FuPS for longer (see Figure 14), it converges to suboptimal performance.
+Appendix H Detailed Results
+H.1 Kaleidoscope Off-Policy Comparison Details
+Our comparison with Kaleidoscope [26], mentioned in Section 5, is conducted using off-policy methods due to its original design.
+Kaleidoscope incorporates intricate mechanisms (e.g., learnable masks, an ensemble of five critics, a specific diversity loss) and numerous specialised hyperparameters (e.g., for critic ensembling: ‘critic_deque_len‘, ‘critic_div_coef‘, ‘reset_interval‘; for mask and threshold parameters: ‘n_masks‘, ‘threshold_init_scale‘, ‘threshold_init_bias‘, ‘weighted_masks‘, ‘sparsity_layer_weights‘, etc.).
+Porting this full complexity to an on-policy PPO backbone would constitute a significant research deviation rather than a direct benchmark of the established method.
+Therefore, we utilised Kaleidoscope’s original off-policy implementation to ensure a meaningful comparison.
+We adopted MATD3 as the algorithmic backbone for this evaluation, as it was the only publicly available Kaleidoscope variant with tuned hyperparameters for Multi-Agent MuJoCo (MaMuJoCo).
+The MaMuJoCo tasks were chosen for alignment with our primary on-policy (IPPO) results and Kaleidoscope’s original evaluation, specifically: Ant-v2, HalfCheetah-v2, Walker2d-v2 (overlapping with our IPPO experiments), and Swimmer-v2-10x2 ( which represents the MaMuJoCo variant with the highest number of agents).
+Comparative results in Table 5 show that HyperMARL achieves competitive results with Kaleidoscope, while only using two critics (standard MATD3) and without additional diversity objectives.
+Table 5: Mean episode return in MAMuJoCo for off-policy MATD3 variants.
+IQM of the mean episode returns with 95% stratified bootstrap CI.
+Bold indicates the highest IQM score; * indicates scores whose confidence intervals overlap with the highest.
+Although Kaleidoscope employs an ensemble of five critics and an explicit diversity loss, HyperMARL (using a standard MATD3 setup with two critics) achieved competitive results without these additional mechanisms.
+Environment Ind. Actors (MATD3) HyperMARL (MATD3) Kaleidoscope (MATD3)
+Ant-v2 5270.38 (4329.73, 5719.78) *5886.58 (5840.00, 5920.66) 6160.70 (5798.02, 6463.83)
+HalfCheetah-v2 *6777.04 (3169.11, 8233.94) 7057.44 (3508.70, 8818.11) *6901.00 (3609.73, 8192.38)
+Walker2d-v2 *5771.87 (5144.84, 8103.34) 7057.68 (5976.50, 8166.09) *6664.32 (5408.95, 8828.11)
+Swimmer-v2-10x2 *453.74 (427.24, 487.86) 465.91 (410.82, 475.77) *462.48 (444.22, 475.64)
+H.2 Dispersion Detailed Results
+H.2.1 Interval Esimates Dispersion
+(a) IPPO
+(b) MAPPO
+Figure 18: Performance of IPPO and MAPPO on Dispersion after 20 million timesteps.
+We show the Interquartile Mean (IQM) of the Mean Episode Return and the 95% Stratified Bootstrap Confidence Intervals (CI) using [1].
+Hypernetworks achieve comparable performance to NoPS, while FuPS struggle with specialisation.
 Figure 19: We see that even if we run MAPPO-FuPS on Dispersion for 40 million timesteps (double the timesteps of MLP Hypernetwork), it converges to suboptimal performance.
-E.2 Detailed MAMujoco Plots
+H.3 Detailed MAMujoco Plots
 (a) Humanoid-v2 17x1
@@
 (d) Ant-v2 4x2
-Figure 15: Performance of Recurrent IPPO and MAPPO on SMAX. HyperMARL performs comparably to these baselines.
-E.3 Detailed Navigation Plots
-(a) Same Goal (8 agents)
-(b) Different Goals (8 agents)
-(c) Four Goals (8 agents)
-(d) Scenarios with 8 Agents
-(e) Same Goal (4 agents)
-(f) Different Goals (4 agents)
-(g) Two Goals (4 agents)
-(h) Scenarios with 4 Agents
-(i) Same Goal (2 agents)
-(j) Different Goals (2 agents)
-(k) Scenarios with 2 Agents
-Figure 17: Sample Efficiency of IPPO variants in the VMAS Navigation environment under different goal configurations and agent numbers.
-E.4 Additional Ablations
+Figure 20: Performance of Recurrent IPPO and MAPPO on MaMoJoCo.
+HyperMARL performs comparably to these baselines, and is the only method with shared actors to demonstrate stable learning in the notoriously difficult 17-agent Humanoid environment.
+H.4 Detailed Navigation Plots
+(a) 8 Agents, Same Goal
+(b) 8 Agents, Different Goals
+(c) 8 Agents, Four Goals
+(d) 4 Agents, Same Goal
+(e) 4 Agents, Different Goals
+(f) 4 Agents, Two Goals
+(g) 2 Agents, Same Goal
+(h) 2 Agents, Different Goals
+Figure 21: Sample efficiency of IPPO variants in the VMAS Navigation environment.
+Plots show IQM and 95% CI (shaded regions) of mean episode return against training steps for different agent counts (rows: 8, 4, 2 agents) and goal configurations (columns, where applicable: Same, Different, Specific Goal Counts).
+Legend shown at bottom applies to all subplots.
+H.5 Interval Esimates - SMAX
+(a) 2s3z
+(b) 3s5z
+(c) SMACv2 10 Units
+(d) SMACv2 20 Units
+Figure 22: Performance comparison in SMAX environments after 10 million timesteps.
+We show the Interquartile Mean (IQM) of the Mean Win Rate and the 95% Stratified Bootstrap Confidence Intervals (CI).
+HyperMARL performs comparably to FuPS baselines across all environments, demonstrating its effectiveness in tasks requiring homogeneous behaviours and using recurrency.
+H.6 Additional Ablations
 Figure 23: Ablation results comparing HyperMARL with its variants in Dispersion. The results highlight that gradient decoupling is essential for maintaining HyperMARL’s performance.
-Appendix F Hyperparameters
-Table 8: Hyperparameters for Specialisation and Synchronisation Game
+Appendix I Hyperparameters
+Table 6: Hyperparameters, Training and Evaluation for Specialisation and Synchronisation Game
 Hyperparameter Value
-Number of agents and foods 2, 4, 8
-Number of foods 3
-Number of seeds 5
-Number of trials 10,000
-Evaluation trials 1,000
-Batch size 10
-Learning rate 0.01
+Environment Configuration
+Number of agents 2, 4, 8, 16
+Maximum steps per episode 10
+Training Protocol
+Number of seeds 10
+Training steps 10,000
+Evaluation episodes 100
+Evaluation interval 1,000 steps
+Batch size 32
+Model Architecture
+Hidden layer size 8, 16, 32, 64
 Activation function ReLU
 Output activation Softmax
+Optimization
+Learning rate 0.01
 Optimizer SGD
-Total Number of Parameters
+Model Parameter Count
 2 Agents
-NoPS: 36
-FuPS: 34
-FuPS+ID: 42
+NoPS: 60
+FuPS: 58
+FuPS+ID: 74
+FuPS+ID (No State): 42
 4 Agents
-NoPS: 112
-FuPS: 100
-FuPS+ID: 148
+NoPS: 352
+FuPS: 240
+FuPS+ID: 404
+FuPS+ID (No State): 148
 8 Agents
-NoPS: 384
-FuPS: 328
-FuPS+ID: 552
+NoPS: 2400
+FuPS: 2344
+FuPS+ID: 2600
+FuPS+ID (No State): 552
+16 Agents
+NoPS: 17728
+FuPS: 17488
+FuPS+ID: 18512
+FuPS+ID (No State): 2128
 Table 7: IPPO and MAPPO Hyperparameters in Dispersion
@@
 continuous_actions True
-Table 17: Default algorithm and model hyperparameters for the Ant-v2-4x2 environment (from  (Zhong et al., 2024)).
+Table 15: Default algorithm and model hyperparameters for the Ant-v2-4x2 environment (from  [49]).
 Parameter Value
@@
 weight_decay 0
-Table 18: Default algorithm and model hyperparameters for the Humanoid-v2-17x1 environment (from  (Zhong et al., 2024)).
+Table 16: Default algorithm and model hyperparameters for the Humanoid-v2-17x1 environment (from  [49]).
 Parameter Value
@@
 weight_decay 0
-Table 19: Default algorithm and model hyperparameters for the Walker2d-v2-2x3 environment (from  (Zhong et al., 2024)).
+Table 17: Default algorithm and model hyperparameters for the Walker2d-v2-2x3 environment (from  [49]).
 Parameter Value
@@
 weight_decay 0
-Table 20: Default algorithm and model hyperparameters for the HalfCheetah-v2-2x3 environment (from  (Zhong et al., 2024)).
+Table 18: Default algorithm and model hyperparameters for the HalfCheetah-v2-2x3 environment (from  [49]).
 Parameter Value
@@
 Table 19: HyperMARL Hyperparameters Across MaMuJoCo Environments
-Parameter Humanoid-v2-17x1 Walker2d-v2-2x3 HalfCheetah-v2-2x3 Ant-v2-4x2 Sweeps
-AGENT_ID_EMBEDDING_DIM 64 64 64 64 None
-HYPERNET_HIDDEN_DIMS 64 64 64 64 None
+Parameter Humanoid Walker2d HalfCheetah Ant Sweeps
+v2-17x1 v2-2x3 v2-2x3 v2-4x2
+AGENT_ID_DIM 64 64 64 64 None
+HNET_HIDDEN_DIMS 64 64 64 64 None
 clip_param 0.075 0.0375 0.0375 0.075 [0.1,0.075,0.05,0.0375]
@@
 Map Algorithm LR Range Chosen LR HNET Embedding Dim HNET Hidden Dims
-2s3z IPPO 0.004 0.004 -
-MAPPO 0.002 0.002 - -
+2s3z IPPO 0.004 0.004
+MAPPO 0.002 0.002
 H-IPPO 0.004 0.004 4 32
 H-MAPPO 0.002 0.002 64 16
-3s5z IPPO 0.004 0.004 - -
-MAPPO 0.002,0.005,0.0003 0.002 - -
+3s5z IPPO 0.004 0.004
+MAPPO 0.002, 0.005, 0.0003 0.002
 H-IPPO 0.004 0.004 64 16
 H-MAPPO 0.002, 0.005, 0.0003 0.0003 64 16
-smacv2_10_units IPPO 0.005,0.001,0.0003,0.004 0.001 - -
-MAPPO 0.002,0.005,0.0003 0.0003 - -
+smacv2_10_units IPPO 0.005, 0.001, 0.0003, 0.004 0.001
+MAPPO 0.002, 0.005, 0.0003 0.0003
 H-IPPO 0.005, 0.001, 0.0003, 0.004 0.005 4 64
 H-MAPPO 0.002, 0.005, 0.0003 0.0003 64 16
-smacv2_20_units IPPO 0.002,0.005,0.0003 0.005 - -
-MAPPO 0.002,0.005,0.0003 0.0003 - -
+smacv2_20_units IPPO 0.002, 0.005, 0.0003 0.005
+MAPPO 0.002, 0.005, 0.0003 0.0003
 H-IPPO 0.002, 0.005, 0.0003 0.005 64 64
@@
 HNET Hidden Dims refers to the hidden layer dimensions of the hypernetwork, chosen from the range {16, 32, 64}.
+Appendix J Computational Resources
+Table 22: Computational Resources by Experiment Type
+Experiment Category Hardware Configuration Execution Time Total Hours
+Specialisation, 8 cores on AMD EPYC 7H12 2-6 hours per run 250
+Synchronisation & Dispersion 64-Core Processor (agent-count dependent)
+Navigation 8 cores on AMD EPYC 7H12 4-10 hours per run 320
+Experiments 64-Core Processor + NVIDIA RTX A4500
+MaMuJoCo 8 cores on AMD EPYC 7H12 8-24 hours per run 1,680
+Experiments 64-Core Processor + NVIDIA RTX A4500 (scenario & algorithm dependent)
+SMAX 8 cores on AMD EPYC 7H12 2-5 hours per run 160
+Experiments 64-Core Processor + NVIDIA RTX A4500

### v4

@@
 Abstract
-Adaptability to specialised or homogeneous behaviours is critical in cooperative multi-agent reinforcement learning (MARL).
-Parameter sharing (PS) techniques, common for efficient adaptation, often limit behavioural diversity due to cross-agent gradient interference, which we show can be exacerbated by the coupling of observations and agent IDs.
-Current remedies typically add complexity through altered objectives, manual preset diversity levels, or sequential updates.
-We ask: can shared policies adapt without these complexities?
-We propose HyperMARL, a PS approach using hypernetworks for dynamic agent-specific parameters, without altering the RL objective or requiring preset diversity levels.
-HyperMARL’s explicit decoupling of observation- and agent-conditioned gradients empirically reduces policy gradient variance, facilitates shared-policy adaptation (including specialisation), and helps mitigate cross-agent interference.
-Across diverse MARL benchmarks (up to 20 agents), requiring homogeneous, heterogeneous, or mixed behaviours, HyperMARL achieves competitive performance against key baselines – fully shared, non-parameter sharing, and three diversity-promoting methods – while preserving behavioural diversity comparable to non-parameter sharing.
-These findings establish HyperMARL as a versatile approach for adaptive MARL.
+Adaptive cooperation in multi-agent reinforcement learning (MARL) requires policies to express homogeneous, specialised, or mixed behaviours, yet achieving this adaptivity remains a critical challenge.
+While parameter sharing (PS) is standard for efficient learning, it notoriously suppresses the behavioural diversity required for specialisation.
+This failure is largely due to cross-agent gradient interference, a problem we find is surprisingly exacerbated by the common practice of coupling agent IDs with observations.
+Existing remedies typically add complexity through altered objectives, manual preset diversity levels, or sequential updates – raising a fundamental question: can shared policies adapt without these intricacies?
+We propose a solution built on a key insight: an agent-conditioned hypernetwork can generate agent-specific parameters and decouple observation- and agent-conditioned gradients, directly countering the interference from coupling agent IDs with observations.
+Our resulting method, HyperMARL, avoids the complexities of prior work and empirically reduces policy gradient variance.
+Across diverse MARL benchmarks (22 scenarios, up to 30 agents), HyperMARL achieves performance competitive with six key baselines while preserving behavioural diversity comparable to non-parameter sharing methods, establishing it as a versatile and principled approach for adaptive MARL.
 The code is publicly available at https://github.com/KaleabTessera/HyperMARL.
@@
 Specialist and generalist behaviours are critical to collective intelligence, enhancing performance and adaptability in both natural and artificial systems [52, 43, 44, 23, 50].
-In Multi-Agent Reinforcement Learning (MARL), this translates to a critical need for policies that can flexibly exhibit specialised, homogeneous, or mixed behaviours to meet diverse task demands [25, 6, 2].
+In Multi-Agent Reinforcement Learning (MARL) [2], this translates to a critical need for policies that can flexibly exhibit specialised, homogeneous, or mixed behaviours to meet diverse task demands [26, 6].
 Figure 1: HyperMARL Policy Architecture.
-Common Agent-ID conditioned shared MARL policy (left) vs HyperMARL (right), which uses a hypernetwork to generate agent-specific policies and decouples observation- and agent-conditioned gradients.
+Common agent-ID conditioned shared MARL policy (FuPS+ID, left) vs HyperMARL (right), which uses an agent-conditioned hypernetwork to generate agent-specific weights and decouples observation- and agent-conditioned gradients.
 Optimal MARL performance thus hinges on being able to represent the required behaviours.
@@
 Conversely, Full Parameter Sharing (FuPS) [45, 18, 15], which trains a single shared network, improves efficiency but typically struggles to foster the behavioural diversity necessary for many complex tasks [24, 17, 26].
-Balancing FuPS efficiency with the capacity for diverse behaviours therefore remains a central open problem in MARL.
-Prior works have explored intrinsic-rewards [25, 21], role-based allocations [43, 44], specialised architectures [23, 26, 6], or sequential updates [49].
-However, these methods often alter the learning objective, require prior knowledge of optimal diversity levels, involve delicate hyperparameter tuning, or necessitate maintaining agent-specific parameters and sequential updates (see Section 7).
-These challenges lead us to ask: Can we design a shared MARL architecture that flexibly supports both specialised and homogeneous behaviours—without altered learning objectives, manual preset diversity levels, or sequential updates?
-A key difficulty with FuPS, particularly for diverse behaviours, was hypothesised to be gradient interference among agents, whereby their updates negatively impact each other’s learning [10, 49].
+This failure of FuPS, particularly for diverse behaviours, was hypothesised to be gradient interference among agents, whereby their updates negatively impact each other’s learning [11, 54].
 We not only empirically validate this hypothesis but also demonstrate a critical insight: this conflict is significantly exacerbated by the common practice of coupling observations with agent IDs within a shared network (Fig.
@@
 3.2 for results).
-To counteract the challenges posed by this coupling, we propose HyperMARL, a novel parameter sharing method that explicitly decouples observation- and agent-conditioned gradients (Section 4.3) through hypernetworks [18].
-Specifically, agent-conditioned hypernetworks, which condition on an agent identifier (or embedding), and generate an agent’s actor and critic weights on the fly (Fig.
-1).
-Hypernetworks have already proved effective for resolving gradient conflicts in multi-task RL [30] and for mitigating catastrophic forgetting in continual learning [42], our work establishes their effectiveness in MARL.
-We further show that HyperMARL empirically attains lower policy gradient variance than FuPS and that this decoupling is critical for specialisation (Sec.
-5.26), suggesting it helps mitigate cross-agent gradient interference in shared architectures.
-We validate HyperMARL on diverse MARL benchmarks—including Dispersion and Navigation (VMAS) [5], Multi-Agent MuJoCo (MAMuJoCo) [34], and SMAX [36]—across environments with two to twenty agents that require homogeneous, heterogeneous, or mixed behaviours.
-HyperMARL consistently matches or outperforms NoPS, FuPS, and diversity-promoting methods such as DiCo [6],HAPPO [49] and Kaleidoscope [26], while achieving NoPS-level behavioural diversity while using a shared architecture.
+Balancing FuPS efficiency with the capacity for diverse behaviours therefore remains a central open problem in MARL.
+Prior works have explored intrinsic-rewards [26, 22], role-based allocations [48, 49], specialised architectures [24, 28, 6], sequential updates [54], and sharing parameters within clusters of agents [11].
+However, these remedies introduce their own intricacies: they often alter the learning objective, require prior knowledge of optimal diversity levels, necessitate maintaining agent-specific parameters or require sequential updates.
+This raises a fundamental question: Can we design a shared MARL architecture that flexibly supports both specialised and homogeneous behaviours—without altered learning objectives, manual preset diversity levels, or sequential updates?
+Guided by our observation-ID coupling insight, we propose HyperMARL, a novel agent-conditioned hypernetwork [19] architecture.
+HyperMARL generates per-agent weights on the fly (Fig.
+1) and explicitly decouples observation- and agent-conditioned gradients (Sec.
+4.3).
+This choice is motivated by hypernetworks’ proven effectiveness at resolving gradient conflicts in multi-task RL [33] and continual learning [47].
+Our work establishes their effectiveness for the problem of cross-agent interference in MARL.
+Indeed, HyperMARL empirically attains lower policy gradient variance than FuPS, and we show this decoupling is critical for specialisation (Sec.
+5.26.1), confirming its role in mitigating interference.
+We validate HyperMARL on diverse MARL benchmarks – including Dispersion and Navigation (VMAS) [5], Multi-Agent MuJoCo (MAMuJoCo) [37], SMAX [41], and Blind-Particle Spread (BPS) [11] – across environments with two to thirty agents that require homogeneous, heterogeneous, or mixed behaviours.
+HyperMARL consistently matches or outperforms NoPS, FuPS, and diversity-promoting methods such as Diversity Control (DiCo) [6], Heterogeneous-Agent Proximal Policy Optimisation (HAPPO) [54], Kaleidoscope [28] and Selective Parameter Sharing (SePS) [11], while achieving NoPS-level behavioural diversity while using a shared architecture.
 Our contributions are as follows:
-  • Motivated by evidence that coupling agent IDs with observations can be detrimental, we propose HyperMARL (Sec.
-4), a method that leverages agent-conditioned hypernetworks to decouple observation- and agent-conditioned gradients.
-HyperMARL adaptively learns diverse, homogeneous, or mixed behaviours without altering the RL objective or requiring preset diversity levels.
-  • We thoroughly evaluate HyperMARL (Section 5) across diverse MARL environments with up to 20 agents.
-HyperMARL consistently achieves competitive mean episode returns compared to NoPS, FuPS, and diversity-focused methods such as DiCo, HAPPO and Kaleidoscope, while achieving NoPS-level behavioural diversity.
-  • We demonstrate that HyperMARL empirically reduces policy gradient variance compared to FuPS and facilitates specialisation, suggesting the importance of gradient decoupling for mitigating cross-agent interference (Section 5.2; Section 6).
+  • We identify that cross-agent gradient interference in shared policies is critically exacerbated by the common practice of coupling agent IDs with observations (Sec. 3.2).
+  • We propose HyperMARL (Sec.
+4), an agent-conditioned hypernetwork architecture, to test the hypothesis that explicitly decoupling these gradients enables adaptive (diverse, homogeneous, or mixed) behaviours without the complexities of prior remedies (e.g., altered objectives, preset diversity levels, or sequential updates).
+  • Our extensive evaluation (Sec.
+5) across 22 diverse scenarios (up to 30 agents) shows HyperMARL achieves competitive returns against six strong baselines, while achieving NoPS-level behavioural diversity.
+We further show this decoupling is empirically linked to reduced policy gradient variance and is critical for specialisation (Sec.
+5.2; Sec.
+6.1).
 2 Background
@@
 These observations are accumulated into an action-observation history $h_{t}^{i}=\left(o_{0}^{i},a_{0}^{i},\ldots,o_{t-1}^{i},a_{t-1}^{i},o_{t}^{i}\right)$.
-Each agent $i$ acts based on their decentralized policies $\pi^{i}(a^{i}|h^{i})$.
+Each agent $i$ acts based on their decentralised policies $\pi^{i}(a^{i}|h^{i})$.
 The joint history and joint action are defined as follows $\mathbf{h_{t}}=(h_{t}^{1},\ldots,h_{t}^{n})$ and $\mathbf{a_{t}}=(a_{t}^{1},\ldots,a_{t}^{n})$.
-The goal is to learn an optimal joint policy $\boldsymbol{\pi^{*}}=(\pi^{1*},...,\pi^{n*})$ that maximizes the expected discounted return as follows, [^1] We use simplified notion here omitting explicit dependence on state transitions and distributions for brevity.
+The goal is to learn an optimal joint policy $\boldsymbol{\pi^{*}}=(\pi^{1*},...,\pi^{n*})$ that maximizes the expected discounted return as follows, [^1] We use simplified notation here omitting explicit dependence on state transitions and distributions for brevity.
 $\boldsymbol{\pi}^{*}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{s_{0}\sim\rho_{0},\;\mathbf{h}\sim\boldsymbol{\pi}}\left[G(\mathbf{h})\right]$, where $G(\mathbf{h})=\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})$.
@@
 NoPS allows for uninhibited agent specialisation but can be sample inefficient and computationally expensive.
-FuPS, often conditioned with an agent ID (FuPS+ID), is more efficient but can struggle when agents must learn diverse behaviours.
+FuPS, often conditioned with an agent ID (FuPS+ID), is more efficient but can struggle when agents must learn diverse behaviours [11].
 This section investigates the limitations of these common policy architectures.
@@
 B (1,1) (0.5,0.5)
-Two-Player Payoff matrix
-$N$-player Interaction
+Two‐Player Payoff matrix
+NN‐player Interaction
 (a) Specialisation Game
@@
 B (0.5,0.5) (1,1)
-Two-Player Payoff matrix
-$N$-player Interaction
+Two‐Player Payoff matrix
+NN‐player Interaction
 (b) Synchronisation Game
@@
 The Specialisation game (left), which encourages distinct actions, and the Synchronisation game (right), where rewards encourage identical actions.
-Depicted are their two-player payoff matrices (pure Nash equilibria in blue) and $N$-player interaction schematics.
+Depicted are their two-player payoff matrices (pure Nash equilibria in blue) and NN-player interaction schematics.
 While simple in form, these games are challenging MARL benchmarks due to non-stationarity and exponentially scaling observation spaces (temporal version).
 To probe these limitations, we introduce two illustrative environments: the Specialisation Game, rewarding distinct actions, and the Synchronisation Game, rewarding identical actions.
-Both are inspired by prior work [16, 5, 32] and extended here to $N$-agent and temporal settings where agents observe prior joint actions (see Appendix E for full definitions).
+Both are inspired by prior work [10, 17, 5, 35] and extended here to $N$-agent and temporal settings where agents observe prior joint actions (see Appendix E for full definitions).
 3.1 Limitations of Fully Shared and Independent Policies
-Figure 3: Average evaluation reward (mean $\pm$ 95% CI) for temporal Specialisation vs.
+Table 1: Average evaluation reward (mean $\pm$ 95% CI) for temporal Specialisation vs.
 Synchronisation using REINFORCE (10 seeds).
@@
 #Ag NoPS FuPS FuPS+ID NoPS FuPS FuPS+ID
-2 0.88$\pm$0.09 0.50$\pm$0.00 0.64$\pm$0.10 0.83$\pm$0.12 1.00$\pm$0.00 0.91$\pm$0.09
-4 0.74$\pm$0.08 0.25$\pm$0.00 0.40$\pm$0.07 0.32$\pm$0.03 1.00$\pm$0.00 0.67$\pm$0.15
-8 0.68$\pm$0.02 0.12$\pm$0.00 0.25$\pm$0.03 0.14$\pm$0.00 1.00$\pm$0.00 0.54$\pm$0.10
-16 0.64$\pm$0.01 0.06$\pm$0.00 0.13$\pm$0.02 0.07$\pm$0.00 1.00$\pm$0.00 0.55$\pm$0.14
+[…]
+[…]
+[…]
+[…]
 (a) Avg. evaluation reward
@@
 Figure 3: Multi-agent policy gradient methods in the Specialisation environment.
-The FuPS+ID (No State) ablation outperforms FuPS+ID, showing near-orthogonal gradients (4(b)), indicating that observation–ID decoupling is important.
+The FuPS+ID (No State) ablation outperforms FuPS+ID, showing near-orthogonal gradients (3(b)), indicating that observation–ID decoupling is important.
 HyperMARL (MLP) enables this decoupling while leveraging state information, and achieves better performance and reduced gradient conflict than FuPS+ID.
@@
 We compare three standard architectures trained with REINFORCE [51]: 1) NoPS: independent policies ($\pi_{\theta^{i}}(a^{i}|o^{i})$); 2) FuPS: a single shared policy ($\pi_{\theta}(a^{i}|o^{i})$); and 3) FuPS+ID: a shared policy incorporating a one-hot agent ID ($\pi_{\theta}(a^{i}|o^{i},id^{i})$).
-All use single-layer networks, 10-step episodes, and $10,000$ training steps (further details in App.
-6).
+All use single-layer networks, 10-step episodes and $10,000$ training steps (further details in Table 9 in App.
+I).
 Empirical Performance.
@@
 E.4).
-Importance of Observation and ID Decoupling To investigate the effect of entangled observation and ID inputs, we introduce an ablation: FuPS+ID (No State), where the policy $\pi_{\theta}(a^{i}\mid id^{i})$ conditions only on the agent ID, ignoring observations.
+Importance of Observation and ID Decoupling.
+To investigate the effect of entangled observation and ID inputs, we introduce an ablation: FuPS+ID (No State), where the policy $\pi_{\theta}(a^{i}\mid id^{i})$ conditions only on the agent ID, ignoring observations.
 Surprisingly, FuPS+ID (No State) outperforms standard FuPS+ID in the Specialisation Game for all tested $N$ (Figure 3(a)), even when $N\leq 4$ (where observation spaces are small, suggesting the issue is not merely observation size).
@@
 The parameters $\theta^{i}$ and $\phi^{i}$ define each agent’s policy $\pi_{\theta^{i}}$ and critic $V_{\phi^{i}}$, dynamically enabling either specialised or homogeneous behaviours as required by the task.
-Linear Hypernetworks Given a one-hot agent ID, $\mathds{1}^{i}\in\mathbb{R}^{1\times n}$, a linear hypernetwork $h_{\psi}^{\pi}$ generates agent-specific parameters $\theta^{i}$ as :
+Linear Hypernetworks Given a one-hot agent ID, $\mathds{1}^{i}\in\mathbb{R}^{1\times n}$, a linear hypernetwork $h_{\psi}^{\pi}$ generates agent-specific parameters $\theta^{i}$ as follows [^3] For conciseness we only show the policy parameters in this section.:
 $\theta^{i}=h_{\psi}^{\pi}(\mathds{1}^{i})=\mathds{1}^{i}\cdot W+b$ (2)
@@
 A core difficulty in FuPS is cross-agent gradient interference [11, 54].
-HyperMARL mitigates this by generating each agent’s parameters through a shared hypernetwork, thereby decoupling agent-conditioned and observation-conditioned components of the gradient.
+HyperMARL mitigates this by generating each agent’s parameters through a shared hypernetwork, thereby decoupling agent-conditioned and observation‑conditioned components of the gradient.
 Hypernetwork gradients. Consider a fully cooperative MARL setting with a centralised critic, we can formulate the policy gradient for agent $i$ as follows [2, 25]:
@@
   • Agent-conditioned factor $\mathbf{J}_{i}$.
-This Jacobian depends only on the fixed embedding $e^{i}$ and the hypernetwork weights $\psi$, therefore, it is deterministic with respect to mini-batch samples (as $e^{i}$ and $psi$ are fixed per gradient step), separating agent identity from trajectory noise.
+This Jacobian depends only on the fixed embedding $e^{i}$ and the hypernetwork weights $\psi$, therefore, it is deterministic with respect to mini-batch samples (as $e^{i}$ and $\psi$ are fixed per gradient step), separating agent identity from trajectory noise.
   • Observation-conditioned factor $Z_{i}$.
@@
 F.2).
-This is the MARL analogue of the task/state decomposition studied by [37, Eq.
+This is the MARL analogue of the task/state decomposition studied by [42, Eq.
 18] in Meta-RL.
-Section 5.2 empirically verifies the predicted variance drop, and ablations in Section 6 demonstrate that disabling decoupling degrades performance, underscoring its critical role.
+Section 5.2 empirically verifies the predicted variance drop, and ablations in Section 6.1 demonstrate that disabling decoupling degrades performance, underscoring its critical role.
 5 Experiments
@@
 Q2: Effectiveness in Homogeneous Tasks: Is HyperMARL competitive in environments that necessitate homogeneous behaviours?
-To address these questions, HyperMARL is evaluated against five representative modern baselines across a carefully selected suite of MARL benchmarks (18 scenarios, up to 20 agents).
+To address these questions, HyperMARL is evaluated against six representative modern baselines across a carefully selected suite of MARL benchmarks (22 scenarios, up to 30 agents).
 All experiments use at least 5 seeds (details in App.
@@
 SMAX 2–20 Discrete Homo.
+BPS 15–30 Discrete Hetero.
 Environments.
-HyperMARL is evaluated across 18 scenarios from four diverse MARL environments (Dispersion [5], Navigation [5], MAMuJoCo [34], and SMAX [36]) (Table 1).
-These were specifically chosen to rigorously test performance across varying complexities, agent counts (2 to 20), and distinct behaviours (heterogeneous, homogeneous, or mixed).
+HyperMARL is evaluated across 22 scenarios from five diverse MARL environments (Dispersion [5], Navigation [5], MAMuJoCo [37], SMAX [41], BPS [11]) (Table 2).
+These were specifically chosen to rigorously test performance across varying complexities, agent counts (2 to 30), and distinct behaviours (heterogeneous, homogeneous, or mixed).
 Full details in Appendix G.1.
@@
 Core PS comparisons use FuPS+ID and NoPS.
-For specialisation, we include privileged baselines: DiCo [6] (shared and non-shared weights, preset diversity levels); HAPPO [49] (shared critic, sequential actor updates); and Kaleidoscope [26] (learnable masks, critic ensembles, diversity loss).
-We use IPPO/MAPPO [12, 48] as the underlying algorithm for all methods except Kaleidoscope (MATD3 backbone due to its intricate off-policy design; see App.
-H.1 for these results).
+For specialisation, we include privileged baselines: DiCo [6] (shared and non-shared weights, preset diversity levels); HAPPO [54] (shared critic, sequential actor updates); Kaleidoscope [28] (learnable masks, critic ensembles, diversity loss); and SePS (pre-training phase, agent clustering).
+We use IPPO/MAPPO [13, 53] as the underlying algorithm for all methods except Kaleidoscope and SePS (see App.
+H.1 and App.
+H.2 for these results).
 Adhering to best evaluation practices [36], we  use original codebases/hyperparameters and environments for which baselines were tuned.
@@
 Interval estimates in App.
-H.2.1.
+H.3.1.
 (c) Actor gradient variance: Hypernetworks achieve lower gradient variance than FuPS+ID.
@@
 5.2 Q1: Specialised Policy Learning
-Learning Diverse Behaviour (Dispersion) Figures 5(a) and 5(b) show that FuPS variants (IPPO-FuPS, MAPPO-FuPS – ($\bullet$)) can struggle to learn the diverse policies required by Dispersion (even when running for longer - Fig.
+Learning Diverse Behaviour (Dispersion) Figures 4(a) and 4(b) show that FuPS variants (IPPO-FuPS, MAPPO-FuPS – ($\bullet$)) can struggle to learn the diverse policies required by Dispersion (even when running for longer - Fig.
 19), while their NoPS counterparts (IPPO-NoPS, MAPPO-NoPS–($\bullet$)) converge to the optimal policy, corroborating standard FuPS limitations to learn diverse behaviour.
@@
 SND policy diversity measurements (Fig.
-5(d)) confirm FuPS variants achieve lower behavioural diversity than NoPS, while HyperMARL notably matches NoPS-level diversity.
+4(d)) confirm FuPS variants achieve lower behavioural diversity than NoPS, while HyperMARL notably matches NoPS-level diversity.
 Figure 5: 17-agent Humanoid learning dynamics (IQM, 95% CI).
@@
 HyperMARL (IPPO and MAPPO variants) also exhibits lower mean policy gradient variance than FuPS+ID across actor parameters (Fig.
-5(c)).
+4(c)).
 This aligns with their ability to learn diverse behaviours and supports the hypothesis that its gradient decoupling mechanism (Sec.
@@
 Notably, unlike HAPPO and MAPPO (independent actors), HyperMARL uses a shared actor and parallel updates, and yet manages strong performance, even in the 17-agent Humanoid-v2 notoriously difficult heterogeneous task[54] (Fig.
-6), matching methods that employ independent actors and sequential updates.
+5), matching methods that employ independent actors and sequential updates.
 Table 3: Mean episode return in MAMuJoCo for MAPPO variants(IQM, 95% CI).
@@
 We compare HyperMARL with baselines including DiCo [6].
-While using DiCo’s optimal preset diversity for n=2 agents, we note that identifying appropriate diversity levels for DiCo with larger teams ($n>2$) via hyperparameter sweeps proved challenging (see Tables 12 and 13).
+While using DiCo’s optimal preset diversity for n=2 agents, we note that identifying appropriate diversity levels for DiCo with larger teams ($n>2$) via hyperparameter sweeps proved challenging (see Tables 15 and 16).
 (a) Shared Goals
@@
 Sample-efficiency curves appear in App.
-H.4.
+H.5.
 Across all tested goal configurations (shared, unique, and mixed), HyperMARL consistently achieves strong performance (Figure 6).
@@
 HyperMARL performs comparably to FuPS baselines across all environments, demonstrating its effectiveness in tasks requiring homogeneous behaviours and using recurrency.
-Inverval estimates in App.
-22.
+Interval estimates in Fig.
+22 in App.
+H.6.
 5.3 Q2: Effectiveness in Homogeneous Tasks
@@
 Figure 7 shows while some FuPS variants might exhibit marginally faster initial convergence on simpler maps, HyperMARL achieves comparable final performance on all maps, using the same GRU backbone for partial observability.
-These results highlights two points: (i) HyperMARL is fully compatible with recurrent architectures essential under partial observability, and (ii) it has no intrinsic bias toward specialisation and can converge homogenous behaviour when it is optimal (also shown with strong same-goal Navigation performance (Fig.
-7(a))), even with large observation spaces and many agents.
+These results highlight two points: (i) HyperMARL is fully compatible with recurrent architectures essential under partial observability, and (ii) it has no intrinsic bias toward specialisation and can converge to homogeneous behaviour when it is optimal (also shown with strong same-goal Navigation performance (Fig.
+6(a))), even with large observation spaces and many agents.
 Summary.
@@
 For Q2 (Homogeneity), HyperMARL demonstrated competitive performance against strong FuPS baselines in SMAX and shared-goal Navigation, confirming its versatility.
-6 Ablations
-We ablate two critical components of HyperMARL, gradient decoupling(Sec.
+6 Ablations and Embedding Analysis
+6.1 Ablations: Gradient Decoupling and Initialisation Scaling
+We ablate two critical components of HyperMARL: gradient decoupling (Sec.
 4.3) and initialisation scaling (Sec.
-4.2) on the complex Humanoid-v2 and simpler Dispersion tasks.
-Removing gradient decoupling (HyperMARL w/o GD), forcing joint observation/embedding processing, universally degrades performance (Figure 9).
-Omitting initialisation scaling (HyperMARL w/o RF (reset fan in/out), which aligns initial generated weights with conventional networks, reveals task-dependent effects: it is crucial for the 17-agent Humanoid, echoing prior findings on hypernetwork initialisation [8], but has minor impact in Dispersion.
-Thus, while principled initialisation becomes vital with increasing complexity, gradient decoupling is fundamentally essential across tested scenarios.
+4.2).
+In HyperMARL w/o GD, the hypernetwork is conditioned on $o_{t},e^{i}_{t}]$, coupling observation and agent-ID gradients.
+In HyperMARL w/o RF, we remove the reset fan-in/out scaling that aligns the scale of generated parameters $(\theta^{i},\phi^{i})$ with standard initialisers.
+Gradient decoupling is essential; initialisation scaling grows with complexity.
+Figure 8 shows that removing GD consistently degrades performance across both Humanoid-v2 (17 agents) and Dispersion, showing that GD is an essential component of HyperMARL.
+Removing RF reveals a task-dependent effect: it is critical on Humanoid-v2, consistent with hypernetwork initialisation results [8], but has a minor impact on Dispersion, indicating that principled initialisation becomes more vital with increased complexity.
+We provide additional ablations in App.
+H.7.
 (a) Humanoid w/o GD
@@
 (d) Dispersion w/o RF
-Figure 9: Ablation results comparing HyperMARL to variants without gradient decoupling (w/o GD) and without reset fan in/out initialization (w/o RF) across environments.
+Figure 8: Ablation results comparing HyperMARL to variants without gradient decoupling (w/o GD) and without reset fan in/out initialisation (w/o RF) across environments.
 Gradient decoupling (a,b) is consistently critical across both environments, while initialisation scaling (c,d) shows greater importance in the complex Humanoid task but less impact in the simpler Dispersion environment.
+Figure 9: Embedding similarity reflects task demands.
+Mean pairwise cosine distance (dashed line = 1.0).
+Same goal: contraction ($0.882\pm 0.042$).
+Different goals: near-orthogonal ($1.010\pm 0.017$).
+Embeddings contract when a shared policy is optimal and maintain separation for specialisation.
+6.2 Analysis of Learned Agent Embeddings
+Recall from 4.2 that each agent $i$ possesses an embedding $e^{i}$, which serves as input to the hypernetworks.
+For the MLP hypernetworks, these embeddings are learned end-to-end and are orthogonally initialised.
+Consequently, at initialisation (step 0), the pairwise cosine distance between any two distinct agent embeddings is $1.0$, mirroring the separability of one-hot IDs.
+We probe how these embeddings $e^{i}$ adapt in the 4-agent Navigation task (same task as in Figure H.5) under two objectives with identical dynamics: (i) same goal (all agents navigate to a single shared target) and (ii) different goals (each agent navigates to its own unique target).
+At the end of training, we compute the mean pairwise cosine distance between agents’ embeddings (lower values imply greater similarity) and compare it to the orthogonal baseline of $1.0$.
+(Cosine distance =1cosine similarity=1-$\text${cosine similarity}; 0 = identical direction, $1$ = orthogonal, $2$ = opposite).
+Figure 9 shows the per-seed distributions, with the $1.0$ initialisation value as a dashed line.
+Agent embeddings adapt to task demands.
+When behaviour should be homogeneous (same goal), embedding directions become significantly more aligned, the mean pairwise cosine distance contracts to $0.882\pm 0.042$ (one-sample $t$-test vs $1.0$: $p=0.0079$).
+Conversely, when behaviour must differ (different goals), the embeddings maintain their separation, remaining near their orthogonal initialisation at $1.010\pm 0.017$.
+These results demonstrate that the hypernetwork actively modulates the agent embeddings based on the task, promoting alignment for homogeneity while preserving separability for specialisation.
+We also conduct a sensitivity analysis on HyperMARL’s hyperparameters in App F.5 and see that agent embedding size can be an important hyperparameter, one that could correspond to the task’s diversity requirements.
 7 Related Work
@@
 In MARL, QMIX [40] used hypernetworks (conditioned on a global state) to mix per-agent Q-values; however, each agent’s own network remained a standard GRU.
-Parallel work, CASH [15], conditions hypernetworks on local observations and team capabilities for zero-shot generalization with heterogeneous action spaces.
-In contrast, we focus on agent-conditioned hypernetworks for adaptive specialisation within a fixed state-action setting, leveraging gradient decoupling (absent in CASH) that we find critical for specialised behaviours.
+Parallel work, CASH [16], conditions a hypernetwork on local observations and predefined capability descriptors to target zero-shot generalisation across heterogeneous action spaces.
+By contrast, we use agent-conditioned hypernetworks in homogeneous action spaces, conditioning only on agent IDs/learned embeddings, and we explicitly decouple agent-conditioned from observation-conditioned gradients – a mechanism absent in CASH – which we find critical for specialisation.
 Variants of Parameter Sharing.
-Selective Parameter Sharing (SePS) [10] shares weights between similar groups of agents, identified via trajectory clustering.
-Pruning methods [23, 26] split a single network into agent-specific subnetworks.
+While Full Parameter Sharing (FuPS) is the most common approach, several other variants exist.
+Selective Parameter Sharing (SePS) [11] shares weights between similar groups of agents, identified by clustering agent trajectories during a pre-training phase.
+Pruning methods (SNP-PS, Kaleidoscope) [24, 28] split a single network into agent-specific subnetworks using learned agent masks.
+AdaPS [27] combines clustering and pruning masks to create shared policies for different groups of agents.
+Concurrent to our work, GradPS [38] identifies neurons with conflicting gradient updates, clones these neurons, and assigns each clone to a group of agents with low gradient conflict.
+Unlike these works, HyperMARL does not rely on pre-training (SePS), clustering algorithms requiring a preset number of clusters (SePS, GradPS, AdaPS), intricate pruning hyperparameters (SNP-PS, Kaleidoscope), auxiliary diversity losses (Kaleidoscope), or gradient conflict thresholds (GradPS).
 Learning Diverse Policies.
 Shared parameters often limit policy diversity [11, 24, 17, 26].
-Proposed solutions include: (1) maximizing mutual information between agent IDs and trajectories [25], (2) role-based methods [43, 44], or (3) methods that use structural modifications or constraints to induce diversity in agent policies [23, 6, 26].
+Proposed solutions include: maximising mutual information between agent IDs and trajectories [26], role-based methods that assign distinct roles to agents [48, 49], best-response [39] methods and approaches that use structural modifications or constraints to induce diversity [24, 6, 28, 27].
 Outside FuPS/NoPS, HAPPO [54] uses a non-shared centralised critic with individual actors updated sequentially to learn heterogeneous behaviours.
+In contrast to these works, HyperMARL does not alter the learning objective, use sequential updates or require preset diversity levels.
 8 Conclusion
-We introduced HyperMARL, an approach that uses agent-conditioned hypernetworks to generate per-agent parameters without modifying the standard RL objective or requiring preset diversity levels.
-Our results show HyperMARL can adaptively learn specialised, homogeneous, or mixed behaviours in settings with up to 20 agents.
-We also observe a link between HyperMARL’s performance and reduced policy gradient variance, underscoring the importance of decoupling observation- and agent-conditioned gradients.
-Overall, these findings establish HyperMARL as a promising architecture for diverse MARL tasks.
+We investigated why standard parameter sharing fails at behavioural diversity, identifying that cross-agent gradient interference is critically exacerbated by coupling agent IDs with observations.
+We hypothesised that explicitly decoupling these gradients would enable adaptivity without prior complexities, and confirmed this using our HyperMARL approach.
+Our results show this decoupling enables adaptive behaviours (up to 30 agents) and is linked to reduced policy gradient variance.
+These findings establish gradient decoupling via HyperMARL as a versatile, principled approach for adaptive MARL.
 We discuss limitations in App.
 A, most notably parameter count, which can be remedied by parameter-efficient hypernetworks (e.g., chunked variants [47, 9]).
+9 Acknowledgements
+We would like to thank Samuel Garcin, Max Tamborski, Dave Abel, Timothy Hospedales, Trevor Mcinroe, Elliot Fosong, and Aris Filos-Ratsikas for fruitful discussions on early versions of this work.
+We also wish to acknowledge the anonymous reviewers for their constructive comments and feedback that helped strengthen this work.
 Appendix A Limitations
@@
 On the other hand, the Jensen-Shannon distance (JSD) [14], which is the square root of the Jensen-Shannon divergence, does not have this problem as it is a smooth distance function.
-Furthermore, it satisfies the conditions for being a metric – it is non-negative, symmetry, and it satisfies the triangle inequality.
+Furthermore, it satisfies the conditions for being a metric – it is non-negative, symmetric, and it satisfies the triangle inequality.
 For continuous policies, as shown in Figure 11, JSD exhibits similar trends to the Wasserstein distance.
@@
 Method Kinds of Actions Metric Smooth Formula
-[…]
+Wasserstein Distance [46] Continuous* Metric Yes $W(p,q)=\left(\inf_{\gamma\in\Gamma(p,q)}\int_{\mathbb{R}\times\mathbb{R}}|x-y|\,d\gamma(x,y)\right)^{1/p}$
 Total Variation Distance Discrete Metric No $TV(p,q)=\frac{1}{2}\sum_{x}|p(x)-q(x)|$
@@
 Jensen-Shannon Distance [14] Both Metric Yes $\sqrt{JSD(p\parallel q)}$
-Table 3: Measure Policy Diversity
+Table 4: Measures for Policy Diversity
 Appendix E Specialisation and Synchronisation Games
@@
 This structure satisfies Definition 1, since optimal joint policies require complementary, non-identical strategies.
-There is also a symmetric mixed-strategy equilibrium in which each agent plays $A$ and $B$ with probability $0.5$.
+There is also a symmetric mixed‐strategy equilibrium in which each agent plays $A$ and $B$ with probability $0.5$.
 Synchronisation Game.
@@
 This creates two pure Nash equilibria along the diagonal of the payoff matrix (Figure 2(b)), and incentivises uniform behaviour across agents.
-$N$-Agent Extension.
+NN-Agent Extension.
 Both games naturally scale to $n$ agents and $n$ possible actions.
@@
 Temporal Extension.
-To model sequential decision-making, we extend each normal-form game into a repeated Markov game, where the state at time $t$ is the joint action at time $t-1$.
+To model sequential decision-making, we extend each normal‑form game into a repeated Markov game, where the state at time $t$ is the joint action at time t1t-1.
 At each step $t$ all agents observe $a^{t-1}$, choose $a_{i}^{t}$, and receive the original Specialisation or Synchronisation payoff.
@@
 In the Specialisation game, an agent’s payoff peaks when it selects a unique action, and then decays as when actions are shared.
-In the Synchronisation game, payoffs follow a hyperbolic scale $1/(n-k+1)$, reaching maximum only under full consensus, thereby incentivising alignment.
+In the Synchronisation game, payoffs follow a hyperbolic scale 1/(nk+1)1/(n-k+1), reaching maximum only under full consensus, thereby incentivising alignment.
 E.3 Proof that FuPS cannot represent the optimal policy in the two-player Specialisation Game
@@
 $=\sum_{i,j}\mathbf{J}_{i}\;\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})\;\mathbf{J}_{j}^{\!\top}$ (pull deterministic matrices out of $\operatorname{Cov}$)
-Equation (LABEL:eq:variance_split_detail) makes explicit that all trajectory-induced covariance is captured $\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})$, while the agent-conditioned Jacobians $\mathbf{J}_{i}$ remain trajectory noise-free.
+Equation (LABEL:eq:variance_split_detail) makes explicit that all trajectory‐induced covariance is captured $\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})$, while the agent-conditioned Jacobians $\mathbf{J}_{i}$ remain trajectory noise-free.
 Mini-batch update and covariance.
@@
 In practice, NoPS incurs additional data transfer and update costs, widening the efficiency gap.
+F.5 Sensitivity Analysis
+We run a sensitivity analysis on the HyperMARL’s hyperparameters in the 20-agent SMAX scenario with recurrent MAPPO, a challenging setting for optimisation. The results are shown Table 5.
+Findings.
+  • Similarly to FuPS, and in general, deep RL methods like PPO, the learning rate is an important parameter for performance.
+  • Agent embedding size is an important hyperparameter.
+In these results, a smaller embedding size outperformed a larger one.
+This could be due to the homogeneous nature of some SMAX tasks, where with smaller embeddings, it could be easier to learn similar agent embeddings and hence similar behaviours.
+This suggests the embedding size could correspond to the task’s diversity requirements.
+  • Other parameters, such as the width, have limited impact beyond a modest size.
+Table 5: IQM and 95% CI of mean win rate across 5 seeds for the 20-agent SMAX scenario with recurrent MAPPO.
+Bold indicates the highest IQM score; * indicates scores whose 95% confidence intervals overlap with the highest score.
+Method Embedding Dim Hidden Dims LR 0.0001 LR 0.0003 LR 0.0005
+FuPS N/A N/A 0.1067 (0.0513, 0.2128) 0.3053 (0.2041, 0.4769) *0.4213 (0.3623, 0.4800)
+HyperMARL 4 16 0.0799 (0.0250, 0.1190) 0.2946 (0.1143, 0.3778) 0.4455 (0.3443, 0.5692)
+64 0.1233 (0.0750, 0.2162) 0.3338 (0.2692, 0.4286) *0.3765 (0.3281, 0.4107)
+128 0.1147 (0.0270, 0.1842) *0.3787 (0.3509, 0.4603) 0.3327 (0.2931, 0.3548)
+64, 64 0.1781 (0.1628, 0.2075) *0.3498 (0.2807, 0.4500) 0.1944 (0.1268, 0.3077)
+HyperMARL 64 16 0.1109 (0.0256, 0.2093) 0.3198 (0.2353, 0.3934) *0.3483 (0.2581, 0.4308)
+64 0.1360 (0.0303, 0.2821) 0.1191 (0.0517, 0.1964) 0.1155 (0.0794, 0.1500)
+128 0.1193 (0.0750, 0.1579) 0.1036 (0.0517, 0.1594) 0.1106 (0.0455, 0.1587)
+64, 64 0.1353 (0.0541, 0.1818) 0.1283 (0.0526, 0.2258) 0.0860 (0.0345, 0.1343)
 Appendix G Experiment Details
@@
 (d) SMAX.
+(e) BPS.
 Figure 17: Multi-Agent environments used in our experiments.
@@
 FuPS baselines have been shown optimal for these settings [53, 17] indicating homogeneous behaviour is preferred here.
+Blind-Particle Spread (BPS) [11] Blind-Particle Spread (BPS) is a modified variant of the Multi-Agent Particle Environment (MPE) Simple Spread task [30] with 15 to 30 agents.
+The environment contains landmarks of different colours, and each agent is assigned a colour.
+Agents are ’blind’ as they cannot observe their own assigned colour or the colours of other agents.
+They must infer the correct landmark to navigate towards based only on their observations and rewards.
+The number of colours thus dictates the number of distinct roles or behaviours the team must learn to successfully complete the task.
 Table 6: Baseline Methods Selection and Justification.
@@
 Included to evaluate HyperMARL’s competitiveness against a method with ensemble critics and diversity loss, in an off-policy setting.
+SePS [11] Selective Parameter Sharing Although this requires a pre-training phase, it is a strong baseline for parameter sharing approaches.
+Tested in: Blind-Particle Spread environments, with 15 to 30 agents.
 SEAC [12] Shared Experience Focuses primarily on experience sharing rather than parameter sharing architecture, falling outside our research scope.
-SePS [10] Selective Parameter Sharing Requires pretraining phase, which extends beyond the scope of our current study focused on end-to-end learning approaches.
 CDAS [26] Intrinsic Reward Only implemented for off-policy methods and has been demonstrated to underperform FuPS/NoPS architectures [17], making it less suitable for our primary on-policy comparisons.
@@
 G.2 HyperMARL Architecture Details
-For the Dispersion and Navigations results (Sec.
+For the Dispersion and Navigation results (Sec.
 5.2) we use feedforward architectures, where we use HyperMARL to generate both the actor and critic networks.
@@
     • For Navigation (5.2), SMAX (5.3), and MaMuJoCo (5.2), we run 5 seeds and train for 10 million timesteps, consistent with the baselines.
+    • For Blind-Particle Spread (BPS), we run 5 seeds and train for 20 million timesteps, consistent with baselines.
   • Evaluation:
@@
 Appendix H Detailed Results
-H.1 Kaleidoscope Off-Policy Comparison Details
+H.1 Comparison with Kaleidoscope using Off-Policy – MATD3
 Our comparison with Kaleidoscope [28], mentioned in Section 5, is conducted using off-policy methods due to its original design.
@@
 Bold indicates the highest IQM score; * indicates scores whose confidence intervals overlap with the highest.
-Although Kaleidoscope employs an ensemble of five critics and an explicit diversity loss, HyperMARL (using a standard MATD3 setup with two critics) achieved competitive results without these additional mechanisms.
-Environment Ind. Actors (MATD3) HyperMARL (MATD3) Kaleidoscope (MATD3)
-Ant-v2 5270.38 (4329.73, 5719.78) *5886.58 (5840.00, 5920.66) 6160.70 (5798.02, 6463.83)
-HalfCheetah-v2 *6777.04 (3169.11, 8233.94) 7057.44 (3508.70, 8818.11) *6901.00 (3609.73, 8192.38)
-Walker2d-v2 *5771.87 (5144.84, 8103.34) 7057.68 (5976.50, 8166.09) *6664.32 (5408.95, 8828.11)
-Swimmer-v2-10x2 *453.74 (427.24, 487.86) 465.91 (410.82, 475.77) *462.48 (444.22, 475.64)
-H.2 Dispersion Detailed Results
-H.2.1 Interval Esimates Dispersion
+Although Kaleidoscope employs an ensemble of five critics and an explicit diversity loss, HyperMARL (using a standard MATD3 setup with two critics) achieves competitive results without these additional mechanisms.
+Environment Ind. Actors, Shared Critic Kaleidoscope HyperMARL
+Ant-v2 5270.38 (4329.73, 5719.78) 6160.70 (5798.02, 6463.83) *5886.58 (5840.00, 5920.66)
+HalfCheetah-v2 *6777.04 (3169.11, 8233.94) *6901.00 (3609.73, 8192.38) 7057.44 (3508.70, 8818.11)
+Walker2d-v2 *5771.87 (5144.84, 8103.34) *6664.32 (5408.95, 8828.11) 7057.68 (5976.50, 8166.09)
+Swimmer-v2-10x2 *453.74 (427.24, 487.86) *462.48 (444.22, 475.64) 465.91 (410.82, 475.77)
+H.2 Comparison with Selective Parameter Sharing (SePS) - A2C
+We benchmark HyperMARL against Selective Parameter Sharing (SePS) [11], a method that shares weights among pre-determined groups of similar agents.
+SePS identifies these groups by clustering agent trajectories via an autoencoder during a pre-training phase.
+Following the original authors’ setup, we use their provided source code and hyperparameters, evaluating all A2C [32] variants on the Blind Particle Spread (BPS) benchmarks (v1-4), which scale up to 30 agents.
+For HyperMARL, we sweep the learning rates (1e-4, 3e-4) and hypernetwork hidden dimension (16, 64) per scenario.
+As shown in Table 8, HyperMARL demonstrates competitive performance against both NoPS and SePS in these challenging settings, which feature up to 30 agents and five distinct agent roles.
+Crucially, HyperMARL achieves this result without resorting to agent-specific weights (like NoPS) or requiring a pre-training and clustering pipeline (like SePS).
+Table 8: Mean episode reward in Blind Particle Spread for A2C variants.
+We report the IQM of the final total reward (with 95% confidence intervals) across 5 seeds on the BPS environments.
+Bold indicates the highest IQM score; * indicates scores whose 95% confidence intervals overlap with the highest score.
+HyperMARL achieves results competitive with SePS without requiring a pre-training phase or explicit agent clustering.
+Environment NoPS FuPS+ID SePS HyperMARL
+BPS-1 (15 agents, 3 groups, 5–5–5) -216.8 (-227.7, -197.8) -228.2 (-247.5, -213.3) *-201.8 (-221.4, -186.8) -190.8 (-220.0, -178.5)
+BPS-2 (30 agents, 3 groups, 10–10–10) *-415.4 (-459.4, -366.3) -429.7 (-507.1, -411.7) *-407.1 (-453.7, -387.6) -397.8 (-423.7, -376.6)
+BPS-3 (30 agents, 5 groups, 6–6–6–6–6) -403.4 (-421.0, -381.0) -835.2 (-1445.6, -534.1) *-422.1 (-466.2, -387.6) *-417.7 (-448.2, -381.1)
+BPS-4 (30 agents, 5 uneven groups, 2–2–2–15–9) *-410.8 (-436.6, -368.3) -780.5 (-1044.1, -593.8) *-411.6 (-446.6, -346.7) -389.5 (-457.6, -366.0)
+H.3 Dispersion Detailed Results
+H.3.1 Interval Estimates Dispersion
 (a) IPPO
@@
 Figure 19: We see that even if we run MAPPO-FuPS on Dispersion for 40 million timesteps (double the timesteps of MLP Hypernetwork), it converges to suboptimal performance.
-H.3 Detailed MAMujoco Plots
+H.4 Detailed MAMujoco Plots
 (a) Humanoid-v2 17x1
@@
 HyperMARL performs comparably to these baselines, and is the only method with shared actors to demonstrate stable learning in the notoriously difficult 17-agent Humanoid environment.
-H.4 Detailed Navigation Plots
+H.5 Detailed Navigation Plots
 (a) 8 Agents, Same Goal
@@
 Legend shown at bottom applies to all subplots.
-H.5 Interval Esimates - SMAX
+H.6 Interval Estimates - SMAX
 (a) 2s3z
@@
 HyperMARL performs comparably to FuPS baselines across all environments, demonstrating its effectiveness in tasks requiring homogeneous behaviours and using recurrency.
-H.6 Additional Ablations
+H.7 Additional Ablations
 Figure 23: Ablation results comparing HyperMARL with its variants in Dispersion. The results highlight that gradient decoupling is essential for maintaining HyperMARL’s performance.
@@
 HYPERNET_HIDDEN_DIMS 64 64
-Table 12: DiCo Algorithm $SND\_des$ Hyperparameter
+Table 15: DiCo Algorithm SND_desSND\_des Hyperparameter
 Goal Configuration Number of Agents SND_des
@@
 OBS_WITH_AGENT_ID - True
-Table 21: Hyperparameter Sweeps and Final Values for Different Maps in SMAX. H- refers to HyperMARL MLP Hypernetworks.
+Table 24: Hyperparameter Sweeps and Final Values for Different Maps in SMAX. H‐ refers to HyperMARL MLP Hypernetworks.
 Map Algorithm LR Range Chosen LR HNET Embedding Dim HNET Hidden Dims
@@
 Experiments 64-Core Processor + NVIDIA RTX A4500
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Please refer to Sections 35 and 6.1.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Please see Sections A and F.3.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: Please see Section E.3.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Please see Sections 4 and 1. We also provide all the code.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We provide all the code and scripts to reproduce results.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Please see Sections 5 and App. I.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Following best practices, we report statistical significance in all results in Section 5.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Please see Section J.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: Read and confirmed.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: See App. B.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification:
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [N/A]
+  59. Justification:
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification:
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification:
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification:
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification:
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.